### v1

# Elliptic curves and Fourier coefficients of meromorphic modular forms

Abstract.
We discuss several congruences satisfied by the coefficients of meromorphic modular forms, or equivalently, $p$-adic behaviors of meromorphic modular forms under the $U_{p}$ operator, that are summarized from numerical experiments, connecting meromorphic modular forms to symmetric powers of elliptic curves. We also provide heuristic explanations for these congruences as well as prove some of them using hypergeometric functions and the Borcherds–Shimura lift.
Key words and phrases:
meromorphic modular forms, symmetric powers of elliptic curves, magnetic modular forms, supercongruences
2020 Mathematics Subject Classification:
Primary 11F37, 11F30, 11F33, 11F23; Secondary 11G05, 11G15, 11F27, 33C20
1. Introduction
Coefficients of modular forms have always been one of the central topics in the field of modular forms. The vast majority of work in this field is restricted to the study of holomorphic or weakly holomorphic modular forms. In this paper, we instead turn to meromorphic modular forms. The main objects of interest are meromorphic modular forms of level $1$ with precisely one pole at a non-cuspidal point. Specifically, we will focus on modular forms of the form
(1) $A_{1}\cdot\frac{g}{j-c}+A_{2}\cdot\frac{g}{(j-c)^{2}}+\cdots+A_{r}\cdot\frac{g}{(j-c)^{r}},$
where $c\in\mathbb{Q}$, each $A_{i}\in\mathbb{Q}$, and $g\in\mathbb{Z}[[q]]$ is a holomorphic modular form of level $1$ such that $\frac{g}{j-c}$ has a pole. We will discuss the (conjectural) $p$-adic properties possessed by these meromorphic modular forms as well as the heuristic reasons for those properties. In particular, we hope to convince readers that these meromorphic modular forms of weight $k$ behave similarly to symmetric (k2)(k-2)-nd powers of elliptic curves. Indeed, the paper by Brown–Fonseca [BF25] already evidences such a connection.
In this introduction, we will motivate the subject through several phenomena satisfied by the modular form $\frac{E_{4}}{j-c}$ for some $c\in\mathbb{Q}$, some provable and some conjectural. We start with the cases when $c\in\{0,1728\}$, i.e., $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$. Here we make a definition that a (general) modular form $\sum_{n=1}^{\infty}a_{n}q^{n}\in\mathbb{Z}[[q]]$ is called magnetic if $n{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}$, and specifically $r$-magnetic for some $r\in\mathbb{Z}_{\geq 0}$ if $n^{r}{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}$ for all $n\in\mathbb{Z}^{+}$. We have the following result by Li–Neururer [LN19, Theorem 1.5] and Paşol–Zudilin [PZ22, Theorem 1].
Theorem 1.1.
Both $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$ are $1$-magnetic.
The first numerically observed example of a magnetic modular form appeared in Broadhurst–Zudilin [BZ19, Conjecture 4] with motivation coming from the Hall effect in electromagnetism. The magnetic property of that example, together with that of $\frac{64E_{4}}{j}$, was proved in Li–Neururer [LN19]. Their proof uses the Borcherds–Shimura lift [Bor98]: one identifies the half-integral weight preimage, proves certain divisibility properties of the coefficients of the preimage, and deduces the magnetic property of the image. This method was later generalized by Paşol–Zudilin [PZ22] to show the magnetic property of other meromorphic modular forms, including $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$, all of which have poles at CM points. For a detailed discussion of magnetic modular forms, we also refer to the paper by Bönisch–Duhr–Maggio [BDM24] and the references therein.
In fact, by modifying the method of using the Borcherds–Shimura lift, one can show that both $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$ also satisfy the following supercongruences  [^1] Here the word “supercongruence” refers to the type of congruence with modulus $p^{rl}$ for some $r>1$., which almost imply their $1$-magnetic property and in particular confirm that they are not $2$-magnetic. In Section 4, we will illustrate that both the magnetic property and the supercongruences occur in general for meromorphic modular forms with CM poles.
Theorem 1.2.
For all primes $p\geq 5$ and all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}\bigg(\frac{E_{4}}{j}\bigg)$ $\;\equiv\;\big(\tfrac{-3}{p}\big)p{\hskip 1.0pt}a_{np^{l-1}}\bigg(\frac{E_{4}}{j}\bigg)$ $\pmod{p^{3l}}$
$a_{np^{l}}\bigg(\frac{E_{4}}{j-1728}\bigg)$ $\;\equiv\;\big(\tfrac{-4}{p}\big)p{\hskip 1.0pt}a_{np^{l-1}}\bigg(\frac{E_{4}}{j-1728}\bigg)$ $\pmod{p^{3l}}$,
where $(\frac{\cdot}{p})$ denotes the Legendre symbol.
When $c\notin\{0,1728\}$, one observes something different, with elliptic curves coming into play. Let $C/\mathbb{Q}$ be an elliptic curve. For a good prime $p$ of $C$, define
$a_{p}(C)\;:=\;p+1-|C(\mathbb{F}_{p})|.$
It is known that $|a_{p}(C)|\leq 2\sqrt{p}$ by the Hasse–Weil bound. A good prime $p$ of $C$ is called an ordinary (resp. supersingular) prime of $C$ if $p\nmid a_{p}(C)$ (resp. $p{\hskip 1.0pt}|{\hskip 1.0pt}a_{p}(C)$). For $p\geq 5$, this is equivalent to the condition that $a_{p}(C)\neq 0$ (resp. $a_{p}(C)=0$).
Theorem 1.3.
Let $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$ and let $p$ be a good prime of $C$ with $p\geq 5$ and $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then,
$a_{p}\bigg(\frac{E_{4}}{j-j(C)}\bigg)\;\equiv\;a_{p}(C)^{2}\pmod{p}$.
Here the appearance of $a_{p}(C)^{2}$ already hints at the link between $\frac{E_{4}}{j-j(C)}$ and $\mathrm{Sym}^{2}C$, since the trace of Frobenius at $p$ of $\mathrm{Sym}^{2}C$ is $a_{p}(C)^{2}-p$, which is congruent to $a_{p}(C)^{2}$ modulo $p$. When $p$ is supersingular, Theorem 1.3 a priori only implies that $a_{p}(\frac{E_{4}}{j-j(C)})\equiv 0\pmod{p}$, while in fact numerical observations suggest the following stronger congruence.
Conjecture 1.4.
Let $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$ and let $p$ be a supersingular prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then,
$a_{p}\bigg(\frac{E_{4}}{j-j(C)}\bigg)\;\equiv\;0\pmod{p^{2}}$.
In general, the connection between $\frac{E_{4}}{j-j(C)}$ and $\mathrm{Sym}^{2}C$ can be visualized via the following longer congruence, where one considers suitable linear combinations of multiple coefficients of $\frac{E_{4}}{j-j(C)}$.
Conjecture 1.5.
Let $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$ and let $p$ be a good prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Let $\alpha_{p},\beta_{p}$ be two roots of $X^{2}-a_{p}(C)X+p$ and write
$(X-\alpha_{p}^{2})(X-\alpha_{p}\beta_{p})(X-\beta_{p}^{2})\;=:\;X^{3}+c_{p,1}X^{2}+c_{p,2}X+c_{p,3}.$
Write $F=\frac{E_{4}}{j-j(C)}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F)+c_{p,1}{\hskip 1.0pt}a_{np^{l-1}}(F)+c_{p,2}{\hskip 1.0pt}a_{np^{l-2}}(F)+c_{p,3}{\hskip 1.0pt}a_{np^{l-3}}(F)\;\equiv\;0\pmod{p^{3l-3}}$.
Congruences of this form are usually called the Atkin–Swinnerton-Dyer (ASD) congruences. This term originally refers to three-term congruences of the form
$a_{np}-A_{p}{\hskip 1.0pt}a_{n}+\chi(p)p^{k-1}a_{n/p}\;\equiv\;0\pmod{p^{(k-1)(v_{p}(n)+1)}}$.
In general, it also refers to congruences of the form
$a_{np^{l}}+c_{p,1}{\hskip 1.0pt}a_{np^{l-1}}+c_{p,2}{\hskip 1.0pt}a_{np^{l-2}}+\cdots+c_{p,s}{\hskip 1.0pt}a_{np^{l-s}}\;\equiv\;0\pmod{p^{\bullet}}$
for some suitable power of $p$. Atkin and Swinnerton-Dyer [ASD71] first came across this type of congruences when studying the coefficients of some noncongruence modular form, whose result was later generalized by Scholl [Sch85]. In Kazalicki–Scholl [KS16], they also demonstrated certain ASD congruences for weakly holomorphic modular forms, and as an example, showed that $F=E_{4}^{6}/\Delta-1464E_{4}^{3}$ satisfies that
$a_{np}(F)-\tau(p)a_{n}(F)+p^{11}a_{n/p}(F)\;\equiv\;0\pmod{p^{11v_{p}(n)}}$,
where $\tau$ is the Ramanujan tau function. It has also come to the author’s attention that Michael Allen, Ling Long, and Hasan Saad are approaching the ASD congruences for meromorphic modular forms on finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$ by extending the methods of Scholl and Kazalicki–Scholl, which may lead to proofs of some of the conjectures described in this paper. For a comprehensive overview of the ASD congruences, we refer to the paper by Li–Long [LL14] and the references therein.
We will end the list of phenomena with an explicit connection between the $p$-th coefficient of $\frac{E_{4}}{j-c}$ and the truncated hypergeometric sum associated to the datum $\big((\frac{1}{2},\frac{1}{6},\frac{5}{6}),(1,1)\big)$
[…]
where we adopt notations from Section 5.3. This connection also constitutes a key part of the proof of Theorem 1.3.
Theorem 1.6.
Let $c\in\mathbb{Q}$ and let $p\geq 5$ be a prime with $v_{p}(c)=0$. Then,
$a_{p}\bigg(\frac{E_{4}}{j-c}\bigg)$ $\;\equiv\;\big(c(c-1728)\big)^{\frac{p-1}{2}}\cdot\sum_{m=0}^{p-1}\frac{(6m)!}{(m!)^{3}(3m)!}{\hskip 1.0pt}c^{-m}\pmod{p}$.
As discussed earlier, the cases when $c\in\{0,1728\}$ behave differently from the rest. This is because both $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$ have a double pole while $\frac{E_{4}}{j-c}$ for $c\notin\{0,1728\}$ only has a simple pole. In general, the cases when $c\in\{0,1728\}$ usually involve more technicalities. We will thus assume for simplicity that $c\notin\{0,1728\}$ throughout most part of the paper, and only briefly discuss those two excluded cases in Section 7.1. When investigating the $p$-adic properties, we will also assume that $v_{p}(c)=0=v_{p}(c-1728)$ throughout the paper.
The reason that all the phenomena have rather neat formulations in weight $4$ is that there is no cusp form. In general, the coefficients of linear combinations of $\big\{\frac{g}{(j-j(C))^{r}}\big\}_{r}$ are an interaction of the elliptic curve $C/\mathbb{Q}$ and the cusp forms of weight $k$, whose formulation is more complicated. For ease of exposition, we will mostly focus on the cases when $k\in\{4,6,8,10,14\}$ so that there is no influence from cusp forms but that there is still a (canonical) choice for $g$, i.e., the Eisenstein series $E_{k}$ of weight $k$. In Section 7.2, we will describe the way of generalizing the whole setting to other weights.
It should also be noted that while this paper heavily adopts the word “conjectures”, they should be more accurately named as numerical observations, as they are summarized from numerical experiments with these meromorphic modular forms.
Acknowledgements
The author would like to thank many people for useful discussions on this project, including Kilian Bönisch, Steven Charlton, Chi Hong Chow, Henri Cohen, Håvard Damm-Johnsen, Vasily Golyshev, Wen-Ching Winnie Li, Yingkun Li, Ling Long, Danylo Radchenko, Berend Ringeling, Fernando Rodriguez Villegas, Hasan Saad, and Jan Vonk. The author is in particular grateful to Don Zagier and Wadim Zudilin for all the useful ideas on proofs and generalizations.
This project originated from discussions during the International Conference on Modular Forms and $q$-Series at University of Cologne, and during the follow-up workshop at the Hausdorff Research Institute for Mathematics to the trimester program “Periods in Number Theory, Algebraic Geometry and Physics” funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813. Part of the later discussions also took place during the author’s visit to the Abdus Salam International Centre for Theoretical Physics hosted by Vasily Golyshev and Don Zagier, and the author’s visit to Louisiana State University hosted by Ling Long and partially supported by the Simons Foundation grant MP-TSM-00002492.
2. Simple poles
We will start with the simplest case, i.e., the case of meromorphic modular forms with precisely one pole that is simple and at a non-cuspidal point.
Let $k\in\{4,6,8,10,14\}$ and let $C/\mathbb{Q}$ be an elliptic curve (with $j(C)\notin\{0,1728\}$). Write
$F_{k,C}\;:=\;\frac{E_{k}}{j-j(C)}\;=\;\sum_{n=1}^{\infty}a_{n}(F_{k,C})q^{n}.$
Conjecture 2.1.
Let $k\in\{4,6,8,10,14\}$, $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$, and $p$ be a good prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n\in\mathbb{Z}^{+}$,
(2) $a_{np}(F_{k,C})\;\equiv\;a_{p}(C)^{k-2}a_{n}(F_{k,C})\pmod{p}$.
Remark 2.2.
2.1 is a theorem if we further assume that $p\geq 5$, from which Theorem 1.3 follows. In fact, the cases when $k\in\{4,6,8,10,14\}$ all follow from the case when $k=4$, and the proof when $k=4$ depends on the aforementioned connection with hypergeometric functions, namely the congruence
$a_{p}(F_{4,C})$ $\;\equiv\;\big(j(C)(j(C)-1728)\big)^{\frac{p-1}{2}}\cdot\sum_{m=0}^{p-1}\frac{(6m)!}{(m!)^{3}(3m)!}\cdot j(C)^{-m}\pmod{p}$.
We will show all above in Section 5, where we will prove a suitable version of 2.1 for elliptic curves over general number fields.
From a different point of view, 2.1 describes the behavior of $F_{k,C}$ under the $U_{p}$ operator. In particular, it says that $F_{k,C}$, when viewed as a mod $p$ modular form, is a $U_{p}$-eigenform with eigenvalue $a_{p}(C)^{k-2}$. In general, one can also describe the behavior of $F_{k,C}$ under the $U_{p}^{l}$ operator if one considers the supersingular and the ordinary cases separately.
Conjecture 2.3.
Let $k\in\{4,6,8,10,14\}$, $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$, and $p$ be a supersingular prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
(3) $a_{np^{l}}(F_{k,C})\;\equiv\;p^{k-2}a_{np^{l-2}}(F_{k,C})\pmod{p^{(k-1)l-1}}$,
where $a_{m}(\;\cdot\;):=0$ if $m\notin\mathbb{Z}$.
Conjecture 2.4.
Let $k\in\{4,6,8,10,14\}$, $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$, and $p$ be an ordinary prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
(4) $a_{np^{l}}(F_{k,C})\;\equiv\;u_{p}(C)^{k-2}a_{np^{l-1}}(F_{k,C})\pmod{p^{l}}$,
where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit.
Moreover, if $C/\mathbb{Q}$ has CM, then for all $n,l\in\mathbb{Z}^{+}$,
(5) $a_{np^{l}}(F_{k,C})\;\equiv\;u_{p}(C)^{k-2}a_{np^{l-1}}(F_{k,C})\pmod{p^{(k-1)l}}$.
Remark 2.5.
While Remark 2.2 already presents a connection between meromorphic modular forms and hypergeometric functions, the type of congruences that appear in 2.3 and 2.4 may also remind one of hypergeometric (super)congruences.
In the case of CM elliptic curves, it is possible to unify the supersingular and ordinary dichotomy via a single, but slightly weaker, congruence, in the form of the three-term ASD congruence as mentioned in Section 1. There are two viewpoints in the CM setting: one can interpret everything in terms of CM elliptic curves or in terms of negative discriminants. We will elaborate on both viewpoints here.
Conjecture 2.6.
Let $k\in\{4,6,8,10,14\}$. Let $C/\mathbb{Q}$ be an elliptic curve with CM by some imaginary quadratic field $K$ and suppose that $j(C)\notin\{0,1728\}$. Let $\psi_{C}$ be the Hecke character of $K$ corresponding to $C$ and let $\Theta$ be the Hecke eigenform of weight k1k-1 corresponding to the Hecke character $\psi_{C}^{k-2}$. Let $p$ be a good prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
(6) $a_{np^{l}}(F_{k,C})-a_{p}(\Theta){\hskip 1.0pt}a_{np^{l-1}}(F_{k,C})+\chi_{K}(p)p^{k-2}a_{np^{l-2}}(F_{k,C})\;\equiv\;0\pmod{p^{(k-1)l-1}}$,
where $\chi_{K}$ is the quadratic character corresponding to the extension $K/\mathbb{Q}$ and $a_{m}(\;\cdot\;):=0$ if $m\notin\mathbb{Z}$.
Remark 2.7.
By Deuring’s criterion (see [Deu41] or [Sil94, Chapter II, Exercise 2.30]), a good prime $p$ of $C$ is ordinary (resp. supersingular) if and only if $\chi_{K}(p)=1$, i.e. $p$ splits in $K$ (resp. $\chi_{K}(p)=-1$, i.e., $p$ is inert in $K$). The coefficient $a_{p}(\Theta)$ for a good prime $p$ of $C$ can then be written down explicitly as
$a_{p}(\Theta)\;=\;\begin{cases}\alpha_{p}^{k-2}+\beta_{p}^{k-2}&\text{ if }\chi_{K}(p)=1,\\ 0&\text{ if }\chi_{K}(p)=-1,\end{cases}$
where $\alpha_{p}$ and $\beta_{p}$ are the two roots of $X^{2}-a_{p}(C)X+p$.
From this explicit description, one can readily deduce 2.6 from the supersingular (2.3) and ordinary (2.4) dichotomy. In the supersingular case, i.e., $\chi_{K}(p)=-1$, 2.6 reduces to that
$a_{np^{l}}(F_{k,C})\;\equiv\;p^{k-2}a_{np^{l-2}}(F_{k,C})\pmod{p^{(k-1)l-1}}$,
which follows from 2.3. In the ordinary case, i.e., $\chi_{K}(p)=1$, 2.6 reduces to that
$a_{np^{l}}(F_{k,C})-(\alpha_{p}^{k-2}+\beta_{p}^{k-2})a_{np^{l-1}}(F_{k,C})+p^{k-2}a_{np^{l-2}}(F_{k,C})\;\equiv\;0\pmod{p^{(k-1)l-1}}$.
If we let $u_{p}(C)$ denote the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit, then we can rewrite the equation above as
$\bigg(a_{np^{l}}(F_{k,C})-u_{p}(C)^{k-2}a_{np^{l-1}}(F_{k,C})\bigg)$
$\hskip 42.67912pt+\bigg(\frac{p}{u_{p}(C)}\bigg)^{k-2}\bigg(a_{np^{l-1}}(F_{k,C})-u_{p}(C)^{k-2}a_{np^{l-2}}(F_{k,C})\bigg)\;\equiv\;0\pmod{p^{(k-1)l-1}}$,
which follows from 2.4.
Now, we turn to the negative discriminants. Let $D<0$ (with $D\neq-3,-4$) be a discriminant with class number $1$, $\mathcal{O}_{D}$ be the imaginary quadratic order of discriminant $D$, and $\alpha_{D}\in\mathcal{H}$ be a fixed CM point of discriminant $D$. For $k\in\{4,6,8,10,14\}$, write
$F_{k,D}\;:=\;\frac{E_{k}}{j-j(\alpha_{D})}\;=\;\sum_{n=1}^{\infty}a_{n}(F_{k,D})q^{n}.$
Conjecture 2.8.
Let $k\in\{4,6,8,10,14\}$ and let D<4D<-4 be a discriminant with class number $1$. Consider the CM modular form (in fact Hecke eigenform)
$\Theta\;:=\;\Theta_{k-1,D}\;:=\;\frac{1}{2}\sum_{\alpha\in\mathcal{O}_{D}}\alpha^{k-2}q^{|\mathrm{N}(\alpha)|}$
of weight k1k-1. Let $p$ be a prime with $p\nmid D$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,D})-a_{p}(\Theta)a_{np^{l-1}}(F_{k,D})+\big(\tfrac{D}{p}\big)p^{k-2}a_{np^{l-2}}(F_{k,D})\;\equiv\;0\pmod{p^{(k-1)l-1}}$,
where $a_{m}(\;\cdot\;):=0$ if $m\notin\mathbb{Z}$.
Remark 2.9.
The coefficients $a_{p}(\Theta)$ for primes $p\nmid D$ can be written down explicitly as
$a_{p}(\Theta)\;=\;\begin{cases}\pi^{k-2}+\overline{\pi}^{k-2}=\mathrm{tr}_{K/\mathbb{Q}}(\pi^{k-2})&\text{ if }\big(\tfrac{D}{p}\big)=1,\\ 0&\text{ if }\big(\tfrac{D}{p}\big)=-1,\end{cases}$
where $K=\mathbb{Q}(\sqrt{D})$ and $p$ with $\big(\frac{D}{p}\big)=1$ factors as $p=\pi\overline{\pi}$ in $\mathcal{O}_{D}$.
One can also write 2.4 in the CM case explicitly in terms of negative discriminants.
Conjecture 2.10.
Let $k\in\{4,6,8,10,14\}$ and let $D<0$ be a discriminant with class number $1$. Let $p$ be a prime with $\big(\frac{D}{p}\big)=1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$, and write $p=\pi\overline{\pi}$ for some $\pi\in\mathcal{O}_{D}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,D})$ $\;\equiv\;\pi^{k-2}a_{np^{l-1}}(F_{k,D})\pmod{\overline{\pi}^{(k-1)l}}$
$a_{np^{l}}(F_{k,D})$ $\;\equiv\;\overline{\pi}^{k-2}a_{np^{l-1}}(F_{k,D})\pmod{\pi^{(k-1)l}}$.
3. Poles of higher order
As suggested constantly, meromorphic modular forms of weight $k$ with precisely one pole at a non-cuspidal point should behave similarly to the symmetric (k2)(k-2)-nd power of the elliptic curve corresponding to the non-cuspidal pole. The discrepancy, however, is that modular forms typically correspond to $2$-dimensional objects from representation theory, while symmetric (k2)(k-2)-nd powers of elliptic curves are (k1)(k-1)-dimensional. Indeed, on the modular form side, the “correct” space of meromorphic modular forms to consider should be
$\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(C)},\;\frac{E_{k}}{(j-j(C))^{2}},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}\bigg\}$
(for $k\in\{4,6,8,10,14\}$ and $j(C)\notin\{0,1728\}$). This is a (k1)(k-1)-dimensional space and should “correspond” to the (k1)(k-1)-dimensional space $\mathrm{Sym}^{k-2}C$ on the elliptic curve side.
The consideration of this particular space is motivated by several reasons. It is easy to check that for $r\geq k$, $\frac{E_{k}}{(j-j(C))^{r}}$ can be written as a linear combination of $\frac{E_{k}}{j-j(C)},\frac{E_{k}}{(j-j(C))^{2}},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}$, and $\mathcal{D}^{k-1}f$ for some modular form $f$ of weight 2k2-k, where $\mathcal{D}=q\frac{d}{dq}$ is the usual derivative. From the viewpoint of the magnetic property, the modular form $\mathcal{D}^{k-1}f$ here is trivially (k1)(k-1)-magnetic, so it suffices to consider the denominator only up to the (k1)(k-1)-st power. More importantly, the discussion above yields a non-canonical isomorphism (for $k\in\{4,6,8,10,14\}$ and $j(C)\notin\{0,1728\}$)
$\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(C)},\;\frac{E_{k}}{(j-j(C))^{2}},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}\bigg\}\;\cong\;\frac{S_{k}^{\mathrm{mero},C}}{\mathcal{D}^{k-1}M_{2-k}^{\mathrm{mero},C}},$
where $M_{k}^{\mathrm{mero},C}$ denotes the space of meromorphic modular forms of level $1$ and weight $k$ with precisely one pole at (the point on $\mathcal{H}$ corresponding to) $C$, and $S_{k}^{\mathrm{mero},C}$ denotes its subspace consisting of elements having vanishing constant terms in the Fourier expansions. In Kazalicki–Scholl [KS16], the similar quotient where $C$ is replaced by the cusp $\infty$ already occurred and was used to prove the ASD congruences satisfied by weakly holomorphic modular forms. Specifically, they provided a cohomological interpretation of the quotient and transferred the ASD congruences to the action of the Frobenius at $p$ on the cohomology side. In our case, a cohomological interpretation of the quotient also exists and was discussed in Brown–Fonseca [BF25], where they also related it to $\mathrm{Sym}^{k-2}C$. This suggests that these meromorphic modular forms should be linked to symmetric powers of elliptic curves through the ASD congruences.
The guiding philosophy in our case is that the action of $U_{p}$ on the modular form side  [^2] Indeed, there is no $U_{p}$ action on the space of modular forms that is considered, since the $U_{p}$ operator increases the level and shifts the pole. Still, we will formally define the $U_{p}$ operator by $\sum a_{n}q^{n}\mapsto\sum a_{pn}q^{n}$. should correspond to the action of the Frobenius at $p$ on the elliptic curve side.
$\mathrm{Span}\bigg\{\dfrac{E_{k}}{j-j(C)},\;\dfrac{E_{k}}{(j-j(C))^{2}},\ldots,\dfrac{E_{k}}{(j-j(C))^{k-1}}\bigg\}$$\mathrm{Sym}^{k-2}C$$\scriptstyle{U_{p}}$$\scriptstyle{\mathrm{Frob}_{p}}$
Let $p$ be a good prime of $C$ and let $P_{p}(X)$ denote the characteristic polynomial of the Frobenius at $p$ on $\mathrm{Sym}^{k-2}C$. Then,
$P_{p}(X)$ $\;=\;(X-\alpha_{p}^{k-2})(X-\alpha_{p}^{k-3}\beta_{p})\cdots(X-\alpha_{p}\beta_{p}^{k-3})(X-\beta_{p}^{k-2})$
$\;=:\;\sum_{i=0}^{k-1}c_{p,k-1-i}^{(k)}X^{i}\;=\;X^{k-1}+\sum_{i=0}^{k-2}c_{p,k-1-i}^{(k)}X^{i},$
where $\alpha_{p}$ and $\beta_{p}$ are the two roots of $X^{2}-a_{p}(C)X+p$. The expectation is that $F=\frac{E_{k}}{(j-j(C))^{r}}$ should satisfy the following ASD congruence
$a_{np^{l}}(F)+c_{p,1}^{(k)}{\hskip 1.0pt}a_{np^{l-1}}(F)+c_{p,2}^{(k)}{\hskip 1.0pt}a_{np^{l-2}}(F)+\cdots+c_{p,k-1}^{(k)}{\hskip 1.0pt}a_{np^{l-k+1}}(F)\;\equiv\;0\pmod{p^{\bullet}}$
for some power of $p$ with exponent depending on $k,l$, and $r$.
For $r\in\mathbb{Z}^{+}$, write
$F_{k,C}^{(r)}\;:=\;\frac{E_{k}}{(j-j(C))^{r}}\;=\;\sum_{n=1}^{\infty}a_{n}(F_{k,C}^{(r)})q^{n}.$
Then, one has the following conjecture.
Conjecture 3.1.
Let $k\in\{4,6,8,10,14\}$, $1\leq r\leq k-1$, and $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$. Let $p$ be a good prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,C}^{(r)})+c_{p,1}^{(k)}{\hskip 1.0pt}a_{np^{l-1}}(F_{k,C}^{(r)})+\cdots+c_{p,k-1}^{(k)}{\hskip 1.0pt}a_{np^{l-k+1}}(F_{k,C}^{(r)})\;\equiv\;0\pmod{p^{(k-1)l-\frac{1}{2}(k-3)k-r}}$,
where $a_{m}(\;\cdot\;)=0$ if $m\notin\mathbb{Z}$.
Remark 3.2.
If one simply considers the cases when $l\geq k-1$ and writes s=lk+1s=l-k+1, then 3.1 can be rewritten as for all $s\in\mathbb{Z}_{\geq 0}$,
$U_{p}^{s}{\hskip 1.0pt}P_{p}(U_{p})(F_{k,C}^{(r)})\;\equiv\;0\pmod{p^{(k-1)s+\frac{1}{2}(k^{2}-k+2)-r}}$.
On the other hand, one may also consider $V_{p}:\sum a_{n}q^{n}\mapsto\sum a_{n}q^{pn}$ and rewrite 3.1 as for all $l\in\mathbb{Z}^{+}$,
$U_{p}^{l}{\hskip 1.0pt}P_{p}^{*}(V_{p})(F_{k,C}^{(r)})\;\equiv\;0\pmod{p^{(k-1)l-\frac{1}{2}(k-3)k-r}}$,
where $P_{p}^{*}$ is the reciprocal polynomial of $P_{p}$, i.e., $P_{p}^{*}(X)=X^{k-1}P_{p}(X^{-1})$. In view of Kazalicki–Scholl [KS16], it appears that the formulation using $V_{p}$ may be the “correct” one, since their proof of the ASD congruences involves identifying a $V_{p}$-related operator with the Frobenius at $p$. It may thus be the case that we should also compare the $V_{p}$ operator on the modular form side with the Frobenius at $p$ on the elliptic curve side in our case.
If one separates the supersingular and ordinary primes as before, then one also observes a generalization of 2.3. However, generalizations of 2.4 seem to only exist in the CM case, which will be discussed in Section 4.
Conjecture 3.3.
Let $k\in\{4,6,8,10,14\}$, $1\leq r\leq k-1$, and $C/\mathbb{Q}$ be an elliptic curve. Let $p$ be a supersingular prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,C}^{(r)})\;\equiv\;p^{k-2}a_{np^{l-2}}(F_{k,C}^{(r)})\pmod{p^{(k-1)l-r}}$,
where $a_{m}(\;\cdot\;):=0$ if $m\notin\mathbb{Z}$.
Remark 3.4.
When $p$ is supersingular, $F_{k,C}^{(r)}$ may be viewed as an overconvergent $p$-adic modular form, since the pole lies in the supersingular locus. Indeed, for $l\geq 2$, 3.3 leads to the consideration of $U_{p}^{l-2}(U_{p}-p^{k-2})(F_{k,C}^{(r)})$. Here one observes that there is no old cusp form by our choice of $k$ and the operator $U_{p}-p^{k-2}$ annihilates all the $p$-new cusp forms of weight $k$. Hence, heuristically one should be able to write $(U_{p}-p^{k-2})(F_{k,C}^{(r)})$ as an asymptotic sum of $U_{p}$-eigenforms with slopes $\geq k-1$, which then also supports the conjectural congruence satisfied by $U_{p}^{l-2}(U_{p}-p^{k-2})(F_{k,C}^{(r)})$. (We refer to [Cal] and [GM95] for discussions on the asymptotic expansions of overconvergent modular forms using $U_{p}$-eigenforms.)
Remark 3.5.
3.3 implies 3.1 in the supersingular case. We sketch a proof here for $l\geq k-1$. For supersingular $p$, the characteristic polynomial $P_{p}(X)$ is given by
$P_{p}(X)\;=\;$ $\prod_{i=0}^{k-2}\big(X-(\sqrt{-p})^{k-2-i}(-\sqrt{-p})^{i}\big)$
$\;=\;$ $(X-(-p)^{\frac{k-2}{2}})(X^{2}-p^{k-2})^{\frac{k-2}{2}}$
$\;=\;$ $\bigg(\sum_{i=0}^{\frac{k-4}{2}}\binom{\frac{k-4}{2}}{i}(-p^{k-2})^{i}X^{k-3-2i}-(-p)^{\frac{k-2}{2}}\sum_{i=0}^{\frac{k-4}{2}}\binom{\frac{k-4}{2}}{i}(-p^{k-2})^{i}X^{k-4-2i}\bigg)(X^{2}-p^{k-2}).$
It then follows from 3.3 and direct computations that for all $s\in\mathbb{Z}_{\geq 0}$,
$U_{p}^{s}{\hskip 1.0pt}P_{p}(U_{p})(F_{k,C}^{(r)})\;\equiv\;0\pmod{p^{(k-1)s+\frac{1}{2}(k^{2}-k+2)-r}}$,
which implies 3.1.
We will finish this section with two comments. The first one is that one should be able to generalize the whole setting to elliptic curves over number fields. That is, one considers an elliptic curve $C$ over a number field $L$. In this case, one should then consider congruences modulo powers of $\mathfrak{p}$, and the action of the Frobenius at $\mathfrak{p}$ on the elliptic curve side should correspond to the action of $U_{\mathrm{N}(\mathfrak{p})}$ on the modular form side, where $\mathfrak{p}$ is some prime of $L$ and $\mathrm{N}(\mathfrak{p})$ denotes its norm. The second one is that the connection between meromorphic modular forms and symmetric powers of elliptic curves should be rather motivic  [^3] This is also why we always write $\mathrm{Sym}^{k-2}C$ instead of $\mathrm{Sym}^{k-2}H^{1}(C)$ with some specified cohomology, since we think $\mathrm{Sym}^{k-2}C$ as being the motive attached to the symmetric power., as discussed in Brown–Fonseca [BF25]. In particular, the congruences that we observe here should be viewed as visualizations of the motivic connection at finite places. While this is not necessarily clear to the author, it may be reasonable to expect visualizations at infinite places linking special values of $L$-functions of symmetric powers of elliptic curves to “periods” of meromorphic modular forms (see also [BF25, Section 9.2]).
4. The CM case and magnetic modular forms
We will now specialize to the CM case. Let $D<0$ (with $D\neq-3,-4$) be a discriminant with class number $1$, $\mathcal{O}_{D}$ be the imaginary quadratic order of discriminant $D$, and $\alpha_{D}\in\mathcal{H}$ be a fixed CM point with discriminant $D$. Let $C/\mathbb{Q}$ be an elliptic curve with CM by $\mathcal{O}_{D}$ so that $j(C)=j(\alpha_{D})$ and let $\psi_{C}$ be the corresponding Hecke character of $K:=\mathbb{Q}(\sqrt{D})$. For the convenience of the later discussion, we will choose $C/\mathbb{Q}$ in a way that all of its bad primes divide $D$. By Deuring’s criterion, the following are equivalent for a prime $p$ with $p\nmid D$:
  1. (i)
    $p$ is a supersingular prime of $C$ (resp. $p$ is an ordinary prime of $C$);
  2. (ii)
    $p$ is inert in $\mathcal{O}_{D}$ (resp. $p$ splits in $\mathcal{O}_{D}$);
  3. (iii)
    $(\frac{D}{p})=-1$ (resp. $(\frac{D}{p})=1$).
We will thus use these three conditions interchangeably.
Now, on the elliptic curve side, we have the following splitting (as Galois representations of $K$)
(7) $\mathrm{Res}_{K/\mathbb{Q}}(\mathrm{Sym}^{k-2}C)\;\cong\;\mathrm{Res}_{K/\mathbb{Q}}\big(\mathrm{Sym}^{k-2}(\mathrm{Ind}_{K/\mathbb{Q}}\psi_{C})\big)\;\cong\;\bigoplus_{\begin{subarray}{c}a+b=k-2\\ 0\leq a,b\leq k-2\end{subarray}}\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}.$
Let $p$ be a good prime of $C$. Then,
  1. (1)
    if $p=\mathfrak{p}$ is inert in $\mathcal{O}_{D}$, then $\mathrm{Frob}_{\mathfrak{p}}(=\mathrm{Frob}_{p}^{2})$ acts on $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ by $p^{k-2}$;
  2. (2)
    if $p$ splits in $\mathcal{O}_{D}$, then for any embedding $\pi:K\hookrightarrow\mathbb{Q}_{p}$, $\mathrm{Frob}_{\pi}(=\mathrm{Frob}_{p})$ acts on $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ by $u_{p}(C)^{a-b}p^{b}$, where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit under the embedding $\pi$. [^4] Technically, $\mathrm{Frob}_{\pi}$ acts on one of $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ and $\psi_{C}^{b}{\hskip 1.0pt}\overline{\psi_{C}^{a}}$ by $u_{p}(C)^{a-b}p^{b}$ depending on the choice of $\psi_{C}$. We will thus replace $\psi_{C}$ by $\overline{\psi_{C}}$ if necessary to ensure that $\mathrm{Frob}_{\pi}$ acts on $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ by $u_{p}(C)^{a-b}p^{b}$.
This suggests that on the modular form side, the space
$\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(\alpha_{D})},\;\frac{E_{k}}{(j-j(\alpha_{D}))^{2}},\ldots,\frac{E_{k}}{(j-j(\alpha_{D}))^{k-1}}\bigg\}$
should also split into $1$-dimensional pieces $\{V_{a,b}\}_{a+b=k-2}$ equipped with corresponding $U_{p}$ actions. In other words, there should exist $\{G_{a,b}\}_{a+b=k-2}$ such that
  1. (1)
    $\mathrm{Span}\{G_{a,b}\mid a+b=k-2\}=\mathrm{Span}\big\{\frac{E_{k}}{j-j(\alpha_{D})},\frac{E_{k}}{(j-j(\alpha_{D}))^{2}},\ldots,\frac{E_{k}}{(j-j(\alpha_{D}))^{k-1}}\big\}$;
  2. (2)
    for all (but finitely many) primes $p$ with $(\frac{D}{p})=-1$, $U_{p}^{2}G_{a,b}\equiv p^{k-2}G_{a,b}$ modulo some power of $p$;
  3. (3)
    for all (but finitely many) primes $p$ with $(\frac{D}{p})=1$, $U_{p}G_{a,b}\equiv u_{p}(C)^{a-b}p^{b}G_{a,b}$ modulo some power of $p$, where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit.
We will now describe the (numerical) construction of these $G_{a,b}$.
Recall that for a (nearly holomorphic) modular form $f(\tau)$ of weight $k_{f}$, we may view $f$ as a function in $\tau$ and $\overline{\tau}$ and define the nonholomorphic derivative as
$(\partial_{\tau}f)(\tau)\;:=\;\frac{1}{2\pi i}\frac{\partial f}{\partial\tau}-\frac{k_{f}}{4\pi\operatorname{Im}(\tau)}\cdot f(\tau).$
For $k\in\{4,6,8,10,14\}$, write  [^5] This two-variable modular form $G_{k}(z,\tau)$ occurred in many places throughout the literature. We refer to [AKN97, p. 98] for a detailed discussion and the references therein. It (and its half-integral weight analogue) also occurred in [Zag02, DJ08a] and plays a key role in Zagier’s duality.
$G_{k}(z,\tau)\;:=\;\frac{E_{k}(z)\cdot\frac{E_{14-k}}{\Delta}(\tau)}{j(z)-j(\tau)},$
and for $1\leq r\leq k-1$, write
$G_{k,D}^{(r)}(z)\;:=\;(\partial_{\tau}^{r-1}G_{k})(z,\alpha_{D}).$
While $G_{k,D}^{(r)}$ a priori depends on the choice of $\alpha_{D}$ and may even be identically zero, we will describe a way to produce a reasonably standard candidate, which we also denote by $G_{k,D}^{(r)}$.
  1. (1)
    If $G_{k,D}^{(r)}$ is identically zero  [^6] In practice, it seems that if $D\neq-3,-4$, then $G_{k,D}^{(r)}$ is never identically zero and in particular, $a_{1}(G_{k,D}^{(r)})\neq 0$., then there should exist some $s\in\mathbb{Z}^{+}$ such that $\frac{(\partial_{\tau}^{r-1}G_{k})(z,\tau)}{(j(\tau)-j(\alpha_{D}))^{s}}$ is not identically zero at $\tau=\alpha_{D}$. We may thus replace $G_{k,D}^{(r)}$ by $\frac{(\partial_{\tau}^{r-1}G_{k})(z,\tau)}{(j(\tau)-j(\alpha_{D}))^{s}}\big|_{\tau=\alpha_{D}}$.
  2. (2)
    By the CM theory, there should exist some nonzero $A=A_{k,D}^{(r)}\in\mathbb{C}$ such that $AG_{k,D}^{(r)}$ has integer coefficients. We may choose $A$ appropriately and replace $G_{k,D}^{(r)}$ by $AG_{k,D}^{(r)}$ so that the coefficients of $G_{k,D}^{(r)}$ have no common divisors and that the first nonzero coefficient of $G_{k,D}^{(r)}$ is positive.
The downside of making a standard choice of $G_{k,D}^{(r)}$ is that every phenomena stated for $G_{k,D}^{(r)}$ in this section may only hold up to a nonzero integer multiple.
Now, each $G_{k,D}^{(r)}$ has a pole at $\alpha_{D}$ of order exactly $r$ and is of the form
$A_{1}\cdot\frac{E_{k}}{j-j(\alpha_{D})}+A_{2}\cdot\frac{E_{k}}{(j-j(\alpha_{D}))^{2}}+\cdots+A_{r}\cdot\frac{E_{k}}{(j-j(\alpha_{D}))^{r}}$
for some $A_{i}\in\mathbb{Z}$. Through the following conjectures, one observes that each $G_{k,D}^{(r)}$ takes the role of $G_{k-1-r,r-1}$. Here the supersingular case is simply a restatement of 3.3 and the ordinary case should be viewed as a natural generalization of 2.4 in the CM case.
Conjecture 4.1.
Let $k\in\{4,6,8,10,14\}$ and let $1\leq r\leq k-1$. Let D<4D<-4 be a discriminant with class number $1$ and let $C/\mathbb{Q}$ be an elliptic curve with $j(C)=j(\alpha_{D})$. Let $p$ be a prime with $(\frac{D}{p})=-1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})\;\equiv\;p^{k-2}a_{np^{l-2}}(G_{k,D}^{(r)})\pmod{p^{(k-1)l-r}}$.
Conjecture 4.2.
Let $k\in\{4,6,8,10,14\}$ and let $1\leq r\leq k-1$. Let D<4D<-4 be a discriminant with class number $1$ and let $C/\mathbb{Q}$ be an elliptic curve with $j(C)=j(\alpha_{D})$. Let $p$ be a prime with $(\frac{D}{p})=1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})\;\equiv\;u_{p}(C)^{k-2r}p^{r-1}a_{np^{l-1}}(G_{k,D}^{(r)})\pmod{p^{(k-1)l}}$,
where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit.
Remark 4.3.
As with the case of 3.3, it is also possible to deduce 3.1 from 4.2 in the ordinary and CM case through involved computations which we omit here.
As mentioned in the introduction, the magnetic property shows up in all the CM cases. Specifically, we have the following conjecture.
Conjecture 4.4.
Let $k\in\{4,6,8,10,14\}$, $1\leq r\leq k-1$, and $r^{\prime}=\min(r,k-r)$. Let D<4D<-4 be a discriminant with class number $1$. Then, $G_{k,D}^{(r)}$ is $(r^{\prime}-1)$-magnetic, i.e., for all $n\in\mathbb{Z}^{+}$,
$n^{r^{\prime}-1}{\hskip 1.0pt}\big|{\hskip 1.0pt}a_{n}(G_{k,D}^{(r)}).$
Remark 4.5.
If one only focuses on $a_{p}=a_{p}(G_{k,D}^{(r)})$, then generically one obtains the following table from 4.1 and 4.2.
$r$ 11 22 33 $\cdots$ k2k-2 k1k-1
$v_{p}(a_{p})$ for $(\frac{D}{p})=-1$ k2k-2 k3k-3 k4k-4 $\cdots$ 11 0
$v_{p}(a_{p})$ for $(\frac{D}{p})=1$ 0 11 22 $\cdots$ k3k-3 k2k-2
To some extent, this also explains the $(r^{\prime}-1)$-magnetic property of $G_{k,D}^{(r)}$.
Before proceeding further with numerical phenomena, we present an example for $k=4$ and D=7D=-7, which may help with understanding the setting.
Example 4.6.
Let $k=4$ and D=7D=-7. Explicitly, we may choose $\alpha_{D}=\frac{1+\sqrt{-7}}{2}$ and $C/\mathbb{Q}$ given by $y^{2}+xy=x^{3}-x^{2}-2x-1$ with LMFDB label 49.a4 [LMF25, Elliptic Curve 49.a4]. In this case, $j(C)=j(\alpha_{D})=-3375$. Then,
$G_{4,-7}^{(1)}$ $\;=\;\frac{E_{4}}{j+3375}$
$G_{4,-7}^{(2)}$ = $\;=\;19\cdot\frac{E_{4}}{j+3375}-91125\cdot\frac{E_{4}}{(j+3375)^{2}}$
$G_{4,-7}^{(3)}$ = $\;=\;1399\cdot\frac{E_{4}}{j+3375}-19008675\cdot\frac{E_{4}}{(j+3375)^{2}}+54251268750\cdot\frac{E_{4}}{(j+3375)^{3}}.$
For all primes $p$ with $(\frac{-7}{p})=1$ and all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{4,-7}^{(1)})$ $\;\equiv\;u_{p}(C)^{2}{\hskip 1.0pt}a_{np^{l-1}}(G_{4,-7}^{(1)})$ $\pmod{p^{3l}}$
$a_{np^{l}}(G_{4,-7}^{(2)})$ $\;\equiv\;p{\hskip 1.0pt}a_{np^{l-1}}(G_{4,-7}^{(2)})$ $\pmod{p^{3l}}$
$a_{np^{l}}(G_{4,-7}^{(3)})$ $\;\equiv\;u_{p}(C)^{-2}p^{2}{\hskip 1.0pt}a_{np^{l-1}}(G_{4,-7}^{(3)})$ $\pmod{p^{3l}}$,
where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit. Moreover, $G_{4,-7}^{(2)}$ is $1$-magnetic.
In fact, one can already observe interesting numerical phenomena for the family $\{G_{k,D}^{(r)}\}_{1\leq r\leq k-1}$ by viewing $\mathrm{Sym}^{k-2}C$ simply as a Galois representation of $\mathbb{Q}$. Indeed, we have the following splitting (as Galois representations of $\mathbb{Q}$)  [^7] An easy way of observing this splitting is through the splitting of the $L$-function of $\mathrm{Sym}^{k-2}C$.
$\mathrm{Sym}^{k-2}C\;\cong\;\big((\tfrac{D}{\cdot}){\hskip 1.0pt}\varepsilon_{\ell}\big)^{\frac{k-2}{2}}\oplus{\hskip 1.0pt}\bigoplus_{r=1}^{\frac{k-2}{2}}\big(\varepsilon_{\ell}^{r-1}\otimes\mathrm{Ind}_{K/\mathbb{Q}}\psi_{C}^{k-2r}\big),$
where $(\frac{D}{\cdot})$ denotes (the Dirichlet character given by) the Kronecker symbol and $\varepsilon_{\ell}$ denotes the $\ell$-adic cyclotomic character. Transferring this splitting to the modular form side, one also obtains certain congruences for $G_{k,D}^{(k/2)}$ and the others. In this case, it is more convenient to write everything in terms of negative discriminants.
Theorem 4.7.
Let $k\in\{4,6,8,10,14\}$ and let D<4D<-4 be a discriminant with class number $1$. Then, $G_{k,D}^{(k/2)}$ satisfies that for all primes $p$ with $p\nmid D$ and all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(k/2)})\;\equiv\;\big(\big(\tfrac{D}{p}\big){\hskip 1.0pt}p\big)^{\frac{k-2}{2}}a_{np^{l-1}}(G_{k,D}^{(k/2)})\pmod{p^{(k-1)l}}$
Remark 4.8.
We will prove this in Section 6. Note that there are no assumptions on $v_{p}(j(C))$ and $v_{p}(j(C)-1728)$ so primes dividing $D$ are the only “bad primes”. In fact, we will also prove some $p$-adic property for $p{\hskip 1.0pt}|{\hskip 1.0pt}D$, from which the $\frac{k-2}{2}$-magnetic property of (some multiple of) $G_{k,D}^{(k/2)}$ as predicted by 4.4 then follows.
Conjecture 4.9.
Let $k\in\{4,6,8,10,14\}$, $1\leq r\leq k-1$ with $r\neq\frac{k}{2}$, and $r^{\prime}=\min(r,k-r)$. Let D<4D<-4 be a discriminant with class number $1$. Consider the CM modular form (in fact Hecke eigenform)
$\Theta_{r^{\prime}}\;:=\;\Theta_{k+1-2r^{\prime},D}\;:=\;\frac{1}{2}\sum_{\alpha\in\mathcal{O}_{D}}\alpha^{k-2r^{\prime}}q^{|\mathrm{N}(\alpha)|}$
of weight $k+1-2r^{\prime}$. Let $p$ be a prime with $p\nmid D$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})-p^{r^{\prime}-1}a_{p}(\Theta_{r^{\prime}})a_{np^{l-1}}(G_{k,D}^{(r)})+\big(\tfrac{D}{p}\big)p^{k-2}a_{np^{l-2}}(G_{k,D}^{(r)})\;\equiv\;0\pmod{p^{(k-1)l-r}}$.
Remark 4.10.
This should be viewed as a natural generalization of the ASD congruence observed in the case of CM elliptic curves and simple poles, i.e., 2.6 or 2.8. As in Remark 2.7, 4.9 also follows from the supersingular (4.1) and ordinary (4.2) dichotomy.
The final observations we have on the family $\{G_{k,D}^{(r)}\}_{1\leq r\leq k-1}$ remain somewhat mysterious to us. In the splitting (7) of the elliptic curve side over $K$, $\psi_{C}^{a}\overline{\psi_{C}^{b}}$ and $\psi_{C}^{b}\overline{\psi_{C}^{a}}$ multiply to $(\psi_{C}\overline{\psi_{C}})^{k-2}=(\mathrm{Res}_{K/\mathbb{Q}}\varepsilon_{\ell})^{k-2}$. It seems that this is also reflected on the modular form side via the following phenomena.
Conjecture 4.11.
Let $k\in\{4,6,8,10,14\}$, $1\leq r\leq k-1$, and $r^{\prime}=\min(r,k-r)$. Let D<4D<-4 be a discriminant with class number $1$. Let $p$ be a prime with $p\nmid D$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})a_{np^{l}}(G_{k,D}^{(k-r)})\;\equiv\;p^{k-2}a_{np^{l-1}}(G_{k,D}^{(r)})a_{np^{l-1}}(G_{k,D}^{(k-r)})\pmod{p^{(k+r^{\prime}-2)l}}$.
Remark 4.12.
It is easy to check that the case when $r=k/2$ follows from Theorem 4.7 and the case when $(\frac{D}{p})=1$ follows from 4.2. When $(\frac{D}{p})=-1$, the numerical observations also seem to suggest that for all $n,l\in\mathbb{Z}^{+}$ with $l\geq 2$,
$a_{np^{l}}(G_{k,D}^{(r)})a_{np^{l}}(G_{k,D}^{(k-r)})\;\equiv\;p^{k-2}a_{np^{l-1}}(G_{k,D}^{(r)})a_{np^{l-1}}(G_{k,D}^{(k-r)})\pmod{p^{(k+r^{\prime}-2)l+(k-2r^{\prime})}}$.
Conjecture 4.13.
Let $k\in\{4,6,8,10,14\}$ and $1\leq r\leq k-1$. Let D<4D<-4 be a discriminant with class number $1$. Then, for all $n\in\mathbb{Z}^{+}$,
$n^{k-2}{\hskip 1.0pt}\big|{\hskip 1.0pt}a_{n}(G_{k,D}^{(r)})a_{n}(G_{k,D}^{(k-r)}).$
Remark 4.14.
Note that a priori 4.4 only implies that $n^{2r^{\prime}}{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}(G_{k,D}^{(r)})a_{n}(G_{k,D}^{(k-r)})$, where $r^{\prime}=\min(r,k-r)$. On the other hand, the case when $(n,D)=1$ should follow from 4.11.
5. Proof of 2.1
We will now prove 2.1 when $p\geq 5$, or equivalently, $p\nmid 6$. As we will work with general number fields, we first introduce some notations.
Let $L$ be a number field. For a prime $\mathfrak{p}$ of $L$, let $\mathbb{F}_{\mathfrak{p}}$ denote its residue field and let $\mathrm{N}(\mathfrak{p}):=|\mathbb{F}_{\mathfrak{p}}|$ denote its norm. For an elliptic curve $C/L$ and a good prime $\mathfrak{p}$ of $C$, define
$a_{\mathfrak{p}}\;:=\;\mathrm{N}(\mathfrak{p})+1-|C(\mathbb{F}_{\mathfrak{p}})|.$
Now, for $k\in\{4,6,8,10,14\}$ and an elliptic curve $C/L$, write
$F_{k,C}\;:=\;\frac{E_{k}}{j-j(C)}.$
For the convenience of the discussion in this section, we also write
$F_{k,c}\;:=\;\frac{E_{k}}{j-c},$
for $k\in\{4,6,8,10,14\}$ and $c\in\overline{\mathbb{Q}}$ or $\overline{\mathbb{Z}_{p}}$. The main theorem that we will prove in this section is the following.
Theorem 5.1.
Let $k\in\{4,6,8,10,14\}$, $L$ be a number field, and $C/L$ be an elliptic curve. Let $\mathfrak{p}$ be a good prime of $C$ with $\mathfrak{p}\nmid 6$ and $v_{\mathfrak{p}}(j(C))=0=v_{\mathfrak{p}}(j(C)-1728)$. Then, for all $n\in\mathbb{Z}^{+}$,a
$a_{n\cdot\mathrm{N}(\mathfrak{p})}(F_{k,C})\;\equiv\;a_{\mathfrak{p}}(C)^{k-2}a_{n}(F_{k,C})\pmod{\mathfrak{p}}$.
As mentioned in Remark 2.2, the proof depends on the hypergeometric interpretation of the coefficients in the case when $k=4$. In particular, we will show the following result.
Theorem 5.2.
Let $p$ be a rational prime with $p\nmid 6$ and let $c\in\overline{\mathbb{Z}_{p}}$ with $v_{p}(c)=0$. Then, for all $l\in\mathbb{Z}^{+}$,
$a_{p^{l}}(F_{4,c})$ $\;\equiv\;\big(c(c-1728)\big)^{\frac{p^{l}-1}{2}}\cdot\sum_{m=0}^{p^{l}-1}\frac{(6m)!}{(m!)^{3}(3m)!}{\hskip 1.0pt}c^{-m}\pmod{p}$.
We start by first recalling the definitions of the Hecke operators (on $q$-series). For $k\in\mathbb{Z}$, a rational prime $p$, and a $q$-series $f=\sum_{n\gg-\infty}a_{n}(f)q^{n}$, define
$f|U_{p}\;:=\;\sum_{n\gg-\infty}a_{np}(f)q^{n},\quad f|V_{p}\;:=\;\sum_{n\gg-\infty}a_{n}(f)q^{np},\quad f|T_{p,k}\;:=\;f|(U_{p}+p^{k-1}V_{p}).$
More generally, for a $q$-series $f=\sum_{n\gg-\infty}a_{n}(f)q^{n}$, $k\in\mathbb{Z}$, and $m\in\mathbb{Z}^{+}$, define $f|T_{m,k}$ by setting
$a_{n}(f|T_{m,k})\;:=\;\sum_{r>0,r|(m,n)}r^{k-1}a_{mn/r^{2}}(f).$
It is easy to check that for $m,n\in\mathbb{Z}^{+}$ and $k\in\mathbb{Z}$,
$T_{m,k}{\hskip 1.0pt}T_{n,k}\;=\;\sum_{r>0,r|(m,n)}r^{k-1}T_{mn/r^{2},k}.$
5.1. A first reinterpretation
For each $k,n$, there exists a polynomial $P_{k,n}\in\mathbb{Z}[X]$ of degree n1n-1 such that
$\frac{E_{k}}{j-X}\;=\;\sum_{n=1}^{\infty}P_{k,n}(X)q^{n}.$
Then,
(8) $a_{n}(F_{k,c})\;=\;P_{k,n}(c).$
Lemma 5.3.
For $k\in\{4,6,8,10,14\}$,
(9) $\frac{n^{k-1}(E_{14-k}/\Delta)|T_{n,2-k}}{E_{14-k}/\Delta}\;=\;P_{k,n}(j),$
where $E_{0}:=1$.
Proof.
By [AKN97, Theorem 3] or [DJ08a, Theorem 2], it is known that
$G_{k}(z,\tau)\;=\;\frac{E_{k}(z)\cdot(E_{14-k}/\Delta)(\tau)}{j(z)-j(\tau)}\;=\;\sum_{n=1}^{\infty}g_{2-k,n}(\tau)e^{2\pi inz}$
with $g_{2-k,n}=n^{k-1}(E_{14-k}/\Delta)|T_{n,2-k}$. The lemma then follows from dividing both sides of the expansion by $(E_{14-k}/\Delta)(\tau)$. ∎
Lemma 5.4.
Let $k\in 2\mathbb{Z}^{+}$, $p$ be a rational prime, and $g\;=\;\sum_{m\gg-\infty}a_{m}(g)q^{m}$ with $a_{m}(g)\in\overline{\mathbb{Z}_{p}}$. For each $n\in\mathbb{Z}^{+}$, let
$g_{n}\;:=\;n^{k-1}g|T_{n,2-k}.$
Then,
$g_{np}\;\equiv\;g_{n}|V_{p}\pmod{p^{k-1}}$
as $q$-series.
Proof.
Since $T_{np,2-k}=T_{n,2-k}T_{p,2-k}-p^{1-k}T_{n/p,2-k}$, where $T_{n/p,2-k}:=0$ if $p\nmid n$, it follows that
$g_{np}\;=\;p^{k-1}g_{n}|T_{p,2-k}-p^{k-1}g_{n/p}\;=\;g_{n}|V_{p}+p^{k-1}g_{n}|U_{p}-p^{k-1}g_{n/p}\;\equiv\;g_{n}|V_{p}\pmod{p^{k-1}}$,
where $g_{n/p}:=0$ if $p\nmid n$. The result then follows. ∎
Corollary 5.5.
Let $k\in\{4,6,8,10,14\}$ and $p$ be a rational prime. Let $g_{2-k}=E_{14-k}/\Delta$ and let $g_{2-k,n}=n^{k-1}g_{2-k}|T_{n,2-k}$. Then,
$P_{k,np}(j)\;\equiv\;\frac{g_{2-k,n}^{p}}{g_{2-k}}\pmod{p}$
as $q$-series. In particular, for $l\in\mathbb{Z}^{+}$,
(10) $P_{k,p^{l}}(j)\;\equiv\;g_{2-k}^{p^{l}-1}\pmod{p}$
as $q$-series.
Proof.
Indeed,
$P_{k,np}(j)\;=\;\frac{g_{2-k,np}}{g_{2-k}}\;\equiv\;\frac{g_{2-k,n}|V_{p}}{g_{2-k}}\;\equiv\;\frac{g_{2-k,n}^{p}}{g_{2-k}}\pmod{p}$,
where we also use the fact that $g_{2-k,n}$ has integer coefficients. The second congruence now follows from induction. ∎
Let $n\in\mathbb{Z}^{+}$ and $p$ be a rational prime. Then, by Corollary 5.5
$\frac{P_{k,np}(j)}{P_{k,p}(j)P_{k,n}(j)}\;\equiv\;\frac{g_{2-k,n}^{p}/g_{2-k}}{(g_{2-k}^{p}/g_{2-k})\cdot(g_{2-k,n}/g_{2-k})}\;\equiv\;\bigg(\frac{g_{2-k,n}}{g_{2-k}}\bigg)^{p-1}\;\equiv\;P_{k,n}(j)^{p-1}\pmod{p}$
as $q$-series. Hence,
$P_{k,np}(j)\;\equiv\;P_{k,p}(j)\cdot P_{k,n}(j)^{p}\pmod{p}$
as polynomials in $j$. Plugging in $j=c\in\overline{\mathbb{Z}_{p}}$ and using that $a_{n}(F_{k,c})=P_{k,n}(c)$, we thus obtain that
$a_{np}(F_{k,c})\;\equiv\;a_{p}(F_{k,c})a_{n}(F_{k,c})^{p}\pmod{p}$.
For $l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,c})\;\equiv\;a_{p}(F_{k,c}){\hskip 1.0pt}a_{np^{l-1}}(F_{k,c})^{p}\;\equiv\;\cdots\;\equiv\;a_{p}(F_{k,c})^{\frac{p^{l}-1}{p-1}}{\hskip 1.0pt}a_{n}(F_{k,c})^{p^{l}}\pmod{p}$.
Now, suppose that $L$ is a number field and $c\in L$. Then, for any prime $\mathfrak{p}$ of $L$ with $v_{\mathfrak{p}}(c)\geq 0$,
$a_{n\cdot\mathrm{N}(\mathfrak{p})}(F_{k,c})\;\equiv\;a_{p}(F_{k,c})^{\frac{\mathrm{N}(\mathfrak{p})-1}{p-1}}a_{n}(F_{k,c})\pmod{\mathfrak{p}}$.
Since $a_{1}(F_{k,c})=1$, this also implies that
$a_{\mathrm{N}(\mathfrak{p})}(F_{k,c})\;\equiv\;a_{p}(F_{k,c})^{\frac{\mathrm{N}(\mathfrak{p})-1}{p-1}}\pmod{\mathfrak{p}}$.
In summary, we obtain the following propositions.
Proposition 5.6.
Let $k\in\{4,6,8,10,14\}$, $p$ be a rational prime, and $c\in\overline{\mathbb{Z}_{p}}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,c})\;\equiv\;a_{p}(F_{k,c})^{\frac{p^{l}-1}{p-1}}{\hskip 1.0pt}a_{n}(F_{k,c})^{p^{l}}\pmod{p}$.
Proposition 5.7.
Let $k\in\{4,6,8,10,14\}$, $L$ be a number field, $c\in L$, and $\mathfrak{p}$ be a prime of $L$ with $v_{\mathfrak{p}}(c)\geq 0$. Then, for all $n\in\mathbb{Z}^{+}$,
$a_{n\cdot\mathrm{N}(\mathfrak{p})}(F_{k,c})\;\equiv\;a_{p}(F_{k,c})^{\frac{\mathrm{N}(\mathfrak{p})-1}{p-1}}a_{n}(F_{k,c})\;\equiv\;a_{\mathrm{N}(\mathfrak{p})}(F_{k,c}){\hskip 1.0pt}a_{n}(F_{k,c})\pmod{\mathfrak{p}}$.
From this, to prove Theorem 5.1, it suffices to show that
$a_{\mathrm{N}(\mathfrak{p})}(F_{k,C})\;\equiv\;a_{\mathfrak{p}}(C)^{k-2}\pmod{\mathfrak{p}}$.
5.2. Reducing to $k=4$
We first prove that all the cases when $k\in\{4,6,8,10,14\}$ in Theorem 5.1 follow from the single case when $k=4$.
Let $p$ be a rational prime and let $l\in\mathbb{Z}^{+}$. By Corollary 5.5,
$P_{k,p^{l}}(j)\;\equiv\;(E_{14-k}/\Delta)^{p^{l}-1}\pmod{p}$
as $q$-series. For $k\in\{6,8,10,14\}$, note that
[…]
Hence, for $k\in\{6,8,10,14\}$,
$P_{4,p^{l}}(j)^{\frac{k-2}{2}}\;\equiv\;\big((E_{10}/\Delta)^{\frac{k-2}{2}}\big)^{p^{l}-1}$ $\;=\;\big(j^{a}(j-1728)^{b}\big)^{p^{l}-1}(E_{14-k}/\Delta)^{p^{l}-1}$
$\;\equiv\;\big(j^{a}(j-1728)^{b}\big)^{p^{l}-1}P_{k,p^{l}}(j)\pmod{p}$,
as $q$-series, for some $a,b\in\mathbb{Z}_{\geq 0}$. This implies that
(11) $P_{4,p^{l}}(j)^{\frac{k-2}{2}}\;\equiv\;\big(j^{a}(j-1728)^{b}\big)^{p^{l}-1}P_{k,p^{l}}(j)\pmod{p}$
as polynomials in $j$. We thus obtain the following result.
Lemma 5.8.
Let $L$ be a number field and $C/L$ be an elliptic curve. Let $\mathfrak{p}$ be a prime of $L$ with $v_{\mathfrak{p}}(j(C))=0=v_{\mathfrak{p}}(j(C)-1728)$. If $a_{\mathrm{N}(\mathfrak{p})}(F_{4,C})\equiv a_{\mathfrak{p}}(C)^{2}\pmod{\mathfrak{p}}$, then $a_{\mathrm{N}(\mathfrak{p})}(F_{k,C})\equiv a_{\mathfrak{p}}(C)^{k-2}\pmod{\mathfrak{p}}$ for all $k\in\{4,6,8,10,14\}$.
Proof.
Let $p$ be the rational prime lying below $\mathfrak{p}$ and write $\mathrm{N}(\mathfrak{p})=p^{l}$. For $k\in\{6,8,10,14\}$, by Equations (8) and (11), there exist $a,b\in\mathbb{Z}_{\geq 0}$ such that
(12) $a_{p^{l}}(F_{4,C})^{\frac{k-2}{2}}\;\equiv\;\big(j(C)^{a}(j(C)-1728)^{b}\big)^{p^{l}-1}a_{p^{l}}(F_{k,C})\;\equiv\;a_{p^{l}}(F_{k,C})\pmod{\mathfrak{p}}$.
The result then follows. ∎
5.3. Hypergeometric reinterpretation of $a_{p^{l}}(F_{4,c})$
Now, we specialize to $a_{p^{l}}(F_{4,c})$ and prove Theorem 5.2.
We first recall the definition of hypergeometric functions. Given a hypergeometric datum $(\boldsymbol{\alpha},\boldsymbol{\beta})$ with
$\boldsymbol{\alpha}$ $\;=\;(\alpha_{1},\ldots,\alpha_{n-1},\alpha_{n})\in\mathbb{C}^{n}$
$\boldsymbol{\beta}$ $\;=\;(\beta_{1},\ldots,\beta_{n-1})\in\mathbb{C}^{n-1},$
the associated hypergeometric function is defined as
[…]
and for $r\in\mathbb{Z}_{\geq 0}$, the truncated hypergeometric sum is defined as
[…]
where $(a)_{m}:=\prod_{i=0}^{m-1}(a+i)$.
Let $p\nmid 6$ be a rational prime and let $l\in\mathbb{Z}^{+}$. Recall the following identity by Fricke and Klein [Fri12, Chapter 5, Section 10]
[…]
and Clausen’s formula
[…]
It follows that
[…]
Now, write
$Q_{n}(X)\;:=\;X^{n-1}P_{4,n}(X^{-1}).$
Then,
[…]
as $q$-series. Let $\alpha=j^{-1}=q+O(q^{2})$. Then,
[…]
$Q_{p^{l}}(\alpha)\;\equiv\;(1-1728\alpha)^{\frac{p^{l}-1}{2}}B(\alpha)^{1-p^{l}}\;\equiv\;(1-1728\alpha)^{\frac{p^{l}-1}{2}}B(\alpha)\cdot(B^{-1})(\alpha^{p^{l}})\pmod{p}$,
where $B^{-1}$ is the inverse of $B$ as a power series. Since $\deg Q_{p^{l}}=\deg P_{4,p^{l}}\leq p^{l}-1$ and $B(0)=1$, this implies that
[…]
Here for a power series $R(z)=\sum_{n=0}^{\infty}a_{n}z^{n}$, we write
$\mathrm{Trun}_{m}(R)(z)\;:=\;\sum_{n=0}^{m}a_{n}z^{n}\quad\text{ and }\quad\mathrm{Trun}_{m,z}\bigg(\sum_{n=0}^{\infty}a_{n}z^{n}\bigg)\;:=\;\sum_{n=0}^{m}a_{n}z^{n}.$
For $p\nmid 6$, it is easy to check that $\frac{(\frac{1}{2})_{r}(\frac{1}{6})_{r}(\frac{5}{6})_{r}}{(r!)^{3}}$ is divisible by $p$ for $\frac{p^{l}-1}{6}<r\leq p^{l}-1$. It follows that
$\mathrm{Trun}_{p^{l}-1,\alpha}\big((1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot B(\alpha)\big)\;=\;$ $\mathrm{Trun}_{p^{l}-1,\alpha}\big((1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot\mathrm{Trun}_{p^{l}-1}(B)(\alpha)\big)$
$\;\equiv\;$ $\mathrm{Trun}_{p^{l}-1,\alpha}\big((1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot\mathrm{Trun}_{\lfloor(p^{l}-1)/6\rfloor}(B)(\alpha)\big)$
$\;\equiv\;$ $(1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot\mathrm{Trun}_{\lfloor(p^{l}-1)/6\rfloor}(B)(\alpha)$
$\;\equiv\;$ $(1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot\mathrm{Trun}_{p^{l}-1}(B)(\alpha)\pmod{p}$.
To summarize, we obtain that
[…]
as polynomials in $\alpha$.
Now, for $c\in\overline{\mathbb{Z}_{p}}$ with $v_{p}(c)=0$,
$a_{p^{l}}(F_{4,c})\;=\;P_{4,p^{l}}(c)$ $\;=\;c^{p^{l}-1}Q_{p^{l}}(c^{-1})$
[…]
$\;\equiv\;\big(c(c-1728)\big)^{\frac{p^{l}-1}{2}}\cdot\sum_{n=0}^{p^{l}-1}\frac{(6m)!}{(m!)^{3}(3m)!}{\hskip 1.0pt}c^{-m}\pmod{p}$,
which proves Theorem 5.2.
5.4. Proof of Theorem 5.1
Let $L$ be a number field, $C/L$ be an elliptic curve, and $\mathfrak{p}$ be a prime of $L$ with $v_{\mathfrak{p}}(j(C))=0=v_{\mathfrak{p}}(j(C)-1728)$. Let $\mathbb{F}_{\mathfrak{p}}$ denote the residue field of $\mathfrak{p}$, let $p$ be the rational prime lying under $\mathfrak{p}$, and write $\mathrm{N}(\mathfrak{p})=p^{l}$. For $a\in L$ with $v_{\mathfrak{p}}(a)\geq 0$, define a quadratic character
[…]
By Theorem 5.2,
[…]
It thus suffices to show that
[…]
This congruence should follow from the hypergeometric machinery, though an explicit reference seems hard to find. As the proof is solely hypergeometric and unrelated to the main discussion of this paper, we will only write a sketch here.  [^8] The author plans to write a separate and self-contained note on the proof of this congruence in details.
Write $c=j(C)$ and consider the elliptic curve
$C_{1}:y^{2}+xy\;=\;x^{3}-\frac{\lambda}{432}$
with $\lambda=\frac{1-\sqrt{1-1728/c}}{2}$ (so that $j(C_{1})=c$). It is known that for a prime $\mathfrak{P}$ of $L(\sqrt{1-1728/c})$ that is a good prime of $C_{1}$,
[…]
Now, one can relate $a_{\mathfrak{p}}(C)$ and $a_{\mathfrak{P}}(C_{1})$ via the fact that $j(C)=j(C_{1})$ and use an appropriate version of the truncated Clausen’s formula to obtain the desired congruence (see e.g. [CDL+13, Lemma 18] for the truncated Clausen’s formula over $\mathbb{Q}$).
6. Proof of Theorem 4.7 and the $\frac{k-2}{2}$-magnetic property
We will now prove Theorem 4.7, i.e. the supercongruences satisfied by $G_{k,D}^{(k/2)}$. With extra work for primes $p$ with $p{\hskip 1.0pt}|{\hskip 1.0pt}D$, it also follows that, up to some integer multiple, $G_{k,D}^{(k/2)}$ is $\frac{k-2}{2}$-magnetic. In fact, we will prove a more general result on supercongruences and the $\frac{k-2}{2}$-magnetic property which holds for any discriminant $D<0$.
Recall that for $k\in\{4,6,8,10,14\}$,
$G_{k}(z,\tau)\;:=\;\frac{E_{k}(z)\cdot\frac{E_{14-k}}{\Delta}(\tau)}{j(z)-j(\tau)}.$
Let $D<0$ be a discriminant. Let $\mathcal{Q}_{D}$ be a complete set of $\mathrm{SL}_{2}(\mathbb{Z})$-inequivalent positive definite quadratic forms $Q(X,Y)\in\mathbb{Z}[X,Y]$ of discriminant $D$ and let $\mathcal{Q}_{D}^{\mathrm{prim}}\subseteq\mathcal{Q}_{D}$ be the subset consisting of primitive elements, i.e., $aX^{2}+bXY+cY^{2}$ with $(a,b,c)=1$. For $Q(X,Y)=aX^{2}+bXY+cY^{2}\in\mathcal{Q}_{D}$, define
$\alpha_{Q}\;:=\;\frac{-b+\sqrt{D}}{2a}\in\mathcal{H}.$
Now, define
$\mathcal{G}_{k,D}(z)\;:=\;\frac{2}{w_{D}}\sum_{Q\in\mathcal{Q}_{D}^{\mathrm{prim}}}(\partial^{(k-2)/2}_{\tau}G_{k})(z,\alpha_{Q}),$
where
$w_{D}\;:=\;\begin{cases}4&\text{ if $D=-4$,}\\ 6&\text{ if $D=-3$,}\\ 2&\text{ otherwise.}\end{cases}$
Note that $(\partial^{(k-2)/2}_{\tau}G_{k})(z,\tau)$ is of weight 0 with respect to $\tau$, so $\mathcal{G}_{k,D}(z)$ is independent of the choice of $\mathcal{Q}_{D}$ (unlike $G_{k,D}^{(r)}$).
Theorem 6.1.
Let $k\in\{4,6,8,10,14\}$ and let $D<0$ be a discriminant. Then, $\mathcal{G}_{k,D}$ satisfies the following.
  1. (1)
    $\mathcal{G}_{k,D}$ has a Fourier series expansion of the form
    $\mathcal{G}_{k,D}(z)\;=\;\sum_{n=1}^{\infty}a_{n}(\mathcal{G}_{k,D})q^{n}\in\mathbb{Q}[[q]].$
  2. (2)
    Write $D=A^{2}D_{0}$ for some fundamental discriminant $D_{0}<0$ and let
[…]
    Then, $\widetilde{\mathcal{G}}_{k,D}$ is $\frac{k-2}{2}$-magnetic, i.e., $\widetilde{\mathcal{G}}_{k,D}\in\mathbb{Z}[[q]]$ and for all $n\in\mathbb{Z}^{+}$,
    $n^{\frac{k-2}{2}}{\hskip 1.0pt}\big|{\hskip 1.0pt}a_{n}(\widetilde{\mathcal{G}}_{k,D}).$
  3. (3)
    Let $p$ be a prime with $p\nmid A$. Then, for all $n,l\in\mathbb{Z}^{+}$,
    $a_{np^{l}}(\widetilde{\mathcal{G}}_{k,D})\;\equiv\;\big(\big(\tfrac{D}{p}\big)p\big)^{\frac{k-2}{2}}a_{np^{l-1}}(\widetilde{\mathcal{G}}_{k,D})\pmod{p^{(k-1)l}}$.
Theorem 6.1 should be viewed as a natural generalization of Li–Neururer [LN19, Theorem 1.5] and Paşol–Zudilin [PZ22, Theorem 1 and 2]. Indeed, the proof uses the same method: one considers the half-integral weight preimage of $\mathcal{G}_{k,D}$ under the Shimura lift, analyzes the $U_{p}$ action on the preimage, and transfers the properties of the preimage to the supercongruences and the magnetic property of $\mathcal{G}_{k,D}$.
6.1. The Shimura lift
We first recall some definitions and notations regarding the Shimura lift. For $k,s\in\mathbb{Z}$, let $M_{k}^{!}$ denote the space of weakly holomorphic modular forms of weight $k$ and level $1$ and let $M_{s+1/2}^{!,+}$ denote the Kohnen plus space, i.e., the space of weakly holomorphic modular forms of weight $s+1/2$ and level $\Gamma_{0}(4)$ whose $n$-th Fourier coefficient is zero if $(-1)^{s}n\not\equiv 0,1\pmod{4}$. For $f\in M_{s+1/2}^{!,+}$ and a fundamental discriminant $d_{0}$ with $(-1)^{s}d_{0}>0$, define the $d_{0}$-th Shimura lift of $f$ by
$\mathcal{S}_{d_{0}}(f)\;:=\;\frac{1}{2}\cdot L\big(1-s,(\tfrac{d_{0}}{\cdot})\big){\hskip 1.0pt}a_{0}(f)+\sum_{n=1}^{\infty}\bigg(\sum_{m{\hskip 1.0pt}|{\hskip 1.0pt}n}\big(\tfrac{d_{0}}{m}\big){\hskip 1.0pt}m^{s-1}{\hskip 1.0pt}a_{|d_{0}|n^{2}/m^{2}}(f)\bigg)q^{n}.$
Let $d$ be a discriminant and $d_{0}$ be a fundamental discriminant with $dd_{0}<0$. Let $\mathcal{Q}_{dd_{0}}$ be a complete set of $\mathrm{SL}_{2}(\mathbb{Z})$-inequivalent positive definite quadratic forms $Q(X,Y)\in\mathbb{Z}[X,Y]$ of discriminant $dd_{0}$. For each $Q(X,Y)=aX^{2}+bXY+cY^{2}\in\mathcal{Q}_{dd_{0}}$, define
$\alpha_{Q}\;:=\;\frac{-b+\sqrt{dd_{0}}}{2a}\in\mathcal{H},$
define
[…]
and let
[…]
For an $\mathrm{SL}_{2}(\mathbb{Z})$-invariant function $f(\tau)$ on $\mathcal{H}$, recall the following definition from [DJ08b]
(14) $\mathrm{Tr}_{d,d_{0}}(f)\;=\;\sum_{Q\in\mathcal{Q}_{dd_{0}}}\frac{1}{w_{Q}}\cdot\chi(Q)f(\alpha_{Q}).$
Proposition 6.2.
Let $2s\in\{4,6,8,10,14\}$. Let $d$ be a discriminant and $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$ and $dd_{0}<0$. Then,
$\big(\mathcal{S}_{d_{0}}f_{s+1/2,|d|}\big)(z)\;=\;-(-1)^{\lfloor\frac{s-1}{2}\rfloor}{\hskip 1.0pt}|d|^{-\frac{s}{2}}{\hskip 1.0pt}|d_{0}|^{\frac{s-1}{2}}\;\mathrm{Tr}_{d,d_{0}}\big((\partial^{s-1}_{\tau}G_{2s})(z,\tau)\big),$
where $\mathrm{Tr}_{d,d_{0}}$ is taken with respect to $\tau$.
Proof.
This is a special case of [DJ08b, Proposition 10]. ∎
6.2. The $U_{p}$ action
For a $q$-series $f=\sum_{n\gg-\infty}a_{n}(f)q^{n}$, we adopt the notations $f|U_{p},f|V_{p}$, and $f|T_{p,k}$ for from Section 5. For a map $\chi:\mathbb{Z}\rightarrow\mathbb{C}$ (in our case a Dirichlet character), define
$f|\chi\;:=\;\sum_{n\gg-\infty}\chi(n)a_{n}(f)q^{n}.$
Now, let $s\in\mathbb{Z}$ and let
$f\;=\;\sum_{(-1)^{s}n\equiv 0,1\;(\mathrm{mod}\;4)}a_{n}(f)q^{n}\in M_{s+1/2}^{+,!}.$
For a prime $p$, define
$f|T_{p,s+1/2}\;:=\;\sum_{(-1)^{s}n\equiv 0,1\;(\mathrm{mod}\;4)}\big(a_{np^{2}}(f)+p^{s-1}\big(\tfrac{(-1)^{s}n}{p}\big)a_{n}(f)+p^{2s-1}a_{n/p^{2}}(f)\big)q^{n}.$
Indeed, if $p$ is odd, then
$f|T_{p,s+1/2}\;=\;f|(U_{p}^{2}+p^{s-1}\chi_{p}^{\prime}+p^{2s-1}V_{p}^{2}),$
where $\chi_{p}^{\prime}(n)=\big(\frac{(-1)^{s}n}{p}\big)$.
Now, let $2s\in\{4,6,8,10,14\}$. For every $m\in\mathbb{Z}$ with $m\geq 0$ and $(-1)^{s-1}m\equiv 0,1\pmod{4}$, there exists a unique $f_{s+1/2,m}\in M_{s+1/2}^{!,+}$ such that
$f_{s+1/2,m}\;=\;q^{-m}+O(q)\in\mathbb{Z}[[q,q^{-1}]].$
In particular, these $f_{s+1/2,m}$ form a basis of $M_{s+1/2}^{!,+}$ (see [DJ08b, Section 2 and Appendix] for a discussion of this fact).
Lemma 6.3.
Let $2s\in\{4,6,8,10,14\}$, $p$ be a prime, and $r\in\mathbb{Z}^{+}$. Let $f\in M_{s+1/2}^{!,+}$ be $p$-integral and suppose that $p^{r}{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}(f)$ for all $n\leq 0$. Then, $f\equiv 0\pmod{p^{r}}$.
Proof.
This essentially follows from the argument in the last part of the proof of [PZ22, Lemma 2], which we will recall here. As $(f_{s+1/2,m})_{m}$ forms an integral basis of $M_{s+1/2}^{!,+}$, we must have
$f\;=\;\sum_{n\leq 0}a_{n}(f)f_{s+1/2,-n},$
so $f\equiv 0\pmod{p^{r}}$. ∎
Lemma 6.4.
Let $2s\in\{4,6,8,10,14\}$, $m\in\mathbb{Z}^{+}$ with $(-1)^{s-1}m\equiv 0,1\pmod{4}$, and $p$ be a prime. Suppose that
  1. (1)
    $p^{2}\nmid m$, or
  2. (2)
    if $p=2$ and $4{\hskip 1.0pt}|{\hskip 1.0pt}m$, then $(-1)^{s-1}{\hskip 1.0pt}\frac{m}{4}\equiv 2,3\pmod{4}$.
Let
$g_{0}$ $\;:=\;p^{s-1}\big(\tfrac{(-1)^{s-1}m}{p}\big)\cdot f_{s+1/2,m}$
$g_{1}$ $\;:=\;f_{s+1/2,m}$
$g_{i+1}$ $\;:=\;\frac{1}{p^{2s-1}}\cdot\big(g_{i}|T_{p,s+1/2}-g_{i-1}\big).$
Then, for all $i\in\mathbb{Z}^{+}$,
$g_{i}\;=\;q^{-mp^{2i-2}}+O(q)\;=\;f_{s+1/2,mp^{2i-2}}\in\mathbb{Z}[[q]].$
Proof.
Write $f_{m}=f_{s+1/2,m}$. The case when $i=1$ follows from the definition of $f_{m}$. For $i=2$, it is easy to check that the assumptions on $p$ and $m$ imply that
$g_{1}|T_{p,s+1/2}-g_{0}$ $\;=\;\big(p^{2s-1}q^{-mp^{2}}+p^{s-1}\big(\tfrac{(-1)^{s-1}m}{p}\big)q^{-m}+O(q)\big)-\big(p^{s-1}\big(\tfrac{(-1)^{s-1}m}{p}\big)q^{-m}+O(q)\big)$
$\;=\;p^{2s-1}q^{-mp^{2}}+O(q).$
It then follows from Lemma 6.3 that
$g_{2}\;=\;\frac{1}{p^{2s-1}}\cdot\big(g_{1}|T_{p,s+1/2}-g_{0}\big)\;=\;q^{-mp^{2}}+O(q)\in\mathbb{Z}[[q,q^{-1}]].$
Now, we proceed by induction. For $i\geq 2$, by induction hypothesis,
$g_{i}|T_{p,s+1/2}-g_{i-1}$ $\;=\;\big(p^{2s-1}q^{-mp^{2i}}+q^{-mp^{2i-4}}+O(q)\big)-\big(q^{-mp^{2i-4}}+O(q)\big)$
$\;=\;p^{2s-1}q^{-mp^{2i}}+O(q).$
It then follows from Lemma 6.3 that
$g_{i+1}\;=\;\frac{1}{p^{2s-1}}\cdot\big(g_{i}|T_{p,s+1/2}-g_{i-1}\big)\;=\;q^{-mp^{2i}}+O(q)\in\mathbb{Z}[[q,q^{-1}]].$
Now, recall that the Shimura lift is Hecke-equivariant. That is, for $s\in\mathbb{Z}$, $f\in M_{s+1/2}^{!,+}$, and a prime $p$
$\mathcal{S}_{d_{0}}(f|T_{p^{2},s+1/2})\;=\;\mathcal{S}_{d_{0}}(f)|T_{p,2s},$
where $d_{0}$ is some fundamental discriminant with $(-1)^{s}d_{0}>0$.
Proposition 6.5.
Let $2s\in\{4,6,8,10,14\}$ and let $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$. Let $m\in\mathbb{Z}$ with $m\geq 0$ and $(-1)^{s-1}m\equiv 0,1\pmod{4}$ and let $p$ be a prime such that
  1. (1)
    $p^{2}\nmid m$, or
  2. (2)
    if $p=2$ and $4{\hskip 1.0pt}|{\hskip 1.0pt}m$, then $(-1)^{s-1}{\hskip 1.0pt}\frac{m}{4}\equiv 2,3\pmod{4}$.
Let
$F_{m}\;:=\;\mathcal{S}_{d_{0}}f_{s+1/2,m}$
Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{m})\;\equiv\;p^{s-1}\big(\tfrac{(-1)^{s-1}m}{p}\big)\cdot a_{np^{l-1}}(F_{m})\pmod{p^{(2s-1)l}}$.
In particular, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{m})\;\equiv\;0\pmod{p^{(s-1)l}}$.
Proof.
Let $g_{i}\in M_{s+1/2}^{!,+}$ be as in Lemma 6.4 and let $G_{i}=\mathcal{S}_{d_{0}}g_{i}$. In particular,
$G_{0}$ $\;=\;p^{s-1}\big(\tfrac{(-1)^{s-1}m}{p}\big)\cdot F_{m}$
$G_{1}$ $\;=\;F_{m}$
$G_{i+1}$ $\;=\;\frac{1}{p^{2s-1}}\cdot\big(G_{i}|T_{p,2s}-G_{i-1}\big)\;=\;\frac{1}{p^{2s-1}}\big(G_{i}|U_{p}-G_{i-1}\big)+G_{i}|V_{p}\in\mathbb{Z}[[q]]$
Write $H_{i}=G_{i}|U_{p}-G_{i-1}$. Then,
$H_{i+1}\;=\;G_{i+1}|U_{p}-G_{i}\;=\;\bigg(\frac{1}{p^{2s-1}}\big(G_{i}|U_{p}-G_{i-1}\big)+G_{i}|V_{p}\bigg)\bigg|U_{p}-G_{i}\;=\;\frac{1}{p^{2s-1}}{\hskip 1.0pt}H_{i}|U_{p}.$
By induction, it follows that
$H_{1}|U_{p}^{l}\;=\;p^{(2s-1)l}H_{l+1}.$
Hence,
$F_{m}|U_{p}^{l}-p^{s-1}\big(\tfrac{(-1)^{s-1}m}{p}\big)\cdot F_{m}|U_{p}^{l-1}\;=\;H_{1}|U_{p}^{l-1}$ $\;=\;p^{(2s-1)(l-1)}H_{l}$
$\;\equiv\;p^{(2s-1)(l-1)}\big(G_{l}|T_{p,2s}-G_{l-1}\big)$
$\;=\;p^{(2s-1)l}G_{l+1}\;\equiv\;0\pmod{p^{(2s-1)l}}$.
The result then follows. ∎
Proposition 6.6.
Let $2s\in\{4,6,8,10,14\}$ and let $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$. Let $m\in\mathbb{Z}$ with $m\geq 0$ and $(-1)^{s-1}m\equiv 0,1\pmod{4}$ and let $p$ be a prime. Let
$F_{m}\;:=\;\mathcal{S}_{d_{0}}f_{s+1/2,m}.$
Then, for all $n,l\in\mathbb{Z}^{+}$,
$p^{(s-1)t}\cdot a_{np^{l}}(F_{m})\;\equiv\;0\pmod{p^{(s-1)l}}$,
where $t$ is the largest integer such that $p^{2t}{\hskip 1.0pt}|{\hskip 1.0pt}m$ and $(-1)^{s-1}{\hskip 1.0pt}m/p^{2t}\equiv 0,1\pmod{4}$.
Proof.
We will treat $t\geq 0$ as a variable and prove this by induction on $t$. Note that the base case when $t=0$ follows from Proposition 6.5. Let $m^{\prime}=m/p^{2t}$. Then, $(-1)^{s-1}m^{\prime}\equiv 0,1\pmod{4}$ and the pair $(p,m^{\prime})$ satisfies the assumption in Proposition 6.5. Define $G_{i}$ as in the proof of Proposition 6.5 with respect to $m^{\prime}$. Then, by Lemma 6.4,
$G_{i+1}$ $\;=\;\mathcal{S}_{d_{0}}f_{s+1/2,{\hskip 1.0pt}m^{\prime}p^{2i}}\;=\;F_{m^{\prime}p^{2i}}.$
Let $H_{i}=G_{i}|U_{p}-G_{i-1}$ as before. Then,
$G_{t+1}|U_{p}^{l}\;=\;(H_{t+1}+G_{t})|U_{p}^{l-1}\;=\;p^{-(2s-1)t}H_{1}|U_{p}^{t+l-1}+G_{t}|U_{p}^{l-1}.$
Suppose that $t\geq 1$. Note that by the last equation in the proof of Proposition 6.5,
$H_{1}|U_{p}^{t+l-1}\;\equiv\;0\pmod{p^{(2s-1)(t+l)}}$
and that by induction hypothesis,
$p^{(s-1)(t-1)}\cdot G_{t}|U_{p}^{l-1}\;\equiv\;0\pmod{p^{(s-1)(l-1)}}$.
It follows that
$p^{(s-1)t}\cdot G_{t+1}|U_{p}^{l}\;\equiv\;0\pmod{p^{(s-1)l}}$.
The result then follows. ∎
Proposition 6.7.
Let $2s\in\{4,6,8,10,14\}$ and let $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$. Let $m\in\mathbb{Z}$ with $m\geq 0$ and $(-1)^{s-1}m\equiv 0,1\pmod{4}$ and let $A$ be the largest integer such that $A^{2}{\hskip 1.0pt}|{\hskip 1.0pt}m$ and $(-1)^{s-1}m/A^{2}\equiv 0,1\pmod{4}$. Then,
$F\;:=\;A^{s-1}\mathcal{S}_{d_{0}}f_{s+1/2,m}$
is (s1)(s-1)-magnetic, i.e., $n^{s-1}{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}(F)$ for all $n\in\mathbb{Z}^{+}$.
Proof.
Fix $n\in\mathbb{Z}^{+}$ and a prime $p$. It suffices to prove that if $p^{l}{\hskip 1.0pt}|{\hskip 1.0pt}n$, then $p^{(s-1)l}{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}(F)$. Let $t=v_{p}(A)$ and we will prove that $t$ is the largest integer such that $p^{2t}{\hskip 1.0pt}|{\hskip 1.0pt}m$ and $(-1)^{s-1}{\hskip 1.0pt}m/p^{2t}\equiv 0,1\pmod{4}$. Indeed, $p^{2t}{\hskip 1.0pt}|{\hskip 1.0pt}m$ since $p^{2t}{\hskip 1.0pt}|{\hskip 1.0pt}A^{2}$ and $A^{2}{\hskip 1.0pt}|{\hskip 1.0pt}m$, and
$(-1)^{s-1}\frac{m}{p^{2t}}\;=\;\bigg(\frac{A}{p^{t}}\bigg)^{2}(-1)^{s-1}{\hskip 1.0pt}\frac{m}{A^{2}}\;\equiv\;0,1\pmod{4}$.
Suppose that $t^{\prime}>t$ with $p^{2t^{\prime}}{\hskip 1.0pt}|{\hskip 1.0pt}m$ and $(-1)^{s-1}{\hskip 1.0pt}m/p^{2t^{\prime}}\equiv 0,1\pmod{4}$. Then,
$(-1)^{s-1}\frac{m}{p^{2t^{\prime}}}\;=\;\bigg(\frac{Ap^{t^{\prime}-t}}{p^{t^{\prime}}}\bigg)^{2}(-1)^{s-1}{\hskip 1.0pt}\frac{m}{(Ap^{t^{\prime}-t})^{2}}.$
By the maximality of $A$, we must have
$(-1)^{s-1}\frac{m}{p^{2t^{\prime}}}\;\equiv\;0\pmod{4}\quad\text{ and }\quad\bigg(\frac{Ap^{t^{\prime}-t}}{p^{t^{\prime}}}\bigg)^{2}\;\equiv\;0\pmod{4}$,
so in particular $2{\hskip 1.0pt}|{\hskip 1.0pt}A$ and $p\neq 2$. This then implies that
$(-1)^{s-1}\frac{m}{(Ap^{t^{\prime}-t})^{2}}\;=\;(-1)^{s-1}\frac{m}{A}\cdot\frac{1}{p^{2t^{\prime}-2t}}\;\equiv\;0,1\pmod{4}$,
contradicting the maximality of $A$.
Now, by Proposition 6.6, if $p^{l}{\hskip 1.0pt}|{\hskip 1.0pt}n$, then $p^{(s-1)l}{\hskip 1.0pt}|{\hskip 1.0pt}p^{(s-1)t}a_{n}(F_{m})$, where $F_{m}=\mathcal{S}_{d_{0}}f_{s+1/2,m}$. In particular, if $p^{l}{\hskip 1.0pt}|{\hskip 1.0pt}n$, then
$p^{(s-1)l}{\hskip 1.0pt}|{\hskip 1.0pt}A^{s-1}a_{n}(F_{m})\;=\;a_{n}(F).$
The result then follows. ∎
6.3. Proof of Theorem 6.1 when $k/2$ is even
Let $s=k/2$. As $s$ is even, one can apply Proposition 6.2 with $d_{0}=1$ and $d=D$ to obtain that
$\big(\mathcal{S}_{1}f_{s+1/2,-D}\big)(z)\;=\;(-1)^{\frac{s}{2}}{\hskip 1.0pt}|D|^{-\frac{s}{2}}{\hskip 1.0pt}\mathrm{Tr}_{D,1}\big((\partial_{\tau}^{s-1}G_{2s})(z,\tau)\big).$
Since $d_{0}=1$, we have that $\chi(Q)=1$ for all $Q\in\mathcal{Q}_{D}$. Hence, for an $\mathrm{SL}_{2}(\mathbb{Z})$-invariant function $f(\tau)$ on $\mathcal{H}$,
(15) $\mathrm{Tr}_{D,1}(f)\;=\;\sum_{Q\in\mathcal{Q}_{D}}\frac{1}{w_{Q}}f(\alpha_{Q}).$
Write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}^{+}$ and some fundamental discriminant $D_{0}<0$. Then,
(16) $\mathcal{Q}_{D}\;=\;\bigsqcup_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\frac{A}{A^{\prime}}\cdot\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}},$
where $\frac{A}{A^{\prime}}\cdot\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}=\{\frac{A}{A^{\prime}}\cdot Q\mid Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}\}$. One can thus further write Equation 15 as
$\mathrm{Tr}_{D,1}(f)\;=\;\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\frac{2}{w_{A^{\prime 2}D_{0}}}f(\alpha_{Q}),$
where we use that $w_{A^{\prime 2}D_{0}}=2w_{Q}$ for all $Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}$.
To summarize, we obtain that
$\big(\mathcal{S}_{1}f_{s+1/2,-D}\big)(z)$ $\;=\;(-1)^{\frac{s}{2}}{\hskip 1.0pt}|D|^{-\frac{s}{2}}{\hskip 1.0pt}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\frac{2}{w_{A^{\prime 2}D_{0}}}{\hskip 1.0pt}(\partial_{\tau}^{s-1}G_{2s})(z,\alpha_{Q})$
$\;=\;(-1)^{\frac{s}{2}}{\hskip 1.0pt}|D|^{-\frac{s}{2}}{\hskip 1.0pt}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\mathcal{G}_{2s,A^{\prime 2}D_{0}}(z).$
By the Möbius inversion formula,
$\mathcal{G}_{2s,A^{2}D_{0}}$ $\;=\;\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\mu\big(\tfrac{A}{A^{\prime}}\big)(-1)^{\frac{s}{2}}{\hskip 1.0pt}|A^{\prime 2}D_{0}|^{\frac{s}{2}}\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}$
$\;=\;(-1)^{\frac{s}{2}}|D_{0}|^{\frac{s}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\mu\big(\tfrac{A}{A^{\prime}}\big)A^{\prime s}{\hskip 1.0pt}\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}.$
As $s$ is even, one can apply Proposition 6.7 to obtain that each $A^{\prime s-1}{\hskip 1.0pt}\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}$ is (s1)(s-1)-magnetic. It thus follows that $|D_{0}|^{-\frac{k}{4}}\mathcal{G}_{k,D}=|D_{0}|^{-\frac{s}{2}}\mathcal{G}_{2s,A^{2}D_{0}}$ has integer Fourier coefficients and is $\frac{k-2}{2}$-magnetic. Part (3) of Theorem 6.1 now follows from Proposition 6.5.
6.4. Proof of Theorem 6.1 when $k/2$ is odd
Let $s=k/2$ and write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}^{+}$ and some fundamental discriminant $D_{0}<0$ as before. As $s$ is odd, one can apply Proposition 6.2 with $d_{0}=D_{0}$ and $d=A^{2}$ to obtain that
$\big(\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}\big)(z)\;=\;(-1)^{\frac{s+1}{2}}A^{-s}{\hskip 1.0pt}|D_{0}|^{\frac{s-1}{2}}{\hskip 1.0pt}\mathrm{Tr}_{A^{2},D_{0}}\big((\partial_{\tau}^{s-1}G_{2s})(z,\tau)\big).$
In this case, by Equation 16 and Lemma 6.8, we have that for an $\mathrm{SL}_{2}(\mathbb{Z})$-invariant function $f(\tau)$ on $\mathcal{H}$,
$\mathrm{Tr}_{A^{2},D_{0}}(f)\;=\;\sum_{Q\in\mathcal{Q}_{D}}\frac{1}{w_{Q}}\chi(Q)f(\alpha_{Q})\;=\;\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\frac{2}{w_{A^{\prime 2}D_{0}}}\bigg(\frac{D_{0}}{A/A^{\prime}}\bigg){\hskip 1.0pt}f(\alpha_{Q}).$
To summarize, we obtain that
$\big(\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}\big)(z)$ $\;=\;(-1)^{\frac{s+1}{2}}A^{-s}{\hskip 1.0pt}|D_{0}|^{\frac{s-1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\frac{2}{w_{A^{\prime 2}D_{0}}}\bigg(\frac{D_{0}}{A/A^{\prime}}\bigg){\hskip 1.0pt}(\partial_{\tau}^{s-1}G_{2s})(z,\alpha_{Q})$
$\;=\;(-1)^{\frac{s+1}{2}}A^{-s}{\hskip 1.0pt}|D_{0}|^{\frac{s-1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\bigg(\frac{D_{0}}{A/A^{\prime}}\bigg)\mathcal{G}_{2s,A^{\prime 2}D_{0}}(z).$
Hence,
$A^{s}{\hskip 1.0pt}\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}\;=\;(-1)^{\frac{s+1}{2}}|D_{0}|^{\frac{s-1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\bigg(\frac{D_{0}}{A/A^{\prime}}\bigg)\mathcal{G}_{2s,A^{\prime 2}D_{0}}(z).$
Here the left hand side is (s1)(s-1)-magnetic by Proposition 6.7. It follows from induction (on $A$) that $|D_{0}|^{\frac{k-2}{4}}\mathcal{G}_{k,D}=|D_{0}|^{\frac{s-1}{2}}\mathcal{G}_{2s,A^{2}D_{0}}$ has integer coefficients and is $\frac{k-2}{2}$-magnetic. Similarly, Part (3) of Theorem 6.1 follows from Proposition 6.5 and induction.
The following is a lemma which should follow from the theory of the genus characters of quadratic forms. We still include a proof for completeness.
Lemma 6.8.
Let $D=A^{2}D_{0}<0$ be a discriminant with $A\in\mathbb{Z}^{+}$ and $D_{0}<0$ being a fundamental discriminant. Let $Q\in\mathcal{Q}_{D}$ and suppose that $Q$ is $\mathrm{SL}_{2}(\mathbb{Z})$-equivalent to $A_{Q}\cdot Q^{\prime}$ for some $A_{Q}\in\mathbb{Z}^{+}$ with $A_{Q}{\hskip 1.0pt}|{\hskip 1.0pt}A$ and $Q^{\prime}\in\mathcal{Q}_{D/A_{Q}^{2}}^{\mathrm{prim}}$. Then, $\chi_{A^{2},D_{0}}(Q)=\big(\frac{D_{0}}{A_{Q}}\big)$.
Proof.
Since $Q^{\prime}$ is primitive, there exists a prime $p^{\prime}$ with $(p^{\prime},D)=1$ such that $p^{\prime}=Q^{\prime}(x^{\prime},y^{\prime})$ for some $x^{\prime},y^{\prime}\in\mathbb{Z}$. Then,
$\chi_{A^{2},D_{0}}(Q)\;=\;\bigg(\frac{D_{0}}{A_{Q}\cdot p^{\prime}}\bigg)\;=\;\bigg(\frac{D_{0}}{A_{Q}}\bigg)\bigg(\frac{D_{0}}{p^{\prime}}\bigg),$
so it suffices to prove that $\big(\frac{D_{0}}{p^{\prime}}\big)=1$. Since $\gcd(x^{\prime},y^{\prime})=1$ (as $p^{\prime}=Q^{\prime}(x^{\prime},y^{\prime})$ is a prime), there exist $u,v\in\mathbb{Z}$ such that $x^{\prime}u-y^{\prime}v=1$. Then,
$Q^{\prime}(x^{\prime}X+vY,y^{\prime}X+uY)\;=\;p^{\prime}X^{2}+BXY+CY^{2}$
for some $B,C\in\mathbb{Z}$. In particular,
$D_{0}\cdot(A/A_{Q})^{2}\;=\;D/A_{Q}^{2}\;=\;B^{2}-4p^{\prime}C,$
so $D_{0}$ is a square modulo $p^{\prime}$, i.e., $\big(\frac{D_{0}}{p^{\prime}}\big)=1$. ∎
7. Other remarks
In the final section, we will discuss the cases when $j(C)\in\{0,1728\}$ and the generalization to other weights.
7.1. The cases when $j(C)\in\{0,1728\}$
While the heuristics in this case are essentially the same as in the other CM cases, we will first briefly discuss the behaviors of $\frac{E_{k}}{j}$ and $\frac{E_{k}}{j-1728}$ for motivation.
The reason why $\frac{E_{k}}{j}$ and $\frac{E_{k}}{j-1728}$ behave differently is that they do not necessarily have a simple pole and thus resemble $G_{k,D}^{(r)}$ for some $r>1$ rather than $r=1$. Indeed, the following holds.
  1. (1)
    $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$ are both $1$-magnetic [LN19]
  2. (2)
    $\frac{E_{6}}{j}$ is $2$-magnetic.
  3. (3)
    $\frac{E_{8}}{j-1728}$ is (numerically) $1$-magnetic.
  4. (4)
    $\frac{E_{10}}{j}$ is (numerically) $1$-magnetic.
When $\frac{E_{k}}{j}$ or $\frac{E_{k}}{j-1728}$ does have a simple pole, all the phenomena in Section 2 still hold for $\frac{E_{k}}{j}$ or $\frac{E_{k}}{j-1728}$, though one would need to replace the assumption $v_{p}(j(C))=0=v_{p}(j(C)-1728)$ by something appropriate.
Example 7.1.
Consider $\frac{E_{14}}{j}$. Let $C/\mathbb{Q}$ be the elliptic curve given by $y^{2}+y=x^{3}$ with LMFDB label 27.a4 [LMF25, Elliptic Curve 27.a4] and $j(C)=0$. Then, we have the following numerical congruences.
  1. (1)
    For all primes $p$ with $(\tfrac{-3}{p})=1$ and all $n,l\in\mathbb{Z}^{+}$,
    $a_{np^{l}}\bigg(\frac{E_{14}}{j}\bigg)\;\equiv\;u_{p}(C)^{12}a_{np^{l-1}}\bigg(\frac{E_{14}}{j}\bigg)\pmod{p^{13l}}$,
    where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ that is a $p$-adic unit.
  2. (2)
    For all primes $p$ with $p\geq 13$ and $(\tfrac{-3}{p})=-1$ and all $n,l\in\mathbb{Z}^{+}$,
    $a_{np^{l}}\bigg(\frac{E_{14}}{j}\bigg)\;\equiv\;p^{12}a_{np^{l-2}}\bigg(\frac{E_{14}}{j}\bigg)\pmod{p^{13l-1}}$.
We speculate that the reason that there is no need to assume $v_{p}(j(C))=0=v_{p}(j(C)-1728)$ is that both $\frac{E_{14}}{j}$ and $\frac{E_{14}}{j-1728}$ have a simple pole, and that the appearance of $13$ in the supersingular case should be k1k-1 in general.
Now, we turn to the general case. Let $k\in\{4,6,8,10,14\}$ and $C/\mathbb{Q}$ be an elliptic curve with $j(C)\in\{0,1728\}$. Let $D\in\{-3,-4\}$ with $\alpha_{-3}=\frac{1+\sqrt{-3}}{2}$ and $\alpha_{-4}=\sqrt{-1}$. Consider
$\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(C)},\;\frac{E_{k}}{(j-j(C))^{2}},\ldots,\frac{E_{k}}{(j-j(C))^{\kappa}}\bigg\},$
where $\kappa=\kappa_{k,C}$ is the largest integer such that $\frac{E_{k}}{(j-j(C))^{\kappa}}$ has a pole of order $\leq k-1$. For $1\leq i\leq\kappa$, let $r_{i}$ denote the order of the pole of $\frac{E_{k}}{(j-j(C))^{i}}$. Then, one still considers for $1\leq i\leq\kappa$,
$G_{k,D}^{(r_{i})}(z)\;=\;(\partial_{\tau}^{r_{i}-1}G_{k})(z,\alpha_{D}).$
As in the other CM cases, we make a choice of $G_{k,D}^{(r_{i})}$ following the process described in Section 4. Now, the family $\{G_{k,D}^{(r_{i})}\}_{1\leq i\leq\kappa}$ should satisfy all the phenomena in Section 4. Indeed, one would also need to replace the condition that $v_{p}(j(C))=0=v_{p}(j(C)-1728)$ by something appropriate. We illustrate this through some examples.
Example 7.2.
Let $k=4$. We have $\kappa=1$ and $r_{1}=2$ for both $j(C)\in\{0,1728\}$. In this case,
$G_{4,-3}^{(2)}\;=\;\frac{E_{4}}{j}\quad\text{ and }\quad G_{4,-4}^{(2)}\;=\;\frac{E_{4}}{j-1728}.$
From these identifications, the $1$-magnetic properties of $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$, i.e., Theorem 1.1, are now accounted for by 4.4, while the supercongruences, i.e., Theorem 1.2, are now explained by Theorem 4.7 (and follow from Theorem 6.1 by plugging in $k=4$ and D=3,4D=-3,-4).
Example 7.3.
Let $k=6$ and $j(C)=0$. We have $\kappa=1$ and $r_{1}=3$. In this case,
$G_{6,-3}^{(3)}\;=\;\frac{E_{6}}{j}.$
Now, 4.4 predicts that $\frac{E_{6}}{j}$ is $2$-magnetic, while Theorem 4.7 predicts the following supercongruence: for all primes $p\geq 5$ and all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}\bigg(\frac{E_{6}}{j}\bigg)\;\equiv\;p^{2}{\hskip 1.0pt}a_{np^{l-1}}\bigg(\frac{E_{6}}{j}\bigg)\pmod{p^{5l}}$.
Here the $2$-magnetic property is proved in [PZ22, Theorem 2], and the supercongruence follows from Theorem 6.1 by plugging in $k=6$ and D=3D=-3.
Example 7.4.
Let $k=6$ and $j(C)=1728$. To be more explicit, let $C/\mathbb{Q}$ be the elliptic curve given by $y^{2}=x^{3}-x$ with LMFDB label 32.a3 [LMF25, Elliptic Curve 32.a3]. We have $\kappa=3$ and $r_{1}=1,r_{2}=3,r_{3}=5$. In this case,
$G_{6,-4}^{(1)}$ $\;=\;\frac{E_{6}}{j-1728}$
$G_{6,-4}^{(3)}$ = $\;=\;13\cdot\frac{E_{6}}{j-1728}+31104\cdot\frac{E_{6}}{(j-1728)^{2}}$
$G_{6,-4}^{(5)}$ = $\;=\;277\cdot\frac{E_{6}}{j-1728}+2571264\cdot\frac{E_{6}}{(j-1728)^{2}}+3869835264\cdot\frac{E_{6}}{(j-1728)^{3}}.$
Then, the family $\{G_{6,-4}^{(1)},G_{6,-4}^{(3)},G_{6,-4}^{(5)}\}$ numerically satisfies all the congruences in Section 4, where one considers primes $p$ with $p\geq 5$. For example, we have the following numerical congruences.
  1. (1)
    For all primes $p$ with $p\geq 5$ and $(\frac{-4}{p})=1$ and all $n,l\in\mathbb{Z}^{+}$,
    $a_{np^{l}}(G_{6,-4}^{(r)})\;\equiv\;u_{p}(C)^{5-r}p^{r-1}a_{np^{l-1}}(G_{6,-4}^{(r)})\pmod{p^{5l}}$
    where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ that is a $p$-adic unit.
  2. (2)
    For all primes $p$ with $p\geq 5$ and $(\frac{-4}{p})=-1$ and all $n,l\in\mathbb{Z}^{+}$,
    $a_{np^{l}}(G_{6,-4}^{(r)})\;\equiv\;p^{4}{\hskip 1.0pt}a_{np^{l-2}}(G_{6,-4}^{(r)})\pmod{p^{5l-r}}$.
  3. (3)
    $G_{6,-4}^{(3)}$ is $2$-magnetic and for all primes $p\geq 5$ and all $n,l\in\mathbb{Z}^{+}$,
    $a_{np^{l}}(G_{6,-4}^{(3)})\;\equiv\;p^{2}{\hskip 1.0pt}a_{np^{l-1}}(G_{6,-4}^{(3)})\pmod{p^{5l}}$.
    This (almost) follows from Theorem 6.1 by plugging in $k=6$ and D=4D=-4.
  4. (4)
    For all primes $p\geq 5$ and all $n,l\in\mathbb{Z}^{+}$,
    $a_{np^{l}}(G_{6,-4}^{(1)}){\hskip 1.0pt}a_{np^{l}}(G_{6,-4}^{(3)})\;\equiv\;p^{4}{\hskip 1.0pt}a_{np^{l-1}}(G_{6,-4}^{(1)}){\hskip 1.0pt}a_{np^{l-1}}(G_{6,-4}^{(3)})\pmod{p^{5l}}$.
7.2. Generalization to other weights
We will now discuss the generalization of the whole setting to other weights. In general, the coefficients of meromorphic modular forms having precisely one pole at a non-cuspidal point are an interplay between cusp forms and the elliptic curve corresponding to the pole.
Let $g\in\mathbb{Z}[[q]]$ be a holomorphic modular form of weight $k$ and level $1$ and let $C/\mathbb{Q}$ be an elliptic curve (with $j(C)\notin\{0,1728\}$) such that $\frac{g}{j-j(C)}$ has a pole. For $r\in\mathbb{Z}^{+}$, write
$F_{g,C}^{(r)}\;:=\;\frac{g}{(j-j(C))^{r}}\;=\;\sum_{n=1}^{\infty}a_{n}(F_{g,C}^{(r)})q^{n}.$
In general, it is necessary to remove the influence from the cusp forms to obtain the same congruences observed for $F_{k,C}^{(r)}$ in Section 2 and Section 3. This is possible via the help of relations. An integer sequence $\lambda=(\lambda_{m})_{m\in\mathbb{Z}^{+}}$ is called a relation for $S_{k}(1)$ if
  1. (1)
    $\lambda_{m}=0$ for all but finitely many $m$;
  2. (2)
    $\sum_{m=1}^{\infty}\lambda_{m}a_{m}\;=\;0$ for any cusp form $\sum_{m=1}^{\infty}a_{m}q^{m}\in S_{k}(1)$.
For a modular form $F$ of weight $k$ and a relation $\lambda$ for $S_{k}(1)$, write
$F|\lambda\;:=\;\sum_{m=1}^{\infty}\lambda_{m}F|T_{m,k},$
where $T_{m,k}$ is the $m$-th Hecke operator of weight $k$. Note that $\sum_{m=1}^{\infty}\lambda_{m}(\sum_{m=1}^{\infty}a_{m}q^{m})|T_{m,k}=0$ for any cusp form $\sum_{m=1}^{\infty}a_{m}q^{m}\in S_{k}(1)$, so replacing $F$ by $F|\lambda$ will “kill the cuspidal part” of $F$. The consideration of such $\lambda$ first appeared in Gross–Zagier [GZ86, p. 316], and this idea was also used in Löbrich–Schwagenscheidt [LS21] to construct examples of magnetic modular forms.
Now, one simply fixes a relation $\lambda$ for $S_{k}(1)$ and considers $F_{g,C}^{(r)}|\lambda$. Then, all the phenomena in Section 2 and Section 3 should still hold after replacing $F_{k,C}^{(r)}$ by $F_{g,C}^{(r)}|\lambda$.
In fact, as in the cases when $k\in\{4,6,8,10,14\}$ where only $g=E_{k}$ is considered, in general, it is also sufficient to consider a fixed holomorphic modular form $g$ of weight $k$ and level $1$. This is because of the following lemma.
Lemma 7.5.
Let $f,g\in\mathbb{Q}[[q]]$ be two holomorphic modular forms of weight $k$ and level $1$. Let $\alpha\in\mathcal{H}$ with $j(\alpha)\in\mathbb{Q}$ and suppose that $g(\alpha)\neq 0$. Let $r\in\mathbb{Z}^{+}$. Then, $\frac{f}{(j-j(\alpha))^{r}}$ is a $\mathbb{Q}$-linear combination of $\frac{g}{j-j(\alpha)},\frac{g}{(j-j(\alpha))^{2}},\ldots,\frac{g}{(j-j(\alpha))^{r}}$, and cusp forms of weight $k$ and level $1$.
Proof.
We will prove this by induction. Note that $f/g$ has rational coefficients and is a modular function of level $1$, so it is a rational polynomial in $j$. In particular, $\lambda:=f(\alpha)/g(\alpha)\in\mathbb{Q}$. Write
$h\;=\;\frac{f-\lambda g}{j-j(\alpha)}\in\mathbb{Q}[[q]].$
Then, $h$ is holomorphic on $\mathcal{H}$ and vanishes at $\infty$, so $h\in S_{k}(1)$. Now, for $r=1$,
$\frac{f}{j-j(\alpha)}\;=\;h+\lambda\cdot\frac{g}{j-j(\alpha)},$
so it satisfies the requirement. For $r>1$,
$\frac{f}{(j-j(\alpha))^{r}}\;=\;\frac{h}{(j-j(\alpha))^{r-1}}+\lambda\cdot\frac{g}{(j-j(\alpha))^{r}}.$
The result then follows by applying the induction hypothesis on $h$ and r1r-1. ∎
This lemma is especially useful in the CM case, since it suffices to consider a particular holomorphic modular form $g$ of weight $k$ and level $1$ such that suitable linear combinations of $\{\frac{g}{(j-j(C))^{r}}\}_{1\leq r\leq k-1}$, after killing the cuspidal part, should still satisfy the congruences in Section 4. As before, we will construct these linear combinations via a two-variable modular form.
For any $k\in 2\mathbb{Z}$, write
$g_{k}\;:=\;E_{k^{\prime}}\Delta^{l},$
where $l,k^{\prime}\in\mathbb{Z}$ with $k=12l+k^{\prime}$ and $k^{\prime}\in\{0,4,6,8,10,14\}$. For $k\in 2\mathbb{Z}_{\geq 2}$, write
$G_{k}(z,\tau)\;:=\;\frac{g_{k}(z)\cdot g_{2-k}(\tau)}{j(z)-j(\tau)}.$
Adopting the notations in Section 4, for $1\leq r\leq k-1$, write
$G_{k,D}^{(r)}(z)\;:=\;(\partial_{\tau}^{r-1}G_{k})(z,\alpha_{D}).$
As before, one can also suitably rescale $G_{k,D}^{(r)}$ so that it has integer coefficients with no common divisors. Now, one simply fixes a relation $\lambda$ for $S_{k}(1)$ and considers $G_{k,D}^{(r)}|\lambda$. Then, all the phenomena in Section 4 should still hold after replacing $G_{k,D}^{(r)}$ with $G_{k,D}^{(r)}|\lambda$.