### v2

# Elliptic curves and Fourier coefficients of meromorphic modular forms

Abstract.
We discuss several congruences satisfied by the coefficients of meromorphic modular forms, or equivalently, the $p$-adic behaviors of meromorphic modular forms under the $U_{p}$ operator, that are summarized from numerical experiments. In the generic case, we observe the connection to symmetric powers of elliptic curves, while in the CM case, we furthermore observe the connection to the $p$-adic analogue of the Chowla–Selberg periods. Along with the discussions, we will provide some heuristic explanations for these congruences as well as prove some of them using hypergeometric functions and the Borcherds–Shimura lift.
Key words and phrases:
meromorphic modular forms, symmetric powers of elliptic curves, magnetic modular forms, supercongruences, Chowla–Selberg periods.
2020 Mathematics Subject Classification:
Primary 11F37, 11F30, 11F33, 11F23; Secondary 11G05, 11G15, 11F27, 33C20
1. Introduction
Coefficients of modular forms have always been one of the central topics in the field of modular forms. The vast majority of work in this field is restricted to the study of holomorphic or weakly holomorphic modular forms. In this paper, we instead turn to meromorphic modular forms. The main objects of interest are meromorphic modular forms of level $1$ with precisely one pole at a non-cuspidal point. We will discuss the conjectural [^1] It should be noted that while this paper heavily adopts the word “conjectures”, they should be more accurately named as numerical observations, as they are summarized from numerical experiments with these meromorphic modular forms. $p$-adic properties possessed by these meromorphic modular forms as well as the heuristic reasons for those properties. In particular, it turns out that these meromorphic modular forms of weight $k$ behave similarly to symmetric (k2)(k-2)-nd powers of elliptic curves. Indeed, the paper by Brown–Fonseca [BF25] already evidences such a connection. We will also discuss the CM case in detail, namely when the non-cuspidal pole is at a CM point, which may be the most interesting part of this paper and deserves thorough study in its own right. In this case, we will construct a family of meromorphic modular forms whose conjectural $p$-adic behaviors can be characterized via the $p$-adic analogue of the Chowla–Selberg periods, which is then related to special values of $p$-adic $L$-functions via Ferrero–Greenberg [FG79].
For ease of exposition, throughout the first half of this paper, we will mostly focus on the cases of weight $k\in\{4,6,8,10,14\}$, and specifically, on modular forms of the form
$A_{1}\cdot\frac{E_{k}}{j-c}+A_{2}\cdot\frac{E_{k}}{(j-c)^{2}}+\cdots+A_{r}\cdot\frac{E_{k}}{(j-c)^{r}},$
where $c\in\mathbb{Q}$, each $A_{i}\in\mathbb{Q}$, and $E_{k}$ is the Eisenstein series of weight $k\in\{4,6,8,10,14\}$. For general weights, we will first remove the influence of cusp forms using relations, or equivalently, certain linear combinations of the Hecke operators, and then investigate the $p$-adic properties of the resulting modular forms. We will discuss this generalization and prove the main results for all weights in the second half of this paper.
1.1. The case of weight 44
We first motivate the subject through several phenomena satisfied by the modular form
$\frac{E_{4}}{j-c}$
for some $c\in\mathbb{Q}$. We start with the cases when $c\in\{0,1728\}$, i.e., $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$. Here we make a definition that a (general) modular form $\sum_{n=1}^{\infty}a_{n}q^{n}\in\mathbb{Z}[[q]]$ is called magnetic if $n{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}$ for all $n\in\mathbb{Z}^{+}$, and specifically $r$-magnetic for some $r\in\mathbb{Z}_{\geq 0}$ if $n^{r}{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}$ for all $n\in\mathbb{Z}^{+}$. We have the following result by Li–Neururer [LN19, Theorem 1.5] and Paşol–Zudilin [PZ22, Theorem 1].
Theorem 1.1.
Both $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$ are $1$-magnetic.
The first numerically observed example of a magnetic modular form appeared in Broadhurst–Zudilin [BZ19, Conjecture 4] with motivation coming from the Hall effect in electromagnetism. The magnetic property of that example, together with that of $\frac{64E_{4}}{j}$, was proved in Li–Neururer [LN19]. Their proof uses the Borcherds–Shimura lift: one identifies the half-integral weight preimage, proves certain divisibility properties for the preimage, and deduces the magnetic property of the image. This method was later generalized by Paşol–Zudilin [PZ22] to show the magnetic property of other meromorphic modular forms, including $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$, all of which have poles at CM points. For a detailed discussion of magnetic modular forms, we also refer to the paper by Bönisch–Duhr–Maggio [BDM24] and the references therein.
By modifying the method of using the Borcherds–Shimura lift, one can also show that both $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$ satisfy the following supercongruences  [^2] Here the word “supercongruences” refers to the type of congruences with modulus $p^{rl}$ for some $r>1$., which almost imply their $1$-magnetic property and in particular confirm that they are not $2$-magnetic. In general, both the magnetic property and the supercongruences occur for meromorphic modular forms with CM poles, as we will see in Section 4.
Theorem 1.2.
For all primes $p\geq 5$ and all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}\bigg(\frac{E_{4}}{j}\bigg)$ $\;\equiv\;\big(\tfrac{-3}{p}\big)p{\hskip 1.0pt}a_{np^{l-1}}\bigg(\frac{E_{4}}{j}\bigg)$ $\pmod{p^{3l}}$
$a_{np^{l}}\bigg(\frac{E_{4}}{j-1728}\bigg)$ $\;\equiv\;\big(\tfrac{-4}{p}\big)p{\hskip 1.0pt}a_{np^{l-1}}\bigg(\frac{E_{4}}{j-1728}\bigg)$ $\pmod{p^{3l}}$,
where $(\frac{\cdot}{p})$ denotes the Legendre symbol.
When $c\notin\{0,1728\}$, one observes something different, with elliptic curves coming into play. Let $C/\mathbb{Q}$ be an elliptic curve. For a good prime $p$ of $C$, define
$a_{p}(C)\;:=\;p+1-|C(\mathbb{F}_{p})|.$
A good prime $p$ of $C$ is called an ordinary (resp. supersingular) prime of $C$ if $p\nmid a_{p}(C)$ (resp. $p{\hskip 1.0pt}|{\hskip 1.0pt}a_{p}(C)$). For $p\geq 5$, this is equivalent to the condition that $a_{p}(C)\neq 0$ (resp. $a_{p}(C)=0$).
Theorem 1.3.
Let $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$ and let $p$ be a good prime of $C$ with $p\geq 5$ and $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then,
$a_{p}\bigg(\frac{E_{4}}{j-j(C)}\bigg)\;\equiv\;a_{p}(C)^{2}\pmod{p}$.
Here the appearance of $a_{p}(C)^{2}$ already hints at the link between $\frac{E_{4}}{j-j(C)}$ and $\mathrm{Sym}^{2}C$, since the trace of Frobenius at $p$ on $\mathrm{Sym}^{2}C$ is $a_{p}(C)^{2}-p$, which is congruent to $a_{p}(C)^{2}$ modulo $p$. When $p$ is supersingular, Theorem 1.3 a priori only implies that $a_{p}(\frac{E_{4}}{j-j(C)})\equiv 0\pmod{p}$, while in fact numerical observations suggest the following stronger congruence.
Conjecture 1.4.
Let $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$ and let $p$ be a supersingular prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then,
$a_{p}\bigg(\frac{E_{4}}{j-j(C)}\bigg)\;\equiv\;0\pmod{p^{2}}$.
In general, the connection between $\frac{E_{4}}{j-j(C)}$ and $\mathrm{Sym}^{2}C$ can be visualized via the following longer congruence, which involves suitable linear combinations of several coefficients of $\frac{E_{4}}{j-j(C)}$.
Conjecture 1.5.
Let $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$ and let $p$ be a good prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Let $\alpha_{p},\beta_{p}$ be two roots of $X^{2}-a_{p}(C)X+p$ and write
$(X-\alpha_{p}^{2})(X-\alpha_{p}\beta_{p})(X-\beta_{p}^{2})\;=:\;X^{3}+c_{p,1}X^{2}+c_{p,2}X+c_{p,3}.$
Write $F=\frac{E_{4}}{j-j(C)}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F)+c_{p,1}{\hskip 1.0pt}a_{np^{l-1}}(F)+c_{p,2}{\hskip 1.0pt}a_{np^{l-2}}(F)+c_{p,3}{\hskip 1.0pt}a_{np^{l-3}}(F)\;\equiv\;0\pmod{p^{3l-3}}$.
Congruences of this form are usually called the Atkin–Swinnerton-Dyer (ASD) congruences. This term originally refers to three-term congruences of the form
$a_{np}-A_{p}{\hskip 1.0pt}a_{n}+\chi(p)p^{k-1}a_{n/p}\;\equiv\;0\pmod{p^{(k-1)(v_{p}(n)+1)}}$
for some sequence $(a_{n})$. In general, it also refers to congruences of the form
$a_{np^{l}}+c_{p,1}{\hskip 1.0pt}a_{np^{l-1}}+c_{p,2}{\hskip 1.0pt}a_{np^{l-2}}+\cdots+c_{p,s}{\hskip 1.0pt}a_{np^{l-s}}\;\equiv\;0\pmod{p^{\bullet}}$
for some suitable power of $p$. Atkin and Swinnerton-Dyer [ASD71] first came across this type of congruences when studying the coefficients of modular forms on non-congruence subgroups, whose result was later generalized by Scholl [Sch85]. In [KS16], Kazalicki and Scholl demonstrated certain ASD congruences for weakly holomorphic modular forms, and as an example, showed that $F=E_{4}^{6}/\Delta-1464E_{4}^{3}$ satisfies that
$a_{np}(F)-\tau(p)a_{n}(F)+p^{11}a_{n/p}(F)\;\equiv\;0\pmod{p^{11v_{p}(n)}}$,
where $\tau$ is the Ramanujan tau function. More recently, Allen, Long, and Saad [ALS25] established certain ASD congruences for meromorphic modular forms on finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$. For an overview of the ASD congruences, we refer to the paper by Li–Long [LL14].
We end the list of phenomena with an explicit connection between the $p$-th coefficient of $\frac{E_{4}}{j-c}$ and the truncated hypergeometric sum associated to the datum $\big((\frac{1}{2},\frac{1}{6},\frac{5}{6}),(1,1)\big)$
[…]
where we adopt notations from Section 7.3. This connection also constitutes a key part of the proof of Theorem 1.3.
Theorem 1.6.
Let $c\in\mathbb{Q}$ and let $p\geq 5$ be a prime with $v_{p}(c)=0$. Then,
$a_{p}\bigg(\frac{E_{4}}{j-c}\bigg)$ $\;\equiv\;\big(c(c-1728)\big)^{\frac{p-1}{2}}\cdot\sum_{m=0}^{p-1}\frac{(6m)!}{(m!)^{3}(3m)!}{\hskip 1.0pt}c^{-m}\pmod{p}$.
1.2. Main results
We now discuss the main results of this paper. We will fix the weight $k\in\{4,6,8,10,14\}$ for ease of exposition, though the results are proved for all weights.
The first result is a generalization of Theorem 1.3 on meromorphic modular forms with a simple pole.
Theorem 1.7.
Let $k\in\{4,6,8,10,14\}$, $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$, and $p\geq 5$ be a good prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n\in\mathbb{Z}^{+}$,
$a_{np}\bigg(\frac{E_{k}}{j-j(C)}\bigg)\;\equiv\;a_{p}(C)^{k-2}a_{n}\bigg(\frac{E_{k}}{j-j(C)}\bigg)\pmod{p}$.
Remark 1.8.
This theorem is conjecturally true for $p=2,3$. In the case of a general weight $k$, conjecturally one can sometimes also remove the assumption $v_{p}(j(C))=0$ or $v_{p}(j(C)-1728)=0$ depending on $k\pmod{12}$ (see Remark 5.2).
Remark 1.9.
We will prove a suitable generalization of this theorem that holds for all weights and for all elliptic curves over number fields (Theorem 7.1). The key ingredient of proof is the hypergeometric connection, i.e., Theorem 1.6.
The second result is on the supercongruences and the magnetic property of certain meromorphic modular forms of weight $k$ with a CM pole of order $k/2$. In order to describe the result, we first introduce some notations.
Recall that for a (nearly holomorphic) modular form $f(\tau)$ of weight $k_{f}$, we may view $f$ as a function in $\tau$ and $\overline{\tau}$ and define the non-holomorphic derivative as
$(\partial_{\tau}f)(\tau)\;:=\;\frac{1}{2\pi i}\frac{\partial f}{\partial\tau}-\frac{k_{f}}{4\pi\operatorname{Im}(\tau)}\cdot f(\tau).$
It is easy to check that $\partial_{\tau}f$ is modular of weight $k_{f}+2$. Thus, it is possible to apply the operator $\partial_{\tau}$ recursively to a modular form.
Let $D<0$ be a discriminant with class number $1$, $\alpha_{D}\in\mathcal{H}$ be a CM point of discriminant $D$, $\mathcal{O}_{D}$ be the imaginary quadratic order of discriminant $D$, and $w_{D}:=|\mathcal{O}_{D}^{\times}|$. For $k\in\{4,6,8,10,14\}$, let  [^3] This two-variable modular form $G_{k}(z,\tau)$ occurred in many places throughout the literature. We refer to [AKN97, p. 98] for a detailed discussion and the references therein. It also plays a key role in Zagier’s duality [Zag02, DJ08a].
$G_{k}(z,\tau)\;:=\;\frac{E_{k}(z)\cdot\frac{E_{14-k}}{\Delta}(\tau)}{j(z)-j(\tau)}$ (1)
and
$\mathcal{G}_{k,D}(z)\;:=\;\frac{2}{w_{D}}\cdot(\partial^{\frac{k-2}{2}}_{\tau}G_{k})(z,\alpha_{D}).$ (2)
Here $G_{k}(z,\tau)$ is modular of weight $k$ and 2k2-k with respect to $z$ and $\tau$ respectively, so $(\partial^{\frac{k-2}{2}}_{\tau}G_{k})(z,\tau)$ is of weight 0 with respect to $\tau$. Hence, $\mathcal{G}_{k,D}(z)$ is independent of the choice of $\alpha_{D}$.
Theorem 1.10.
Let $k\in\{4,6,8,10,14\}$. Let $D<0$ be a discriminant of class number $1$ and write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}$ and some fundamental discriminant $D_{0}<0$. Let
[…]
Then, $\widetilde{\mathcal{G}}_{k,D}$ satisfies the following.
  1. (1)
    $\widetilde{\mathcal{G}}_{k,D}$ has integer Fourier coefficients and is $\frac{k-2}{2}$-magnetic, i.e., for all $n\in\mathbb{Z}^{+}$,
    $n^{\frac{k-2}{2}}{\hskip 1.0pt}\big|{\hskip 1.0pt}a_{n}(\widetilde{\mathcal{G}}_{k,D}).$
  2. (2)
    Let $p$ be a prime with $p\nmid A$. Then, for all $n,l\in\mathbb{Z}^{+}$,
    $a_{np^{l}}(\widetilde{\mathcal{G}}_{k,D})\;\equiv\;\big(\big(\tfrac{D}{p}\big)p\big)^{\frac{k-2}{2}}a_{np^{l-1}}(\widetilde{\mathcal{G}}_{k,D})\pmod{p^{(k-1)l}}$,
    where $(\frac{\cdot}{p})$ denotes the Legendre symbol.
Remark 1.11.
We will prove a suitable generalization of this theorem that holds for all weights and for all discriminants $D<0$ (Theorem 8.1).
Remark 1.12.
This theorem implies the supercongruences satisfied by $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$, namely Theorem 1.2, by choosing $k=4$ and D=3D=-3 or D=4D=-4 respectively. If one chooses $k=6$ and D=3D=-3, then one obtains the following nice supercongruence for $\frac{E_{6}}{j}$: for all primes $p\geq 5$ and all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}\bigg(\frac{E_{6}}{j}\bigg)\;\equiv\;p^{2}{\hskip 1.0pt}a_{np^{l-1}}\bigg(\frac{E_{6}}{j}\bigg)\pmod{p^{5l}}$.
In this case, this theorem also implies that $\frac{384E_{6}}{j}$ is $2$-magnetic, though the $2$-magnetic property of $\frac{E_{6}}{j}$ was already proved in [PZ22, Theorem 2]. We refer to Example 5.4 for more examples of magnetic modular forms deduced from this theorem.
1.3. Structure of the paper
As mentioned before, the first four sections will only focus on the cases of weight $k\in\{4,6,8,10,14\}$, and the remaining four sections will concern general weights. We will also mostly focus on the case when the elliptic curve $C/\mathbb{Q}$ satisfies that $j(C)\notin\{0,1728\}$, and we will carry the assumption that $v_{p}(j(C))=0=v_{p}(j(C)-1728)$ when investigating the $p$-adic properties.
Section 2 will focus on the case of having a simple pole, where we raise the conjectural generalizations of Theorem 1.7 through a supersingular-ordinary dichotomy. We will also discuss the case when the simple pole is at a CM point, where one can already observe some three-term ASD congruences.
Section 3 will discuss the case of having a pole of higher order, and more specifically, will focus on the following space of modular forms
$\mathrm{MMF}_{k,C}\;:=\;\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(C)},\;\frac{E_{k}}{(j-j(C))^{2}},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}\bigg\}.$
We will make comparisons between the space $\mathrm{MMF}_{k,C}$ with the $U_{p}$ action and the space $\mathrm{Sym}^{k-2}C$ with the Frobenius action, and as a generalization of 1.5, we will conjecture a suitable ASD congruence satisfied by modular forms in $\mathrm{MMF}_{k,C}$, which is given by the characteristic polynomial of the Frobenius action on $\mathrm{Sym}^{k-2}C$.
Section 4 will focus on the CM case. For a discriminant D<4D<-4 (of class number $1$), we will specifically focus on the case of having a pole (of order $\leq k-1$) at a CM point $\alpha_{D}$ of discriminant $D$, or equivalently, the space $\mathrm{MMF}_{k,D}:=\mathrm{MMF}_{k,C}$ when $C/\mathbb{Q}$ is a CM elliptic curve with CM by the imaginary quadratic order of discriminant $D$. We will construct a (rational) basis $\{G_{k,D}^{(r)}\}_{r=1}^{k-1}$ for the space $\mathrm{MMF}_{k,D}$, with each $G_{k,D}^{(r)}$ having a pole at $\alpha_{D}$ of order $r$, such that, roughly speaking, the basis is conjecturally
  1. (1)
    diagonal under the action of (powers of) $U_{p}$ for ordinary primes $p$, and
  2. (2)
    anti-diagonal under the action of (powers of) $U_{p}$ for supersingular primes $p$
via explicit supercongruences, where the (anti-)eigenvalues are described explicitly using the $p$-adic analogue of the Chowla–Selberg periods. Moreover, up to some nonzero integer multiple, each $G_{k,D}^{(r)}$ is conjecturally $(r^{\prime}-1)$-magnetic, where $r^{\prime}=\min(r,k-r)$. At the end of Section 4, we will also discuss the case when $D\in\{-3,-4\}$, or equivalently, $j(C)\in\{0,1728\}$.
The last four sections will focus on general weights and proofs of the main results. Section 5 will discuss the generalization to other weights via the help of relations and restate the main results and conjectures for general weights. Section 6 will recall the notations and preliminaries that are needed for the proofs of the main results. Section 7 will discuss the hypergeometric connection, namely Theorem 1.6, and use this to prove a generalization of Theorem 1.7 for all weights and for elliptic curves over number fields. Section 8 will discuss the proof of a generalization of Theorem 1.10 for all weights and all discriminants $D<0$ using the Borcherds–Shimura lift.
Throughout the paper, we also include many numerical examples under the help of LMFDB [LMF26] and PARI/GP [PAR23].
1.4. Related work
One of the main motivations of this paper is to obtain an explicit (though possibly conjectural) connection between meromorphic modular forms and symmetric powers of elliptic curves through numerically verifiable congruences. As mentioned before, the connection itself was already discussed in Brown–Fonseca [BF25], which built a comprehensive theory for meromorphic modular forms that is very closely related to this paper.
Indeed, in view of Kazalicki–Scholl [KS16], it is natural to expect the congruences that link meromorphic modular forms to symmetric powers of elliptic curves to be in the form of the ASD congruences. In this direction, the work of Allen–Long–Saad [ALS25] established certain ASD congruences for meromorphic modular forms on finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$ by extending the methods of Scholl [Sch85] and Kazalicki–Scholl [KS16]. They also treated the CM case in detail and deduced the existence of the (anti-)diagonal basis, though without explicitly identifying the basis and the (anti-)eigenvalues. Indeed, their main results should partially imply some of the congruences conjectured in this paper, at least when the weight $k\in\{4,6,8,10,14\}$ and when $p\geq k-1$. [^4] One should compare the appearances of $p\geq k-1$ here and below with the appearance of factorials in 5.3 and 5.8. However, it is not immediately clear whether the remaining cases follow as well, since our approach to general weights is different from theirs.
Another main issue that this paper encounters is the lack of Hecke operators on the space of meromorphic modular forms, or specifically, on the space $\mathrm{MMF}_{k,C}$. A reasonable workaround that one may expect is via the theory of $p$-adic modular forms. From this viewpoint, the work of Bordignon [Bor26] related meromorphic modular forms to overconvergent $p$-adic modular forms and described the explicit action of the $U_{p}$ operator via comparison theorems between rigid and de Rham cohomology, making it possible to deduce certain congruences of meromorphic modular forms by using the integrality arising from crystalline cohomology, and hence potentially resolving some of the conjectures in this paper when $p\geq k-1$.
Acknowledgements
The author would like to thank many people for useful discussions on this project, including Kilian Bönisch, Steven Charlton, Chi Hong Chow, Henri Cohen, Håvard Damm-Johnsen, Vasily Golyshev, Wen-Ching Winnie Li, Yingkun Li, Ling Long, Alexandre Maksoud, Danylo Radchenko, Berend Ringeling, Fernando Rodriguez Villegas, Hasan Saad, and Jan Vonk. The author is in particular grateful to Don Zagier and Wadim Zudilin for all the useful ideas on proofs and generalizations.
This project originated from discussions during the International Conference on Modular Forms and $q$-Series at University of Cologne, and during the follow-up workshop at the Hausdorff Research Institute for Mathematics to the trimester program “Periods in Number Theory, Algebraic Geometry and Physics” funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813. Part of the later discussions also took place during the author’s visit to the Abdus Salam International Centre for Theoretical Physics hosted by Vasily Golyshev and Don Zagier, and during the author’s visit to Louisiana State University hosted by Ling Long and partially supported by the Simons Foundation grant MP-TSM-00002492.
2. Simple poles
We will start with the simplest case, i.e., the case of meromorphic modular forms with precisely one pole that is simple and at a non-cuspidal point.
Let $k\in\{4,6,8,10,14\}$ and let $C/\mathbb{Q}$ be an elliptic curve (with $j(C)\notin\{0,1728\}$). Write
$F_{k,C}\;:=\;\frac{E_{k}}{j-j(C)}\;=\;\sum_{n=1}^{\infty}a_{n}(F_{k,C})q^{n}.$
We first recall that Theorem 1.7 implies that for all good primes $p\geq 5$ of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$ and all $n\in\mathbb{Z}^{+}$,
$a_{np}(F_{k,C})\;\equiv\;a_{p}(C)^{k-2}a_{n}(F_{k,C})\pmod{p}$.
From a different point of view, this describes the behavior of $F_{k,C}$ under the $U_{p}$ operator. In particular, it says that $F_{k,C}$, when viewed as a mod $p$ modular form, is a $U_{p}$-eigenform with eigenvalue $a_{p}(C)^{k-2}$. In general, one can also describe the behavior of $F_{k,C}$ under the $U_{p}^{l}$ operator if one considers the supersingular and the ordinary cases separately.
Conjecture 2.1.
Let $k\in\{4,6,8,10,14\}$, $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$, and $p$ be a supersingular prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,C})\;\equiv\;p^{k-2}a_{np^{l-2}}(F_{k,C})\pmod{p^{(k-1)l-1}}$,
where $a_{m}(\;\cdot\;):=0$ if $m\notin\mathbb{Z}$.
Conjecture 2.2.
Let $k\in\{4,6,8,10,14\}$, $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$, and $p$ be an ordinary prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,C})\;\equiv\;u_{p}(C)^{k-2}a_{np^{l-1}}(F_{k,C})\pmod{p^{l}}$,
where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit.
Moreover, if $C/\mathbb{Q}$ has CM, then for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,C})\;\equiv\;u_{p}(C)^{k-2}a_{np^{l-1}}(F_{k,C})\pmod{p^{(k-1)l}}$.
Remark 2.3.
While Theorem 1.6 already presents a connection between meromorphic modular forms and hypergeometric functions, the type of congruences that appear in 2.1 and 2.2 may also remind readers of hypergeometric (super)congruences.
In the case of CM elliptic curves, it is possible to unify the supersingular and ordinary dichotomy via a single, but slightly weaker, three-term ASD congruence. There are two viewpoints in the CM setting: one can interpret everything in terms of CM elliptic curves or in terms of negative discriminants. We will elaborate on both viewpoints here.
Conjecture 2.4.
Let $k\in\{4,6,8,10,14\}$. Let $C/\mathbb{Q}$ be an elliptic curve with CM by some imaginary quadratic field $K$ and suppose that $j(C)\notin\{0,1728\}$. Let $\psi_{C}$ be the Hecke character of $K$ corresponding to $C$ and let $\Theta$ be the Hecke eigenform of weight k1k-1 corresponding to the Hecke character $\psi_{C}^{k-2}$. Let $p$ be a good prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,C})-a_{p}(\Theta){\hskip 1.0pt}a_{np^{l-1}}(F_{k,C})+\chi_{K}(p)p^{k-2}a_{np^{l-2}}(F_{k,C})\;\equiv\;0\pmod{p^{(k-1)l-1}}$,
where $\chi_{K}$ is the quadratic character corresponding to the extension $K/\mathbb{Q}$ and $a_{m}(\;\cdot\;):=0$ if $m\notin\mathbb{Z}$.
Remark 2.5.
By Deuring’s criterion (see [Deu41] or [Sil94, Chapter II, Exercise 2.30]), a good prime $p$ of $C$ is ordinary (resp. supersingular) if and only if $\chi_{K}(p)=1$, i.e., $p$ splits in $K$ (resp. $\chi_{K}(p)=-1$, i.e., $p$ is inert in $K$). The coefficient $a_{p}(\Theta)$ for a good prime $p$ of $C$ can then be written down explicitly as
$a_{p}(\Theta)\;=\;\begin{cases}\alpha_{p}^{k-2}+\beta_{p}^{k-2}&\text{ if }\chi_{K}(p)=1,\\ 0&\text{ if }\chi_{K}(p)=-1,\end{cases}$
where $\alpha_{p}$ and $\beta_{p}$ are the two roots of $X^{2}-a_{p}(C)X+p$ when $\chi_{K}(p)=1$.
From this explicit description, one can readily deduce 2.4 from the supersingular (2.1) and ordinary (2.2) dichotomy. In the supersingular case, i.e., $\chi_{K}(p)=-1$, 2.4 reduces to that
$a_{np^{l}}(F_{k,C})\;\equiv\;p^{k-2}a_{np^{l-2}}(F_{k,C})\pmod{p^{(k-1)l-1}}$,
which follows from 2.1. In the ordinary case, i.e., $\chi_{K}(p)=1$, 2.4 reduces to that
$a_{np^{l}}(F_{k,C})-(\alpha_{p}^{k-2}+\beta_{p}^{k-2})a_{np^{l-1}}(F_{k,C})+p^{k-2}a_{np^{l-2}}(F_{k,C})\;\equiv\;0\pmod{p^{(k-1)l-1}}$.
If we let $u_{p}(C)$ denote the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit, then we can rewrite the equation above as
$\left(a_{np^{l}}(F_{k,C})-u_{p}(C)^{k-2}a_{np^{l-1}}(F_{k,C})\right)$
$\hskip 42.67912pt+\bigg(\frac{p}{u_{p}(C)}\bigg)^{k-2}\left(a_{np^{l-1}}(F_{k,C})-u_{p}(C)^{k-2}a_{np^{l-2}}(F_{k,C})\right)\;\equiv\;0\pmod{p^{(k-1)l-1}}$,
which then follows from 2.2.
Now, we turn to the negative discriminants. Let D<4D<-4 be a discriminant with class number $1$, $\mathcal{O}_{D}$ be the imaginary quadratic order of discriminant $D$, and $\alpha_{D}\in\mathcal{H}$ be a fixed CM point of discriminant $D$. For $k\in\{4,6,8,10,14\}$, write
$F_{k,D}\;:=\;\frac{E_{k}}{j-j(\alpha_{D})}\;=\;\sum_{n=1}^{\infty}a_{n}(F_{k,D})q^{n}.$
Conjecture 2.6.
Let $k\in\{4,6,8,10,14\}$ and let D<4D<-4 be a discriminant with class number $1$. Consider the CM modular form
$\Theta\;:=\;\Theta_{k-1,D}\;:=\;\frac{1}{2}\sum_{\alpha\in\mathcal{O}_{D}}\alpha^{k-2}q^{|\mathrm{N}(\alpha)|}$
of weight k1k-1. Let $p$ be a prime with $p\nmid D$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,D})-a_{p}(\Theta)a_{np^{l-1}}(F_{k,D})+\big(\tfrac{D}{p}\big)p^{k-2}a_{np^{l-2}}(F_{k,D})\;\equiv\;0\pmod{p^{(k-1)l-1}}$,
where $a_{m}(\;\cdot\;):=0$ if $m\notin\mathbb{Z}$.
Remark 2.7.
The coefficients $a_{p}(\Theta)$ for primes $p\nmid D$ can be written down explicitly as
$a_{p}(\Theta)\;=\;\begin{cases}\pi^{k-2}+\overline{\pi}^{k-2}=\mathrm{tr}_{K/\mathbb{Q}}(\pi^{k-2})&\text{ if }\big(\tfrac{D}{p}\big)=1,\\ 0&\text{ if }\big(\tfrac{D}{p}\big)=-1,\end{cases}$
where $K=\mathbb{Q}(\sqrt{D})$ and $p$ with $\big(\frac{D}{p}\big)=1$ factors as $p=\pi\overline{\pi}$ in $\mathcal{O}_{D}$.
One can also write 2.2 in the CM case explicitly in terms of negative discriminants.
Conjecture 2.8.
Let $k\in\{4,6,8,10,14\}$ and let D<4D<-4 be a discriminant with class number $1$. Let $p$ be a prime with $\big(\frac{D}{p}\big)=1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$, and write $p=\pi\overline{\pi}$ for some $\pi\in\mathcal{O}_{D}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,D})$ $\;\equiv\;\pi^{k-2}a_{np^{l-1}}(F_{k,D})\pmod{\overline{\pi}^{(k-1)l}}$.
3. Poles of higher order
As suggested constantly, meromorphic modular forms of weight $k$ with precisely one pole at a non-cuspidal point should behave similarly to the symmetric (k2)(k-2)-nd power of the elliptic curve corresponding to the non-cuspidal pole. The discrepancy, however, is that modular forms typically correspond to $2$-dimensional objects from representation theory, while symmetric (k2)(k-2)-nd powers of elliptic curves are (k1)(k-1)-dimensional. Indeed, on the modular form side, the “correct” space of meromorphic modular forms to consider should be
$\mathrm{MMF}_{k,C}\;:=\;\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(C)},\;\frac{E_{k}}{(j-j(C))^{2}},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}\bigg\}$ (3)
(for $k\in\{4,6,8,10,14\}$ and $j(C)\notin\{0,1728\}$). This is a (k1)(k-1)-dimensional space and should “correspond” to the (k1)(k-1)-dimensional space $\mathrm{Sym}^{k-2}C$ on the elliptic curve side.
The consideration of this particular space is motivated by several reasons. It is easy to check that for $r\geq k$, $\frac{E_{k}}{(j-j(C))^{r}}$ can be written as a linear combination of $\frac{E_{k}}{j-j(C)},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}$, and $\mathcal{D}^{k-1}f$ for some modular form $f$ of weight 2k2-k, where $\mathcal{D}=q\frac{\mathrm{d}}{\mathrm{d}q}$ is the usual derivative. As the congruence properties of $\mathcal{D}^{k-1}f$ are easy to understand  [^5] From the viewpoint of the magnetic property, $\mathcal{D}^{k-1}f$ is trivially (k1)(k-1)-magnetic, and hence not interesting either., it suffices to consider the denominator only up to the (k1)(k-1)-st power. More importantly, the discussion above yields a non-canonical isomorphism (for $k\in\{4,6,8,10,14\}$ and $j(C)\notin\{0,1728\}$)
$\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(C)},\;\frac{E_{k}}{(j-j(C))^{2}},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}\bigg\}\;\cong\;\frac{S_{k}^{\mathrm{mero},C}}{\mathcal{D}^{k-1}M_{2-k}^{\mathrm{mero},C}},$ (4)
where $M_{k}^{\mathrm{mero},C}$ denotes the space of meromorphic modular forms of weight $k$ and level $1$ with precisely one pole at (the point on $\mathcal{H}$ corresponding to) $C$, and $S_{k}^{\mathrm{mero},C}$ denotes its subspace consisting of elements with vanishing constant terms in the Fourier expansions. Here a similar quotient where $C$ is replaced by the cusp $\infty$ occurred in Kazalicki–Scholl [KS16], where they interpreted the quotient cohomologically and proved the ASD congruences of weakly holomorphic modular forms. In our case, a cohomological interpretation of the quotient was discussed in Brown–Fonseca [BF25], which was related to $\mathrm{Sym}^{k-2}C$. This suggests that these meromorphic modular forms should be linked to symmetric powers of elliptic curves via the ASD congruences.
The guiding philosophy in our case is that the action of $U_{p}$ on the modular form side  [^6] Indeed, there is no $U_{p}$ action on the space $\mathrm{MMF}_{k,C}$, since the $U_{p}$ operator increases the level and shifts the pole. Still, we will formally consider the $U_{p}$ operator as defined by $\sum a_{n}q^{n}\mapsto\sum a_{pn}q^{n}$. should correspond to the action of the Frobenius at $p$ on the elliptic curve side.
$\mathrm{Span}\bigg\{\dfrac{E_{k}}{j-j(C)},\;\dfrac{E_{k}}{(j-j(C))^{2}},\ldots,\dfrac{E_{k}}{(j-j(C))^{k-1}}\bigg\}$$\mathrm{Sym}^{k-2}C$$\scriptstyle{U_{p}}$$\scriptstyle{\mathrm{Frob}_{p}}$
Let $p$ be a good prime of $C$ and let $P_{p}(X)$ denote the characteristic polynomial of the Frobenius at $p$ on $\mathrm{Sym}^{k-2}C$. Then,
$P_{p}(X)$ $\;=\;(X-\alpha_{p}^{k-2})(X-\alpha_{p}^{k-3}\beta_{p})\cdots(X-\alpha_{p}\beta_{p}^{k-3})(X-\beta_{p}^{k-2})$
$\;=:\;\sum_{i=0}^{k-1}c_{p,k-1-i}^{(k)}X^{i}\;=\;X^{k-1}+\sum_{i=0}^{k-2}c_{p,k-1-i}^{(k)}X^{i},$
where $\alpha_{p}$ and $\beta_{p}$ are the two roots of $X^{2}-a_{p}(C)X+p$. The expectation is that $F=\frac{E_{k}}{(j-j(C))^{r}}$ should satisfy the following ASD congruence
$a_{np^{l}}(F)+c_{p,1}^{(k)}{\hskip 1.0pt}a_{np^{l-1}}(F)+c_{p,2}^{(k)}{\hskip 1.0pt}a_{np^{l-2}}(F)+\cdots+c_{p,k-1}^{(k)}{\hskip 1.0pt}a_{np^{l-k+1}}(F)\;\equiv\;0\pmod{p^{\bullet}}$
for some power of $p$ with exponent depending on $k,l$, and $r$.
For $r\in\mathbb{Z}^{+}$, write
$F_{k,C}^{(r)}\;:=\;\frac{E_{k}}{(j-j(C))^{r}}\;=\;\sum_{n=1}^{\infty}a_{n}(F_{k,C}^{(r)})q^{n}.$
Then, one has the following conjecture.
Conjecture 3.1.
Let $k\in\{4,6,8,10,14\}$, $1\leq r\leq k-1$, and $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$. Let $p$ be a good prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,C}^{(r)})+c_{p,1}^{(k)}{\hskip 1.0pt}a_{np^{l-1}}(F_{k,C}^{(r)})+\cdots+c_{p,k-1}^{(k)}{\hskip 1.0pt}a_{np^{l-k+1}}(F_{k,C}^{(r)})\;\equiv\;0\pmod{p^{(k-1)l-\frac{1}{2}(k-3)k-r}}$,
where $a_{m}(\;\cdot\;)=0$ if $m\notin\mathbb{Z}$.
Remark 3.2.
If one simply considers the cases when $l\geq k-1$ and writes s=lk+1s=l-k+1, then 3.1 can be rewritten as for all $s\in\mathbb{Z}_{\geq 0}$,
$U_{p}^{s}{\hskip 1.0pt}P_{p}(U_{p})(F_{k,C}^{(r)})\;\equiv\;0\pmod{p^{(k-1)s+\frac{1}{2}(k^{2}-k+2)-r}}$.
On the other hand, one may also consider $V_{p}:\sum a_{n}q^{n}\mapsto\sum a_{n}q^{pn}$ and rewrite 3.1 as for all $l\in\mathbb{Z}^{+}$,
$U_{p}^{l}{\hskip 1.0pt}P_{p}^{*}(V_{p})(F_{k,C}^{(r)})\;\equiv\;0\pmod{p^{(k-1)l-\frac{1}{2}(k-3)k-r}}$,
where $P_{p}^{*}$ is the reciprocal polynomial of $P_{p}$, i.e., $P_{p}^{*}(X)=X^{k-1}P_{p}(X^{-1})$.  [^7] In view of Kazalicki–Scholl [KS16], the formulation using $V_{p}$ may be the “correct” one, since their proof of the ASD congruences involves identifying a $V_{p}$-related operator with the Frobenius at $p$.
Remark 3.3.
Note that $(k-1)l-\frac{1}{2}(k-3)k-r$ is never positive when $l=1$ and $r=1$. It is thus not possible to recover Theorem 1.7 from 3.1 by choosing $l=1$ and $r=1$. This is not the only case where the actual congruences appear stronger than the ones predicted by 3.1. For example, let $k=6$ and assume the same conditions as in 3.1. Then, it appears that the following are true: for all $n\in\mathbb{Z}^{+}$ and $1\leq r\leq 5$,
$a_{np^{2}}(F_{6,C}^{(r)})+c_{p,1}^{(6)}{\hskip 1.0pt}a_{np}(F_{6,C}^{(r)})+c_{p,2}^{(6)}{\hskip 1.0pt}a_{n}(F_{6,C}^{(r)})$  $\;\equiv\;0\pmod{p^{4-r}}$
$a_{np^{3}}(F_{6,C}^{(r)})+c_{p,1}^{(6)}{\hskip 1.0pt}a_{np^{2}}(F_{6,C}^{(r)})+c_{p,2}^{(6)}{\hskip 1.0pt}a_{np}(F_{6,C}^{(r)})+c_{p,3}^{(6)}{\hskip 1.0pt}a_{n}(F_{6,C}^{(r)})$  $\;\equiv\;0\pmod{p^{7-r}}$.
We give an example of 3.1 when $k=6$ and $r=2$, which also illustrates the stronger congruences described in Remark 3.3.
Example 3.1.
Let $k=6$, $r=2$, and $C/\mathbb{Q}$ be the elliptic curve given by $y^{2}+y=x^{3}-x^{2}$ with LMFDB label 11.a3 and j(C)=4096/11j(C)=-4096/11. We will look at
$F\;=\;F_{6,C}^{(2)}\;=\;\frac{E_{6}}{(j+4096/11)^{2}}.$
In this case,
$(k-1)l-\tfrac{1}{2}(k-3)k-r\;=\;5l-11.$
Now, let $p=5$, which is a good prime of $C$ and satisfies that $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. In this case, we have $a_{5}(C)=1$, and one computes the characteristic polynomial of the Frobenius at $5$ on $\mathrm{Sym}^{4}C$ to be
$P_{5}(X)\;=\;X^{5}-11X^{4}-495X^{3}+12375X^{2}+171875X-5^{10}.$
Then, one observes numerically that for all $n\in\mathbb{Z}^{+}$,
$0$ ${\hskip 1.0pt}\equiv{\hskip 1.0pt}a_{5^{2}n}(F)-11{\hskip 1.0pt}a_{5n}(F)-495{\hskip 1.0pt}a_{n}(F)$ $\pmod{5^{2}}$
$0$ ${\hskip 1.0pt}\equiv{\hskip 1.0pt}a_{5^{3}n}(F)-11{\hskip 1.0pt}a_{5^{2}n}(F)-495{\hskip 1.0pt}a_{5n}(F)+12375{\hskip 1.0pt}a_{n}(F)$ $\pmod{5^{5}}$
$0$ ${\hskip 1.0pt}\equiv{\hskip 1.0pt}a_{5^{4}n}(F)-11{\hskip 1.0pt}a_{5^{3}n}(F)-495{\hskip 1.0pt}a_{5^{2}n}(F)+12375{\hskip 1.0pt}a_{5n}(F)+171875{\hskip 1.0pt}a_{n}(F)$ $\pmod{5^{9}}$
$0$ ${\hskip 1.0pt}\equiv{\hskip 1.0pt}a_{5^{5}n}(F)-11{\hskip 1.0pt}a_{5^{4}n}(F)-495{\hskip 1.0pt}a_{5^{3}n}(F)+12375{\hskip 1.0pt}a_{5^{2}n}(F)+171875{\hskip 1.0pt}a_{5n}(F)-5^{10}\cdot a_{n}(F)$ $\pmod{5^{15}}$.
Here the congruences when $l=2,3$ are stronger than the ones predicted by 3.1, as discussed in Remark 3.3; the one when $l=4$ is the same as predicted by 3.1; the one when $l=5$ is stronger again, but for a general prime, it does seem that the one when $l=5$ is usually the same as predicted by 3.1.
If one separates the supersingular and ordinary primes as before, then one also observes a generalization of 2.1. However, generalizations of 2.2 seem to only exist in the CM case, which will be discussed in Section 4.
Conjecture 3.4.
Let $k\in\{4,6,8,10,14\}$, $1\leq r\leq k-1$, and $C/\mathbb{Q}$ be an elliptic curve. Let $p$ be a supersingular prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{k,C}^{(r)})\;\equiv\;p^{k-2}a_{np^{l-2}}(F_{k,C}^{(r)})\pmod{p^{(k-1)l-r}}$,
where $a_{m}(\;\cdot\;):=0$ if $m\notin\mathbb{Z}$.
Remark 3.5.
When $p$ is supersingular, $F_{k,C}^{(r)}$ may be viewed as an overconvergent $p$-adic modular form, since the pole lies in the supersingular locus. Indeed, for $l\geq 2$, 3.4 leads to the consideration of $U_{p}^{l-2}(U_{p}-p^{k-2})(F_{k,C}^{(r)})$. Here one observes that there is no old cusp form by our choice of $k$ and the operator $U_{p}-p^{k-2}$ annihilates all the newforms of weight $k$ and level $p$. Hence, heuristically one should be able to write $(U_{p}-p^{k-2})(F_{k,C}^{(r)})$ as an asymptotic sum of $U_{p}$-eigenforms  [^8] We refer to [Cal, Section 3.5] and [GM95] for discussions on the asymptotic expansions of overconvergent modular forms using $U_{p}$-eigenforms. with slopes $\geq k-1$, which then also supports the conjectural congruence satisfied by $U_{p}^{l-2}(U_{p}-p^{k-2})(F_{k,C}^{(r)})$.
Remark 3.6.
3.4 implies 3.1 in the supersingular case. We sketch a proof here for $l\geq k-1$. For a supersingular prime $p$, the characteristic polynomial $P_{p}(X)$ is given by
$P_{p}(X)\;=\;$ $\prod_{i=0}^{k-2}\big(X-(\sqrt{-p})^{k-2-i}(-\sqrt{-p})^{i}\big)$
$\;=\;$ $(X-(-p)^{\frac{k-2}{2}})(X^{2}-p^{k-2})^{\frac{k-2}{2}}$
$\;=\;$ $\bigg(\sum_{i=0}^{\frac{k-4}{2}}\binom{\frac{k-4}{2}}{i}(-p^{k-2})^{i}X^{k-3-2i}-(-p)^{\frac{k-2}{2}}\sum_{i=0}^{\frac{k-4}{2}}\binom{\frac{k-4}{2}}{i}(-p^{k-2})^{i}X^{k-4-2i}\bigg)(X^{2}-p^{k-2}).$
It then follows from 3.4 and direct computations that for all $s\in\mathbb{Z}_{\geq 0}$,
$U_{p}^{s}{\hskip 1.0pt}P_{p}(U_{p})(F_{k,C}^{(r)})\;\equiv\;0\pmod{p^{(k-1)s+\frac{1}{2}(k^{2}-k+2)-r}}$,
which implies 3.1.
We give an example of 3.4 when $k=4$ and $l=1$, In this case, one expects that
$a_{np}(F_{4,C}^{(1)})\;\equiv\;0\pmod{p^{2}}\quad\text{ and }\quad a_{np}(F_{4,C}^{(2)})\;\equiv\;0\pmod{p}$
for all $n\in\mathbb{Z}^{+}$, or equivalently,
$U_{p}F_{4,C}^{(1)}\;\equiv\;0\pmod{p^{2}}\quad\text{ and }\quad U_{p}F_{4,C}^{(2)}\;\equiv\;0\pmod{p}$.
This example is also related to the heuristic described in Remark 3.5.
Example 3.2.
Let $k=4$. Let $p=13$ and let $C/\mathbb{Q}$ be the elliptic curve given by $y^{2}=x^{3}-x^{2}-4$ with LMFDB label 56.b2. In particular, $C$ is supersingular at $13$ and j(C)=4/7j(C)=-4/7 satisfies that $v_{13}(j(C))=0=v_{13}(j(C)-1728)$. We will now try to write
$\frac{E_{4}}{j+4/7}\pmod{13^{2}}\quad\text{ and }\quad\frac{E_{4}}{(j+4/7)^{2}}\pmod{13}$
in terms of linear combinations of Hecke eigenforms of level $13$ and weight $4$.
Note that there are exactly three Hecke eigenforms in $S_{4}(\Gamma_{0}(13))$: the one $f_{1}$ with rational coefficients and with LMFDB label 13.4.a.a, and two embeddings $f_{2},f_{3}$ of the one with coefficients in $\mathbb{Q}(\sqrt{17})$ and with LMFDB label 13.4.a.b. In particular,
$U_{13}f_{1}\;=\;-13f_{1},\quad U_{13}f_{2}\;=\;13f_{2},\quad U_{13}f_{3}\;=\;13f_{3},$
and
$f_{1}$ $\;\equiv\;q-5q^{2}-7q^{3}+17q^{4}+O(q^{5})$ $\pmod{13^{2}}$
$f_{2}$ $\;\equiv\;q-7q^{2}+25q^{3}-11q^{4}+O(q^{5})$ $\pmod{13^{2}}$
$f_{3}$ $\;\equiv\;q-5q^{2}+19q^{3}-9q^{4}+O(q^{5})$ $\pmod{13^{2}}$.
Moreover, it is easy to check that
$f_{1}\;\equiv\;f_{3}\pmod{13}$.
For $r=1$, we find that
$\frac{E_{4}}{j+4/7}\;\equiv\;q+99q^{2}+32q^{3}+108q^{4}+O(q^{5})\;\equiv\;72f_{1}+117f_{2}+150f_{3}\pmod{13^{2}}$. (5)
Hence,
$U_{13}\bigg(\frac{E_{4}}{j+4/7}\bigg)\;\equiv\;13\cdot(-72f_{1}+117f_{2}+150f_{3})\;\equiv\;13\cdot(78f_{1}+117f_{2})\;\equiv\;0\pmod{13^{2}}$.
For $r=2$, we find that
$\frac{E_{4}}{(j+4/7)^{2}}\;\equiv\;q^{2}+10q^{3}+q^{4}+O(q^{5})\;\equiv\;7f_{1}+6f_{2}\pmod{13}$. (6)
Hence,
$U_{13}\bigg(\frac{E_{4}}{(j+4/7)^{2}}\bigg)\;\equiv\;13\cdot(-7f_{1}+6f_{2})\;\equiv\;0\pmod{13}$.
Here the congruences (5) and (6) are provable by multiplying both sides by $j+4/7$ or $(j+4/7)^{2}$ and comparing the Fourier coefficients up to a suitable number of terms, so we in fact obtain a proof of 3.4 in this special case.
We finish this section with two comments. The first comment is that one should be able to generalize the whole setting to elliptic curves over number fields. In this case, one should then consider congruences modulo powers of some fixed prime ideal $\mathfrak{p}$, and the action of the Frobenius at $\mathfrak{p}$ on the elliptic curve side should correspond to the action of $U_{\mathrm{N}(\mathfrak{p})}$ on the modular form side, where $\mathrm{N}(\mathfrak{p})$ denotes the norm. The second comment is that one should expect a motivic interpretation of the connection between meromorphic modular forms and symmetric powers of elliptic curves  [^9] This is also why we always write $\mathrm{Sym}^{k-2}C$ instead of $\mathrm{Sym}^{k-2}H^{1}(C)$ with some specified cohomology, since we would like to think of $\mathrm{Sym}^{k-2}C$ as being the motive attached to the symmetric power of the elliptic curve., as discussed in Brown–Fonseca [BF25]. In particular, the congruences that we observe here should be viewed as visualizations of the motivic connection at finite places, and it may be reasonable to expect visualizations at infinite places linking special values of $L$-functions of symmetric powers of elliptic curves to “periods” of meromorphic modular forms.
4. The CM case, supercongruences, and magnetic modular forms
We will now specialize to the CM case. Let D<4D<-4 be a discriminant with class number $1$, $\mathcal{O}_{D}$ be the imaginary quadratic order of discriminant $D$, $\alpha_{D}\in\mathcal{H}$ be a fixed CM point with discriminant $D$. Let $C/\mathbb{Q}$ be an elliptic curve with CM by $\mathcal{O}_{D}$ so that $j(C)=j(\alpha_{D})$ and let $\psi_{C}$ be the corresponding Hecke character of $K:=\mathbb{Q}(\sqrt{D})$. For the convenience of the later discussion, we will choose $C/\mathbb{Q}$ in a way that all of its bad primes divide $D$. By Deuring’s criterion, the following are equivalent for a prime $p$ with $p\nmid D$:
  1. (i)
    $p$ is an ordinary prime of $C$ (resp. $p$ is a supersingular prime of $C$);
  2. (ii)
    $p$ splits in $\mathcal{O}_{D}$ (resp. $p$ is inert in $\mathcal{O}_{D}$);
  3. (iii)
    $(\frac{D}{p})=1$ (resp. $(\frac{D}{p})=-1$).
We will thus use these three conditions interchangeably.
4.1. A “canonical” basis
As mentioned briefly in the Section 1.3, we will construct a basis for the space
$\mathrm{MMF}_{k,D}\;:=\;\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(\alpha_{D})},\;\frac{E_{k}}{(j-j(\alpha_{D}))^{2}},\ldots,\frac{E_{k}}{(j-j(\alpha_{D}))^{k-1}}\bigg\}$ (7)
that is diagonal (resp. anti-diagonal) with respect to the action of (powers of) $U_{p}$ via explicit supercongruences for all ordinary (resp. supersingular) primes $p$, where the (anti-)eigenvalues will be described explicitly using the $p$-adic analogue of the Chowla–Selberg periods.
To motivate the construction, we first focus on the ordinary case. Note that on the elliptic curve side, we have the following splitting (as Galois representations of $K$)
$\mathrm{Res}_{K/\mathbb{Q}}(\mathrm{Sym}^{k-2}C)\;\cong\;\mathrm{Res}_{K/\mathbb{Q}}\big(\mathrm{Sym}^{k-2}(\mathrm{Ind}_{K/\mathbb{Q}}\psi_{C})\big)\;\cong\;\bigoplus_{\begin{subarray}{c}a+b=k-2\\ 0\leq a,b\leq k-2\end{subarray}}\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}.$ (8)
Let $p$ be a prime that splits in $\mathcal{O}_{D}$. Then, for any embedding $\pi:K\hookrightarrow\mathbb{Q}_{p}$, $\mathrm{Frob}_{\pi}(=\mathrm{Frob}_{p})$ acts on $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ by $p^{b}u_{p}(C)^{a-b}$, where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit under the embedding $\pi$. [^10] Technically, $\mathrm{Frob}_{\pi}$ acts on one of $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ and $\psi_{C}^{b}{\hskip 1.0pt}\overline{\psi_{C}^{a}}$ by $p^{b}u_{p}(C)^{a-b}$ depending on the choice of $\psi_{C}$. We will thus replace $\psi_{C}$ by $\overline{\psi_{C}}$ if necessary to ensure that $\mathrm{Frob}_{\pi}$ acts on $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ by $p^{b}u_{p}(C)^{a-b}$. This suggests that on the modular form side, the space $\mathrm{MMF}_{k,D}$ should also split into $1$-dimensional pieces $\{V_{a,b}\}_{a+b=k-2}$ equipped with corresponding $U_{p}$-actions. In other words, there should exist $\{G_{a,b}\}_{a+b=k-2}$ such that
  1. (1)
    $\mathrm{Span}\{G_{a,b}\mid a+b=k-2\}=\mathrm{MMF}_{k,D}$;
  2. (2)
    for all (but finitely many) primes $p$ with $(\frac{D}{p})=1$, $U_{p}G_{a,b}\equiv p^{b}u_{p}(C)^{a-b}G_{a,b}$ modulo some power of $p$, where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit.
In order to construct these $G_{a,b}$, the idea is to view the original $\frac{E_{k}}{j-j(C)}$ as a two variable function $\frac{E_{k}(z)}{j(z)-j(\tau)}$, multiply it by some auxiliary function $f^{\mathrm{aux}}(\tau)$, differentiate the product (several times) with respect to the auxiliary variable $\tau$, and evaluate $\tau$ at the CM point $\alpha_{D}$.
In this case, the choice of the auxiliary function was already revealed in Section 1.2. That is, for $k\in\{4,6,8,10,14\}$, we choose the auxiliary function as $f_{k}^{\mathrm{aux}}(\tau)=\frac{E_{14-k}}{\Delta}(\tau)$ and consider
$G_{k}(z,\tau)\;:=\;\frac{E_{k}(z)}{j(z)-j(\tau)}\cdot f_{k}^{\mathrm{aux}}(\tau)\;=\;\frac{E_{k}(z)\cdot\frac{E_{14-k}}{\Delta}(\tau)}{j(z)-j(\tau)}.$ (9)
We also recall from Section 1.2 that the non-holomorphic derivative of a (nearly holomorphic) modular form $f(\tau)$ of weight $k_{f}$ is defined as
$(\partial_{\tau}f)(\tau)\;:=\;\frac{1}{2\pi i}\frac{\partial f}{\partial\tau}-\frac{k_{f}}{4\pi\operatorname{Im}(\tau)}\cdot f(\tau).$
Now, if one naively considers the derivative $(\partial_{\tau}^{r}G_{k})(z,\tau)$ and evaluates $\tau$ at $\alpha_{D}$, then the resulting modular form will depend on the choice of $\alpha_{D}$, since $(\partial_{\tau}^{r}G_{k})(z,\tau)$ is of weight 2k+2r2-k+2r with respect to $\tau$. In order to remove the dependency, we adopt the following normalization.
For $1\leq r\leq k-1$, write
$G_{k,D}^{(r)}(z)\;:=\;\bigg(\frac{\Delta}{E_{14-k}}(\tau)\cdot\bigg(-\frac{E_{4}}{E_{6}}(\tau)\bigg)^{r-1}\cdot(\partial_{\tau}^{r-1}G_{k})(z,\tau)\bigg)\bigg|_{\tau=\alpha_{D}}.$ (10)
In this way, $G_{k,D}^{(r)}$ is independent of the choice of $\alpha_{D}$, has rational coefficients  [^11] The rationality is because $D$ is assumed to have class number $1$ (see also Example 4.1 and Example 4.3)., and is of the form
$G_{k,D}^{(r)}\;=\;\sum_{i=1}^{r-1}A_{i}\cdot\frac{E_{k}}{(j-j(\alpha_{D}))^{i}}+(r-1)!\cdot j(\alpha_{D})^{r-1}\cdot\frac{E_{k}}{(j-j(\alpha_{D}))^{r}}.$ (11)
Through the following conjecture, one observes that each $G_{k,D}^{(r)}$ takes the role of $G_{k-1-r,r-1}$. The conjecture should also be viewed as a natural generalization of 2.2 in the CM case.
Conjecture 4.1.
Let $k\in\{4,6,8,10,14\}$ and let $1\leq r\leq k-1$. Let D<4D<-4 be a discriminant with class number $1$ and let $C/\mathbb{Q}$ be an elliptic curve with $j(C)=j(\alpha_{D})$. Let $p$ be a prime with $(\frac{D}{p})=1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})\;\equiv\;p^{r-1}u_{p}(C)^{k-2r}a_{np^{l-1}}(G_{k,D}^{(r)})\pmod{p^{(k-1)l}}$,
where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit.
Remark 4.2.
It is possible to rewrite this purely in terms of negative discriminants as in 2.8. Since $(\frac{D}{p})=1$, we have that $p=\pi\overline{\pi}$ for some $\pi\in\mathcal{O}_{D}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})\;\equiv\;\pi^{k-1-r}{\hskip 1.0pt}\overline{\pi}^{r-1}a_{np^{l-1}}(G_{k,D}^{(r)})\pmod{\overline{\pi}^{(k-1)l}}$.
Remark 4.3.
As with the case of 3.4, it is also possible to deduce the ASD congruence (3.1) in the ordinary and CM case from 4.1 through involved computations which we omit here.
Remark 4.4.
One can also “explain” 3.4 in this CM case through the same heuristic. If $p=\mathfrak{p}$ is inert in $\mathcal{O}_{D}$, then $\mathrm{Frob}_{\mathfrak{p}}(=\mathrm{Frob}_{p}^{2})$ acts on $\psi_{C}^{a}{\hskip 1.0pt}\overline{\psi_{C}^{b}}$ by $p^{k-2}$. Hence, for the potential basis $\{G_{a,b}\}_{a+b=k-2}$, one would expect that $U_{p}^{2}G_{a,b}\equiv p^{k-2}G_{a,b}$ modulo some power of $p$ for supersingular primes $p$, as observed in 3.4.
Now we turn to the supersingular primes. In this case, we can already inherit certain property from 3.4. However, through the help of the $p$-adic Gamma function, one can moreover observe that for a supersingular prime $p$, the family $\{G_{k,D}^{(r)}\}$ behaves like an anti-diagonal basis with respect to the $U_{p}$ operator, i.e., $U_{p}G_{k,D}^{(r)}$ is a multiple of $G_{k,D}^{(k-r)}$ modulo some power of $p$. Indeed, such phenomenon was already hinted at in [ALS25, Corollary 1.2 and Example 3].
For a fundamental discriminant $D<0$ and a prime $p$ with $p\nmid D$, define
$\Omega_{p,D}\;:=\;-\bigg(\prod_{a=1}^{|D|-1}\Gamma_{p}\bigg(\frac{a}{|D|}\bigg)^{(\frac{D}{a})}\bigg)^{\frac{w_{D}}{2}}\;\in\;\mathbb{Z}_{p}^{\times},$ (12)
where $\Gamma_{p}$ denotes the $p$-adic Gamma function, $(\frac{D}{a})$ denotes the Kronecker symbol, and $w_{D}=|\mathcal{O}_{D}^{\times}|$. One should view this definition as a $p$-adic analogue of (the square of) the Chowla–Selberg period [CS67] (see also Remark 4.10).
Conjecture 4.5.
Let $k\in\{4,6,8,10,14\}$ and let $1\leq r\leq k-1$. Let $D<0$ be a discriminant with class number $1$ and write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}$ and some fundamental discriminant $D_{0}<0$. Suppose that $D\notin\{-3,-4,-12\}$. Let $p$ be a prime with $(\frac{D}{p})=-1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$(r-1)!\cdot a_{np^{l}}(G_{k,D}^{(k-r)})\;\equiv\;(k-1-r)!{\hskip 1.0pt}c_{D}^{\frac{k}{2}-r}\cdot(-p)^{r-1}\Omega_{p,D_{0}}^{-(\frac{k}{2}-r)}a_{np^{l-1}}(G_{k,D}^{(r)})\pmod{p^{(k-1)l}}$,
where we write $j_{D}=j(\alpha_{D})$ and
$c_{D}\;=\;\begin{cases}(-1)^{D}\cdot|j_{D}|^{2/3}\cdot(j_{D}-1728)^{-1}\cdot D\;(\in\mathbb{Q})&\text{if }D\neq-16,-27,\\ (-1)^{D}\cdot|j_{D}|^{2/3}\cdot\big(2(j_{D}-1728)\big)^{-1}\cdot D\;=\;-\frac{484}{3969}&\text{if }D=-16,\\ (-1)^{D}\cdot|\frac{j_{D}}{3}|^{2/3}\cdot(j_{D}-1728)^{-1}\cdot D\;=\;-\frac{3600}{64009}&\text{if }D=-27.\end{cases}$
Remark 4.6.
The value of $c_{D}$ is in fact always a rational square in all the cases we consider here.
$D$ 7-7 8-8 11-11 16-16 19-19 27-27 28-28 43-43 67-67 163-163
$c_{D}$ $(5/9)^{2}$ $(5/7)^{2}$ $(4/7)^{2}$ $(22/63)^{2}$ $(4/9)^{2}$ $(60/253)^{2}$ $(170/513)^{2}$ $(40/189)^{2}$ $(220/1953)^{2}$ $(26680/1672209)^{2}$
Table 1. The value of $c_{D}$.
It is also possible to describe the $p$-adic behaviors of $\{G_{k,D}^{(r)}\}$ for ordinary primes $p$ using the $p$-adic Gamma function, though we would need to exclude the case when D=12D=-12 as before.
Conjecture 4.7.
Let $k\in\{4,6,8,10,14\}$ and let $1\leq r\leq k-1$. Let $D<0$ be a discriminant with class number $1$ and write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}$ and some fundamental discriminant $D_{0}<0$. Suppose that $D\notin\{-3,-4,-12\}$. Let $p$ be a prime with $(\frac{D}{p})=1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})\;\equiv\;p^{r-1}\Omega_{p,D_{0}}^{\frac{k}{2}-r}a_{np^{l-1}}(G_{k,D}^{(r)})\pmod{p^{(k-1)l}}$.
Remark 4.8.
If one renormalizes $G_{k,D}^{(r)}$ by setting
$\widehat{G}_{k,D}^{(r)}\;:=\;(k-1-r)!c_{D}^{-\frac{r-1}{2}}\cdot G_{k,D}^{(r)}$
and renumbers them by setting $H_{k,D}^{(i)}:=\widehat{G}_{k,D}^{(\frac{k}{2}+i)}$ for $|i|\leq\frac{k}{2}-1$, then one can write 4.5 and 4.7 in the following uniform way: for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(H_{k,D}^{(\epsilon_{p}i)})\;\equiv\;\big(\epsilon_{p}p)^{\frac{k}{2}+i-1}\Omega_{p,D_{0}}^{-\epsilon_{p}i}{\hskip 1.0pt}a_{np^{l-1}}(H_{k,D}^{(i)})\pmod{p^{(k-1)l}}$,
where $\epsilon_{p}=(\frac{D}{p})$. While the issue here is that $c_{D}$ is not a square, a naive workaround is to simply view the congruence as describing that the difference of both sides has $p$-adic valuation $\geq(k-1)l$.
Remark 4.9.
Let $D<0$ be a fundamental discriminant with class number $1$. By a result of Gross–Koblitz [GK79, Theorem 4.13], one can give an explicit characterization of $\Omega_{p,D}$ when $(\frac{D}{p})=1$. Suppose that $p$ splits as $p=\pi\overline{\pi}$ in $\mathcal{O}_{D}$ and let $\iota_{\pi}:\mathcal{O}_{D}\rightarrow\mathbb{Z}_{p}$ be the $p$-adic embedding induced by $\pi$. Then,
$\Omega_{p,D}\;=\;\iota_{\pi}(\overline{\pi}^{2}).$
Equivalently, if one chooses an elliptic curve $C/\mathbb{Q}$ with CM by $\mathcal{O}_{D}$ in a way that all of its bad primes divide $D$, then
$\Omega_{p,D}\;=\;u_{p}(C)^{2},$
where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit. In this way, 4.7 is equivalent to 4.1 when $D$ is fundamental.
When $D$ is non-fundamental, i.e., D=16,27,28D=-16,-27,-28 in our case, one observes that the elliptic curve $C_{D}/\mathbb{Q}$ with CM by $\mathcal{O}_{D}$ such that all of its bad primes divide $D$ is isogenous to the one $C_{D_{0}}/\mathbb{Q}$ with CM by $\mathcal{O}_{D_{0}}$ such that all of its bad primes divide $D_{0}$, where $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}$ and some fundamental discriminant $D_{0}<0$. In this way, it follows that
$\Omega_{p,D_{0}}\;=\;u_{p}(C_{D_{0}})^{2}\;=\;u_{p}(C_{D})^{2}.$
Hence, 4.7 is also equivalent to 4.1 in the non-fundamental case.
Remark 4.10.
One can also relate $\Omega_{p,D}$ with the special value of a $p$-adic $L$-function by a result of Ferrero–Greenberg [FG79, Theorem 1]. Let $D<0$ be a fundamental discriminant and fix a prime $p$ with $p\nmid D$. Let $\chi=(\frac{D}{\cdot})$ be the (primitive) Dirichlet character associated to $D$ and let $\varepsilon_{p}$ denote the $p$-adic cyclotomic character. Then,
$L_{p}^{\prime}(0,\chi\varepsilon_{p})\;=\;\sum_{a=1}^{|D|-1}\chi(a)\log_{p}\Gamma_{p}\bigg(\frac{a}{|D|}\bigg)-L_{p}(0,\chi\varepsilon_{p})\cdot\log_{p}|D|,$ (13)
and hence
$L_{p}^{\prime}(0,\chi\varepsilon_{p})\;=\;\frac{2}{w_{D}}\log_{p}\Omega_{p,D}-L_{p}(0,\chi\varepsilon_{p})\cdot\log_{p}|D|,$
where $L_{p}$ denotes the $p$-adic $L$-function and $\log_{p}$ denotes the $p$-adic logarithm. Here $L_{p}(0,\chi\varepsilon_{p})$ can be expressed explicitly as the following
$L_{p}(0,\chi\varepsilon_{p})\;=\;-\big(1-\chi(p)\big)\cdot\frac{1}{|D|}\sum_{a=1}^{|D|-1}\chi(a)a.$
The formula (13) should be compared with the version of the complex $L$-functions. In this case, Lerch’s formula [Ler97, p. 303, (25)] says that
$L^{\prime}(0,\chi)$ $\;=\;\sum_{a=1}^{|D|-1}\chi(a)\log\Gamma\bigg(\frac{a}{|D|}\bigg)-L(0,\chi)\cdot\log|D|.$
4.2. The magnetic property and other observations
In fact, one can already observe interesting numerical phenomena for the family $\{G_{k,D}^{(r)}\}$ by viewing $\mathrm{Sym}^{k-2}C$ simply as a Galois representation of $\mathbb{Q}$. Indeed, we have the following splitting (as Galois representations of $\mathbb{Q}$) [^12] A possibly easier way of observing this splitting is through the splitting of the $L$-function of $\mathrm{Sym}^{k-2}C$.
$\mathrm{Sym}^{k-2}C\;\cong\;\big((\tfrac{D}{\cdot}){\hskip 1.0pt}\varepsilon_{\ell}\big)^{\frac{k-2}{2}}\oplus{\hskip 1.0pt}\bigoplus_{r=1}^{\frac{k-2}{2}}\big(\varepsilon_{\ell}^{r-1}\otimes\mathrm{Ind}_{K/\mathbb{Q}}\psi_{C}^{k-2r}\big),$ (14)
where $(\frac{D}{\cdot})$ denotes (the Dirichlet character given by) the Kronecker symbol and $\varepsilon_{\ell}$ denotes the $\ell$-adic cyclotomic character. Here the term $\big((\tfrac{D}{\cdot}){\hskip 1.0pt}\varepsilon_{\ell}\big)^{\frac{k-2}{2}}$ should correspond to $G_{k,D}^{(k/2)}$ and the term $\varepsilon_{\ell}^{r-1}\otimes\mathrm{Ind}_{K/\mathbb{Q}}\psi_{C}^{k-2r}$ should correspond to $\mathrm{Span}\{G_{k,D}^{(r)},G_{k,D}^{(k-r)}\}$. We also observe that the appearance of powers of $\varepsilon_{\ell}$ seems to account for the magnetic property of $G_{k,D}^{(r)}$. Specifically, we have the following.
Conjecture 4.11.
Let $k\in\{4,6,8,10,14\}$ and let $1\leq r\leq k-1$. Let D<4D<-4 be a discriminant with class number $1$. Let $A=A_{k,D}^{(r)}$ be a nonzero integer such that $\widetilde{G}_{k,D}^{(r)}=AG_{k,D}^{(r)}$ has integer coefficients with no common divisors. Then, $\widetilde{G}_{k,D}^{(r)}$ is $(r^{\prime}-1)$-magnetic, i.e., for all $n\in\mathbb{Z}^{+}$,
$n^{r^{\prime}-1}{\hskip 1.0pt}\big|{\hskip 1.0pt}a_{n}(\widetilde{G}_{k,D}^{(r)}),$
where $r^{\prime}=\min(r,k-r)$.
Remark 4.12.
If one only focuses on $a_{p}=a_{p}(G_{k,D}^{(r)})$, then generically one obtains the following table from 3.4 (or 4.5) and 4.1.
$r$ 11 22 33 $\cdots$ k2k-2 k1k-1
$v_{p}(a_{p})$ for $(\frac{D}{p})=-1$ k2k-2 k3k-3 k4k-4 $\cdots$ 11 0
$v_{p}(a_{p})$ for $(\frac{D}{p})=1$ 0 11 22 $\cdots$ k3k-3 k2k-2
To some extent, this provides a visual explanation of the $(r^{\prime}-1)$-magnetic property of $G_{k,D}^{(r)}$.
Theorem 4.13.
Let $k\in\{4,6,8,10,14\}$ and let D<4D<-4 be a discriminant with class number $1$. Let $p$ be a prime with $p\nmid D$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, $G_{k,D}^{(k/2)}$ satisfies that for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(k/2)})\;\equiv\;\big(\big(\tfrac{D}{p}\big){\hskip 1.0pt}p\big)^{\frac{k-2}{2}}a_{np^{l-1}}(G_{k,D}^{(k/2)})\pmod{p^{(k-1)l}}$.
Remark 4.14.
We include this result here for the completeness of the discussion. It simply follows from Theorem 1.10, since $G_{k,D}^{(k/2)}$ is a nonzero rational multiple of $\widetilde{\mathcal{G}}_{k,D}$, and the rational multiple involves only divisors of $D$, $j(\alpha_{D})$, and $j(\alpha_{D})-1728$.
Conjecture 4.15.
Let $k\in\{4,6,8,10,14\}$, $1\leq r\leq k-1$ with $r\neq\frac{k}{2}$, and $r^{\prime}=\min(r,k-r)$. Let D<4D<-4 be a discriminant with class number $1$. Consider the CM modular form
$\Theta_{r^{\prime}}\;:=\;\Theta_{k+1-2r^{\prime},D}\;:=\;\frac{1}{2}\sum_{\alpha\in\mathcal{O}_{D}}\alpha^{k-2r^{\prime}}q^{|\mathrm{N}(\alpha)|}$
of weight $k+1-2r^{\prime}$. Let $p$ be a prime with $p\nmid D$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})-p^{r^{\prime}-1}a_{p}(\Theta_{r^{\prime}})a_{np^{l-1}}(G_{k,D}^{(r)})+\big(\tfrac{D}{p}\big)p^{k-2}a_{np^{l-2}}(G_{k,D}^{(r)})\;\equiv\;0\pmod{p^{(k-1)l-r}}$.
Remark 4.16.
This should be viewed as a generalization of the three-term ASD congruence observed in the case of CM elliptic curves and simple poles, i.e., 2.4 or 2.6. As in Remark 2.5, this conjecture also follows from the supersingular (3.4) and ordinary (4.1) dichotomy.
The last two observations are related to tensor products. In the splitting (8) of the elliptic curve side over $K$, one has that $\psi_{C}^{a}\overline{\psi_{C}^{b}}\otimes\psi_{C}^{b}\overline{\psi_{C}^{a}}=(\psi_{C}\overline{\psi_{C}})^{k-2}=(\mathrm{Res}_{K/\mathbb{Q}}\varepsilon_{\ell})^{k-2}$. On the modular form side, this suggests that the “tensor product” of $G_{k,D}^{(r)}$ and $G_{k,D}^{(k-r)}$ should resemble the character $\varepsilon_{\ell}^{k-2}$. Here we form the “tensor product” of two modular forms via termwise multiplication, motivated by the Rankin–Selberg theory.
Conjecture 4.17.
Let $k\in\{4,6,8,10,14\}$, $r\in\mathbb{Z}$ with $1\leq r\leq k-1$, and $r^{\prime}=\min(r,k-r)$. Let D<4D<-4 be a discriminant with class number $1$. Let $p$ be a prime with $p\nmid D$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)})a_{np^{l}}(G_{k,D}^{(k-r)})\;\equiv\;p^{k-2}a_{np^{l-1}}(G_{k,D}^{(r)})a_{np^{l-1}}(G_{k,D}^{(k-r)})\pmod{p^{(k+r^{\prime}-2)l}}$. (15)
Remark 4.18.
When generalizing this conjecture to other weights in Section 5, it is possible that in the case when $(\frac{D}{p})=-1$, one may need to multiply both sides by $(r-1)!\cdot(k-1-r)!$ to account for the factors that appear in 4.5. Also, when $(\frac{D}{p})=-1$, the numerical observations suggest that the congruence (15) holds modulo $p^{(k+r^{\prime}-2)l+(k-2r^{\prime})}$ for $l\geq 2$.
Conjecture 4.19.
Let $k\in\{4,6,8,10,14\}$ and $1\leq r\leq k-1$. Let D<4D<-4 be a discriminant with class number $1$. Then, there exists a nonzero integer $A$ such that for all $n\in\mathbb{Z}^{+}$,
$n^{k-2}{\hskip 1.0pt}\big|{\hskip 1.0pt}A{\hskip 1.0pt}a_{n}(G_{k,D}^{(r)})a_{n}(G_{k,D}^{(k-r)}).$
4.3. Examples
We will now discuss several examples in weight $4$ and $6$. We will mostly focus on 4.1 and 4.5, since other phenomena are more or less consequences of these two conjectures. Throughout this subsection, we define
$\psi\;:=\;\frac{E_{2}^{*}E_{4}}{E_{6}}\quad\text{ with }\quad E_{2}^{*}(\tau)\;:=\;E_{2}(\tau)-\frac{3}{\pi\operatorname{Im}(\tau)}.$
Here $E_{2}^{*}$ is called the non-holomorphic Eisenstein series of weight $2$, which satisfies the modular transformation of weight $2$. As a consequence, $\psi$ is invariant under the modular transformation, and hence is a (non-holomorphic) modular function. For a discriminant $D<0$ of class number $1$, we write
$j_{D}\;=\;j(\alpha_{D})\quad\text{ and }\quad\psi_{D}\;=\;\psi(\alpha_{D}).$
It follows from a result by Masser [Mas75, Theorem A1] that $\psi_{D}\in\mathbb{Q}$. [^13] In general, Masser showed that $\psi(\alpha)\in\mathbb{Q}(j(\alpha))$ for any CM point $\alpha$ on the upper half plane. Here we tabulate the exact values of $j_{D}$ and $\psi_{D}$ and we refer to [Cox13, §12] and [Mas75, p. 121] (see also [Mil22]) for the proofs of these exact values.
$D$ 7-7 8-8 11-11 12-12 16-16 19-19 27-27 28-28 43-43 67-67 163-163
$j_{D}$ $15^{3}$ $20^{3}$ $32^{3}$ $2\cdot 30^{3}$ $66^{3}$ $96^{3}$ $3\cdot 160^{3}$ $255^{3}$ $960^{3}$ $5280^{3}$ $640320^{3}$
$\psi_{D}$ $5/21$ $5/14$ $32/77$ $5/11$ $11/21$ $32/57$ $160/253$ $85/133$ $640/903$ $33440/43617$ $77265280/90856689$
Table 2. The values of $j_{D}$ and $\psi_{D}$.
Example 4.1.
Let $k=4$. By computations, we have
$G_{4,D}^{(1)}$ $\;=\;\frac{E_{4}}{j-j_{D}}$
$G_{4,D}^{(2)}$ $\;=\;\frac{E_{4}}{j-j_{D}}\cdot\left(\tfrac{1}{2}\cdot\tfrac{j_{D}}{j_{D}-1728}+\big(\tfrac{1}{3}+\tfrac{\psi_{D}}{6}\big)\right)+\frac{E_{4}}{(j-j_{D})^{2}}\cdot j_{D}$
$G_{4,D}^{(3)}$ $\;=\;\tfrac{E_{4}}{j-j_{D}}\cdot{\scriptstyle\left(\big(\tfrac{61}{72}+\tfrac{\psi_{D}}{12}\big)\cdot\tfrac{j_{D}}{j_{D}-1728}+\big(\tfrac{\psi_{D}}{18}+\tfrac{\psi_{D}^{2}}{72}\big)\right)}+\tfrac{E_{4}}{(j-j_{D})^{2}}\cdot{\scriptstyle j_{D}\left(\tfrac{3}{2}\cdot\tfrac{j_{D}}{j_{D}-1728}+\big(\tfrac{4}{3}+\tfrac{\psi_{D}}{6}\big)\right)}+\tfrac{E_{4}}{(j-j_{D})^{3}}\cdot 2j_{D}^{2}.$
Example 4.2.
Let $k=4$ and let D=7D=-7 so that $j_{D}=-3375$ and $\psi_{D}=5/21$. By Example 4.1, we have
$G_{4,-7}^{(1)}$ $\;=\;\frac{E_{4}}{j+3375}$
$G_{4,-7}^{(2)}$ $\;=\;\frac{19}{27}\cdot\frac{E_{4}}{j+3375}-3375\cdot\frac{E_{4}}{(j+3375)^{2}}$
$G_{4,-7}^{(3)}$ $\;=\;\frac{6995}{11907}\cdot\frac{E_{4}}{j+3375}-\frac{55875}{7}\cdot\frac{E_{4}}{(j+3375)^{2}}+22781250\cdot\frac{E_{4}}{(j+3375)^{3}}.$
Here it is easy to check numerically that $G_{4,-7}^{(2)}$ is $1$-magnetic after clearing the denominators. [^14] After clearing the denominators, this is the same example discussed in [Bor26, Section 5.2].
To give concrete examples, we use the formulations via $\Omega_{p,D}$ and consider $p=2$ and $p=13$, with $(\frac{-7}{2})=1$ and $(\frac{-7}{13})=-1$. Note also that primes that divide $D$, $j_{D}$, or $j_{D}-1728$ are $3,5,7$.
For $p=2$, we have
$\Omega_{2,-7}\;=\;1+O(2^{3})\;\equiv\;1\pmod{2^{3}}$,
and one observes numerically that for all $n\in\mathbb{Z}^{+}$,
$a_{2n}(G_{4,-7}^{(1)})$ $\;\equiv\;a_{n}(G_{4,-7}^{(1)})$ $\pmod{2^{3}}$
$a_{2n}(G_{4,-7}^{(2)})$  $\;\equiv\;2{\hskip 1.0pt}a_{n}(G_{4,-7}^{(2)})$ $\pmod{2^{5}}$
$a_{2n}(G_{4,-7}^{(3)})$  $\;\equiv\;4{\hskip 1.0pt}a_{n}(G_{4,-7}^{(3)})$ $\pmod{2^{10}}$.
Here 4.1 only predicts congruences modulo $2^{3}$, though we observe stronger congruences for the latter two.
For $p=13$, we have
$\Omega_{13,-7}\;=\;9+8\cdot 13+4\cdot 13^{2}+O(13^{3})\;\equiv\;789\pmod{13^{3}}$,
and one notes that $c_{-7}=-\tfrac{25}{81}$ by Remark 4.6. Then, one observes numerically that for all $n\in\mathbb{Z}^{+}$,
$a_{13n}(G_{4,-7}^{(1)})$ $\;\equiv\;{\scriptstyle(2!)^{-1}\cdot(-\frac{25}{81})^{-1}\cdot(-13)^{2}\cdot 789}\cdot a_{n}(G_{4,-7}^{(3)})$  $\;\equiv\;1183{\hskip 1.0pt}a_{n}(G_{4,-7}^{(3)})$ $\pmod{13^{3}}$
$a_{13n}(G_{4,-7}^{(2)})$ $\;\equiv\;-13{\hskip 1.0pt}a_{n}(G_{4,-7}^{(2)})$ $\pmod{13^{3}}$
$a_{13n}(G_{4,-7}^{(3)})$  $\;\equiv\;{\scriptstyle 2!\cdot(-\frac{25}{81})\cdot 789^{-1}}\cdot a_{n}(G_{4,-7}^{(1)})$  $\;\equiv\;67{\hskip 1.0pt}a_{n}(G_{4,-7}^{(1)})$ $\pmod{13^{3}}$.
Example 4.3.
Let $k=6$. By computations, we have
$G_{6,D}^{(1)}$ $\;=\;{\scriptstyle\frac{E_{6}}{j-j_{D}}}$
$G_{6,D}^{(2)}$ $\;=\;{\scriptstyle\frac{E_{6}}{j-j_{D}}\cdot\big(\frac{2}{3}+\frac{\psi_{D}}{3}\big)+\frac{E_{6}}{(j-j_{D})^{2}}\cdot j_{D}}$
$G_{6,D}^{(3)}$ $\;=\;{\scriptstyle\frac{E_{6}}{j-j_{D}}\cdot\left(\frac{13}{36}\cdot\frac{j_{D}}{j_{D}-1728}+\big(\frac{2}{9}+\frac{\psi_{D}}{3}+\frac{\psi_{D}^{2}}{12}\big)\right)+\frac{E_{6}}{(j-j_{D})^{2}}\cdot j_{D}\left(\frac{1}{2}\cdot\frac{j_{D}}{j_{D}-1728}+\big(2+\frac{\psi_{D}}{2}\big)\right)+\frac{E_{6}}{(j-j_{D})^{3}}\cdot 2j_{D}^{2}}$
$G_{6,D}^{(4)}$ $\;=\;{\scriptstyle\frac{E_{6}}{j-j_{D}}\cdot\left(\big(\frac{11}{18}+\frac{13\psi_{D}}{72}\big)\cdot\frac{j_{D}}{j_{D}-1728}+\big(\frac{\psi_{D}}{9}+\frac{\psi_{D}^{2}}{12}+\frac{\psi_{D}^{3}}{72}\big)\right)+\frac{E_{6}}{(j-j_{D})^{2}}\cdot j_{D}\left(\big(\frac{245}{72}+\frac{\psi_{D}}{4}\big)\cdot\frac{j_{D}}{j_{D}-1728}+\big(\frac{20}{9}+\psi_{D}+\frac{\psi_{D}^{2}}{8}\big)\right)}$
${\scriptstyle\hskip 28.45274pt+\frac{E_{6}}{(j-j_{D})^{3}}\cdot j_{D}^{2}\left(3\cdot\frac{j_{D}}{j_{D}-1728}+\big(8+\psi_{D}\big)\right)+\frac{E_{6}}{(j-j_{D})^{4}}\cdot 6j_{D}^{3}}$
$G_{6,D}^{(5)}$ $\;=\;{\scriptstyle\frac{E_{6}}{j-j_{D}}\cdot\left(\frac{277}{864}\cdot\big(\frac{j_{D}}{j_{D}-1728}\big)^{2}+\big(\frac{5}{12}+\frac{11\psi_{D}}{54}+\frac{13\psi_{D}^{2}}{432}\big)\cdot\frac{j_{D}}{j_{D}-1728}+\big(\frac{\psi_{D}^{2}}{54}+\frac{\psi_{D}^{3}}{108}+\frac{\psi_{D}^{4}}{864}\big)\right)}$
${\scriptstyle\hskip 28.45274pt+\frac{E_{6}}{(j-j_{D})^{2}}\cdot j_{D}\left(\frac{31}{18}\cdot\big(\frac{j_{D}}{j_{D}-1728}\big)^{2}+\big(\frac{2095}{216}+\frac{245\psi_{D}}{216}+\frac{\psi_{D}^{2}}{24}\big)\cdot\frac{j_{D}}{j_{D}-1728}+\big(\frac{40}{27}+\frac{20\psi_{D}}{27}+\frac{\psi_{D}^{2}}{6}+\frac{\psi_{D}^{3}}{72}\big)\right)}$
${\scriptstyle\hskip 28.45274pt+\frac{E_{6}}{(j-j_{D})^{3}}\cdot j_{D}^{2}\left(\frac{3}{2}\cdot\big(\frac{j_{D}}{j_{D}-1728}\big)^{2}+\big(\frac{242}{9}+\psi_{D}\big)\cdot\frac{j_{D}}{j_{D}-1728}+\big(\frac{160}{9}+\frac{8\psi_{D}}{3}+\frac{\psi_{D}^{2}}{6}\big)\right)}$
${\scriptstyle\hskip 28.45274pt+\frac{E_{6}}{(j-j_{D})^{4}}\cdot j_{D}^{3}\left(18\cdot\frac{j_{D}}{j_{D}-1728}+(40+2\psi_{D})\right)+\frac{E_{6}}{(j-j_{D})^{5}}\cdot 24j_{D}^{4}}.$
Example 4.4.
Let $k=6$ and let D=28D=-28 so that $j_{D}=255^{3}$ and $\psi_{D}=85/133$. We will focus on $r=2,4$. By Example 4.3, we have
$G_{6,-28}^{(2)}$ $\;=\;{\scriptstyle\frac{117}{133}\cdot\frac{E_{6}}{j-255^{3}}+255^{3}\cdot\frac{E_{6}}{(j-255^{3})^{2}}}$
$G_{6,-28}^{(4)}$ $\;=\;{\scriptstyle\frac{477506455}{571690791}\cdot\frac{E_{6}}{j-255^{3}}+\frac{2922956049071875}{27223371}\cdot\frac{E_{6}}{(j-255^{3})^{2}}+\frac{1155258750095437500}{361}\cdot\frac{E_{6}}{(j-255^{3})^{3}}+6\cdot 255^{9}\cdot\frac{E_{6}}{(j-255^{3})^{4}}}.$
Here it is easy to check numerically that both $G_{6,-28}^{(2)}$ and $G_{6,-28}^{(4)}$ are $1$-magnetic after clearing the denominators.
As before, we consider concrete examples by choosing $p=11$ and $p=13$, with $(\frac{-28}{11})=1$ and $(\frac{-28}{13})=-1$. Note also that primes that divide $D$, $j_{D}$, or $j_{D}-1728$ are $2,3,5,7,17,19$.
For $p=11$, we have
$\Omega_{11,-7}\;=\;5+10\cdot 11+11^{2}+9\cdot 11^{3}+9\cdot 11^{4}+O(11^{5})\;\equiv\;143984\pmod{11^{5}}$,
and one observes numerically that for all $n\in\mathbb{Z}^{+}$,
$a_{11n}(G_{6,-28}^{(2)})$  $\;\equiv\;{\scriptstyle 11\cdot 143984}\cdot a_{n}(G_{6,-28}^{(2)})$  $\;\equiv\;134365{\hskip 1.0pt}a_{n}(G_{6,-28}^{(2)})$ $\pmod{11^{5}}$
$a_{11n}(G_{6,-28}^{(4)})$  $\;\equiv\;{\scriptstyle 11^{3}\cdot 143984^{-1}}\cdot a_{n}(G_{6,-28}^{(4)})$  $\;\equiv\;26620{\hskip 1.0pt}a_{n}(G_{6,-28}^{(4)})$ $\pmod{11^{5}}$.
For $p=13$, we have
$\Omega_{13,-7}\;=\;9+8\cdot 13+4\cdot 13^{2}+4\cdot 13^{3}+11\cdot 13^{4}+O(13^{5})\;\equiv\;323748\pmod{13^{5}}$,
and one notes that $c_{-28}=-\tfrac{28900}{263169}$ by Remark 4.6. Then, one observes numerically that for all $n\in\mathbb{Z}^{+}$,
$a_{13n}(G_{6,-28}^{(2)})$ $\;\equiv\;{\scriptstyle(3!)^{-1}\cdot(-\frac{28900}{263169})^{-1}\cdot(-13)^{3}\cdot 323748}\cdot a_{n}(G_{6,-28}^{(4)})$  $\;\equiv\;4394{\hskip 1.0pt}a_{n}(G_{6,-28}^{(4)})$ $\pmod{13^{5}}$
$a_{13n}(G_{6,-28}^{(4)})$  $\;\equiv\;{\scriptstyle 3!\cdot(-\frac{28900}{263169})\cdot(-13)\cdot 323748^{-1}}\cdot a_{n}(G_{6,-28}^{(2)})$  $\;\equiv\;185653{\hskip 1.0pt}a_{n}(G_{6,-28}^{(2)})$ $\pmod{13^{5}}$.
Example 4.5.
Let $k=6$ and let D=27D=-27 so that $j_{D}=-3\cdot 160^{3}$ and $\psi_{D}=160/253$. We will focus on $r=1,5$. By Example 4.3, we have
$G_{6,-27}^{(1)}$ $\;=\;{\scriptstyle\frac{E_{6}}{j+3\cdot 160^{3}}}$
$G_{6,-27}^{(5)}$ $\;=\;{\scriptstyle\frac{3637833600}{4097152081}\cdot\frac{E_{6}}{j+3\cdot 160^{3}}-\frac{713572831887360000}{4097152081}\cdot\frac{E_{6}}{(j+3\cdot 160^{3})^{2}}+\frac{30034105244414115840000000}{4097152081}\cdot\frac{E_{6}}{(j+3\cdot 160^{3})^{3}}}$
$\hskip 28.45274pt{\scriptstyle-\frac{7038224068289495040000000000}{64009}\cdot\frac{E_{6}}{(j+3\cdot 160^{3})^{4}}+24\cdot 3^{4}\cdot 160^{12}\cdot\frac{E_{6}}{(j+3\cdot 160^{3})^{5}}}.$
We now consider $p=7$ and $p=17$, with $(\frac{-27}{7})=1$ and $(\frac{-27}{17})=-1$. Note also that primes that divide $D$, $j_{D}$, or $j_{D}-1728$ are $2,3,5,11,23$.
For $p=7$, we have
$\Omega_{7,-3}\;=\;1+5\cdot 7+5\cdot 7^{2}+4\cdot 7^{3}+7^{4}+O(7^{5})\;\equiv\;4054\pmod{7^{5}}$,
and one observes numerically that for all $n\in\mathbb{Z}^{+}$,
$a_{7n}(G_{6,-27}^{(1)})$  $\;\equiv\;4054^{2}\cdot a_{n}(G_{6,-27}^{(1)})$  $\;\equiv\;14477{\hskip 1.0pt}a_{n}(G_{6,-27}^{(1)})$ $\pmod{7^{5}}$
$a_{7n}(G_{6,-27}^{(5)})$  $\;\equiv\;7^{4}\cdot 4054^{-2}\cdot a_{n}(G_{6,-27}^{(5)})$  $\;\equiv\;2401{\hskip 1.0pt}a_{n}(G_{6,-27}^{(5)})$ $\pmod{7^{5}}$.
For $p=17$, we have
$\Omega_{17,-3}\;=\;16+2\cdot 17+5\cdot 17^{2}+6\cdot 17^{3}+10\cdot 17^{4}+O(17^{5})\;\equiv\;866183\pmod{17^{5}}$,
and one notes that $c_{-27}\;=\;-\tfrac{3600}{64009}$. Then, one observes numerically that for all $n\in\mathbb{Z}^{+}$,
$a_{17n}(G_{6,-27}^{(1)})$ $\;\equiv\;{\scriptstyle(4!)^{-1}\cdot(-\frac{3600}{64009})^{-2}\cdot(-17)^{4}\cdot 866183^{2}}\cdot a_{n}(G_{6,-27}^{(5)})$  $\;\equiv\;417605{\hskip 1.0pt}a_{n}(G_{6,-27}^{(5)})$ $\pmod{17^{5}}$
$a_{17n}(G_{6,-28}^{(5)})$  $\;\equiv\;{\scriptstyle 4!\cdot(-\frac{28900}{263169})^{2}\cdot 866183^{-2}}\cdot a_{n}(G_{6,-27}^{(1)})$  $\;\equiv\;850024{\hskip 1.0pt}a_{n}(G_{6,-27}^{(1)})$ $\pmod{17^{5}}$.
4.4. The cases when $D\in\{-3,-4\}$
We will now turn to the cases when $D\in\{-3,-4\}$, or equivalently, when $j(C)\in\{0,1728\}$. While the heuristics in these cases are essentially the same as in the other CM cases, we first briefly discuss the behaviors of $\frac{E_{k}}{j}$ and $\frac{E_{k}}{j-1728}$ for motivation.
The reason why $\frac{E_{k}}{j}$ and $\frac{E_{k}}{j-1728}$ behave differently is that they do not necessarily have a simple pole and thus resemble $G_{k,D}^{(r)}$ for some $r>1$ rather than $r=1$. For example, the following holds.
  1. (1)
    $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$ are both $1$-magnetic.
  2. (2)
    $\frac{E_{6}}{j}$ is $2$-magnetic.
  3. (3)
    $\frac{E_{8}}{j-1728}$ is (numerically) $1$-magnetic.
  4. (4)
    $\frac{E_{10}}{j}$ is (numerically) $1$-magnetic.
When $\frac{E_{k}}{j}$ or $\frac{E_{k}}{j-1728}$ does have a simple pole, all the phenomena in Section 2 still hold for $\frac{E_{k}}{j}$ or $\frac{E_{k}}{j-1728}$, though one would need to replace the condition that $v_{p}(j(C))=0=v_{p}(j(C)-1728)$ by some other appropriate condition.
Example 4.6.
Consider $\frac{E_{14}}{j}$. Let $C/\mathbb{Q}$ be the elliptic curve given by $y^{2}+y=x^{3}$ with LMFDB label 27.a4 and $j(C)=0$. Then, we have the following numerical congruences.
  1. (1)
    For all primes $p$ with $(\tfrac{-3}{p})=1$ and all $n,l\in\mathbb{Z}^{+}$,
    $a_{np^{l}}\bigg(\frac{E_{14}}{j}\bigg)\;\equiv\;u_{p}(C)^{12}a_{np^{l-1}}\bigg(\frac{E_{14}}{j}\bigg)\pmod{p^{13l}}$,
    where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ that is a $p$-adic unit.
  2. (2)
    For all primes $p$ with $p\geq 13$ and $(\tfrac{-3}{p})=-1$ and all $n,l\in\mathbb{Z}^{+}$,
    $a_{np^{l}}\bigg(\frac{E_{14}}{j}\bigg)\;\equiv\;p^{12}a_{np^{l-2}}\bigg(\frac{E_{14}}{j}\bigg)\pmod{p^{13l-1}}$.
Here one could remove the assumption that $p\geq 13$ in the supersingular case by considering $12!\cdot\frac{E_{4}}{j}$ instead. The appearance of $12!$ should be related to the appearance of (k2)!(k-2)! in 4.5 when $r=1$ and r=k2r=k-2 (see also 5.3).
We now turn to the general case. We will describe everything in terms of $\Omega_{p,D}$ so that there is no need to make a choice of an elliptic curve $C/\mathbb{Q}$ with CM by $\mathcal{O}_{D}$. Let $k\in\{4,6,8,10,14\}$. Let $D\in\{-3,-4\}$ and define $\alpha_{D}\in\mathcal{H}$ and $\mathcal{O}_{D}$ accordingly. Consider
$\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(\alpha_{D})},\;\frac{E_{k}}{(j-j(\alpha_{D}))^{2}},\ldots,\frac{E_{k}}{(j-j(\alpha_{D}))^{\kappa}}\bigg\},$
where $\kappa=\kappa_{k,D}$ is the largest integer such that $\frac{E_{k}}{(j-j(\alpha_{D}))^{\kappa}}$ has a pole of order $\leq k-1$. For $1\leq i\leq\kappa$, let $r_{i}$ denote the order of the pole of $\frac{E_{k}}{(j-j(\alpha_{D}))^{i}}$. Here it is important to note that for $1\leq i\leq\kappa$,
$r_{i}+r_{\kappa+1-i}\;=\;k.$
Now, we still consider $(\partial_{\tau}^{r}G_{k})(z,\tau)$ but with different normalizations so that the evaluation at $\alpha_{D}$ makes sense, does not vanish, and is independent of the choice of $\alpha_{D}$.
For D=3D=-3, one considers for $1\leq i\leq\kappa$ that
$G_{k,-3}^{(r_{i})}(z)\;:=\;\bigg(j(\tau)^{-\lfloor\frac{r_{i}-1}{3}\rfloor}\cdot\frac{\Delta}{E_{14-k}}(\tau)\cdot\bigg(-\frac{E_{4}}{E_{6}}(\tau)\bigg)^{r_{i}-1}\cdot(\partial_{\tau}^{r_{i}-1}G_{k})(z,\tau)\bigg)\bigg|_{\tau=\alpha_{-3}}.$ (16)
For D=4D=-4, one considers for $1\leq i\leq\kappa$ that
$G_{k,-4}^{(r_{i})}(z){\hskip 1.0pt}:={\hskip 1.0pt}\bigg((j(\tau)-1728)^{\lfloor\frac{r_{i}-1}{2}\rfloor}\cdot\frac{\Delta}{E_{14-k}}(\tau)\cdot\bigg(-\frac{E_{4}}{E_{6}}(\tau)\bigg)^{r_{i}-1}\cdot(\partial_{\tau}^{r_{i}-1}G_{k})(z,\tau)\bigg)\bigg|_{\tau=\alpha_{-4}}.$ (17)
The correction factors here should be related to the normalizing constant $c_{D}$ in 4.5, as we will see in Remark 4.22. This indicates the potential existence of a uniform normalization of $G_{k,D}^{(r)}$ that leads to a uniform description of the $p$-adic behaviors of $G_{k,D}^{(r)}$.
Conjecture 4.20.
Let $k\in\{4,6,8,10,14\}$, D=3D=-3 or 4-4, and $1\leq i\leq\kappa_{k,D}$. Let $p\geq 5$ be a prime with $(\frac{D}{p})=1$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r_{i})})\;\equiv\;p^{r_{i}-1}\Omega_{p,D}^{\frac{k}{2}-r_{i}}a_{np^{l}}(G_{k,D}^{(r_{i})})\pmod{p^{(k-1)l}}$.
For supersingular primes, we need to consider different normalizing constants from the one considered in 4.5.
Conjecture 4.21.
Let $k\in\{4,6,8,10,14\}$, D=3D=-3 or 4-4, and $1\leq i\leq\kappa_{k,D}$. Let $p\geq 5$ be a prime with $(\frac{D}{p})=1$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$(r_{i}-1)!\cdot a_{np^{l}}(G_{k,D}^{(k-r_{i})})\;\equiv\;(k-1-r_{i})!{\hskip 1.0pt}c_{D}^{\frac{k}{2}-r_{i}}\cdot(-p)^{r_{i}-1}\Omega_{p,D}^{-(\frac{k}{2}-r_{i})}a_{np^{l-1}}(G_{k,D}^{(r_{i})})\pmod{p^{(k-1)l}}$,
where we write $j_{D}=j(\alpha_{D})$ and
$c_{D}\;=\;\begin{cases}(-1)^{D}\cdot(j_{D}-1728)^{-1}\cdot D\;=\;-\frac{1}{576}&\text{if }D=-3,\\ (-1)^{D}\cdot|j_{D}|^{2/3}\cdot D\;=\;-576&\text{if }D=-4.\end{cases}$
Remark 4.22.
This constant $c_{D}$ is essentially the same as the one in 4.5. To make the comparison, let $G_{k,D}^{(r)}$ and $c_{D}$ denote the modular form and the normalizing constant given in 4.5 and let $\widehat{G}_{k,D}^{(r_{i})}$ and $\widehat{c}_{D}$ denote the ones given in 4.21. For D=3D=-3, one observes that formally
$\widehat{G}_{k,D}^{(r_{i})}\;=\;j_{D}^{-\lfloor\frac{r_{i}-1}{3}\rfloor}\cdot G_{k,D}^{(r_{i})}.$
Plugging this in 4.5, one formally obtains that
$(r_{i}-1)!a_{np^{l}}(\widehat{G}_{k,D}^{(k-r_{i})})\;\equiv\;(k-1-r_{i})!{\hskip 1.0pt}\widetilde{c}\cdot(-p)^{r_{i}-1}\Omega_{p,D}^{-(\frac{k}{2}-r_{i})}a_{np^{l-1}}(\widehat{G}_{k,D}^{(r_{i})})\pmod{p^{(k-1)l}}$,
where
$\widetilde{c}\;=\;j_{D}^{\lfloor\frac{r_{i}-1}{3}\rfloor-\lfloor\frac{k-r_{i}-1}{3}\rfloor}\cdot c_{D}^{\frac{k}{2}-r_{i}}.$
Now, one notices that $k-r_{i}=r_{\kappa+1-i}$ and that $r_{i}\equiv r_{j}\pmod{3}$ for all $1\leq i,j\leq\kappa$, so
$\lfloor\tfrac{r_{i}-1}{3}\rfloor-\lfloor\tfrac{k-r_{i}-1}{3}\rfloor\;=\;-\tfrac{k-2r_{i}}{3}.$
Hence,
$\widetilde{c}\;=\;\big((-1)^{D}\cdot(j_{D}-1728)^{-1}\cdot D\big)^{\frac{k}{2}-r_{i}}\;=\;\widehat{c}_{D}^{\frac{k}{2}-r_{i}},$
yielding 4.21 when D=3D=-3. It is easy to check that the same formal argument works for D=4D=-4.
Conjecture 4.23.
Let $k\in\{4,6,8,10,14\}$, D=3D=-3 or 4-4, and $1\leq i\leq\kappa_{k,D}$. Let $A=A_{k,D}^{(r_{i})}$ be a nonzero integer such that $\widetilde{G}_{k,D}^{(r_{i})}=AG_{k,D}^{(r_{i})}$ has integer coefficients with no common divisors. Then, $\widetilde{G}_{k,D}^{(r_{i})}$ is $(r_{i}^{\prime}-1)$-magnetic, i.e., for all $n\in\mathbb{Z}^{+}$,
$n^{r_{i}^{\prime}-1}{\hskip 1.0pt}\big|{\hskip 1.0pt}a_{n}(\widetilde{G}_{k,D}^{(r_{i})}),$
where $r_{i}^{\prime}=\min(r_{i},k-r_{i})$.
The family $\{G_{k,D}^{(r_{i})}\}_{1\leq i\leq\kappa}$ should also satisfy other phenomena observed in Section 4.2 which we will not rewrite here. We will now end this subsection with some examples.
Example 4.7.
Let $k=4$. We have $\kappa=1$ and $r_{1}=2$ for both $D\in\{-3,-4\}$. In this case,
$G_{4,-3}^{(2)}\;=\;\frac{1}{3}\cdot\frac{E_{4}}{j}\quad\text{ and }\quad G_{4,-4}^{(2)}\;=\;864\cdot\frac{E_{4}}{j-1728}.$
These identifications now provide a heuristic explanation for the properties of $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$:
  1. (1)
    the $1$-magnetic properties, i.e., Theorem 1.1, should be explained by 4.11;
  2. (2)
    the supercongruences, i.e., Theorem 1.2, should be explained by Theorem 4.13.
Similarly, the $2$-magnetic property and the supercongruence of $\frac{E_{6}}{j}$ described in Remark 1.12 can be explained through this heuristic by considering $k=6$ and D=3D=-3.
Example 4.8.
Let $k=6$ and D=4D=-4. We have $\kappa=3$ and $r_{1}=1,r_{2}=3,r_{3}=5$. In this case,
$G_{6,-4}^{(1)}$ $\;=\;\frac{E_{6}}{j-1728}$
$G_{6,-4}^{(3)}$ = $\;=\;624\cdot\frac{E_{6}}{j-1728}+1492992\cdot\frac{E_{6}}{(j-1728)^{2}}$
$G_{6,-4}^{(5)}$ = $\;=\;957312\cdot\frac{E_{6}}{j-1728}+8886288384\cdot\frac{E_{6}}{(j-1728)^{2}}+13374150672384\cdot\frac{E_{6}}{(j-1728)^{3}}.$
Here it is easy to check numerically that $G_{6,-4}^{(3)}$ (in fact $\frac{1}{48}G_{6,-4}^{(3)}$) is $2$-magnetic.
To give concrete examples, we choose $p=5$ and $p=7$, with $(\frac{-4}{5})=1$ and $(\frac{-4}{7})=-1$. Since the primes are small in this case, we will look at the $np^{2}$-th coefficient.
For $p=5$, we have
$\Omega_{5,-4}\;=\;4+3\cdot 5+4\cdot 5^{3}+5^{4}+5^{5}+3\cdot 5^{6}+2\cdot 5^{8}+2\cdot 5^{9}+O(5^{10})\;\equiv\;4738644\pmod{5^{10}}$,
and one observes numerically that for all $n\in\mathbb{Z}^{+}$,
$a_{25n}(G_{6,-4}^{(1)})$  $\;\equiv\;4738644^{2}\cdot a_{5n}(G_{6,-4}^{(1)})$  $\;\equiv\;864986{\hskip 1.0pt}a_{5n}(G_{6,-4}^{(1)})$ $\pmod{5^{10}}$
$a_{25n}(G_{6,-4}^{(3)})$  $\;\equiv\;5^{2}\cdot a_{5n}(G_{6,-4}^{(3)})$ $\pmod{5^{11}}$
$a_{25n}(G_{6,-4}^{(5)})$  $\;\equiv\;5^{4}\cdot 4738644^{-2}\cdot a_{5n}(G_{6,-4}^{(5)})$  $\;\equiv\;8900625{\hskip 1.0pt}a_{5n}(G_{6,-4}^{(5)})$ $\pmod{5^{11}}$.
For $p=7$, we have
$\Omega_{7,-4}$ = $\;=\;6+2\cdot 7+4\cdot 7^{2}+4\cdot 7^{3}+5\cdot 7^{4}+5\cdot 7^{5}+5\cdot 7^{6}+4\cdot 7^{7}+7^{8}+2\cdot 7^{9}+O(7^{10})$
 $\;\equiv\;90452060\pmod{7^{10}}$,
and one notes that $c_{-4}=-576$. Then, one observes numerically that for all $n\in\mathbb{Z}^{+}$,
$a_{49n}(G_{6,-4}^{(1)})$ $\;\equiv\;{\scriptstyle(4!)^{-1}\cdot(-576)^{-2}\cdot(-7)^{4}\cdot 90452060^{2}}\cdot a_{7n}(G_{6,-4}^{(5)})$  $\;\equiv\;165035136{\hskip 1.0pt}a_{7n}(G_{6,-4}^{(5)})$ $\pmod{7^{10}}$
$a_{49n}(G_{6,-4}^{(3)})$ $\;\equiv\;(-7)^{2}\cdot a_{7n}(G_{6,-4}^{(3)})$ $\pmod{7^{11}}$
$a_{49n}(G_{6,-4}^{(5)})$  $\;\equiv\;{\scriptstyle 4!\cdot(-576)^{2}\cdot 90452060^{-2}}\cdot a_{7n}(G_{6,-4}^{(1)})$  $\;\equiv\;199720267{\hskip 1.0pt}a_{7n}(G_{6,-4}^{(1)})$ $\pmod{7^{11}}$.
In both cases, 4.20 or 4.21 only predicts congruences modulo $p^{10}$, though we observe stronger congruences for the latter two.
Example 4.9.
Let $k=8$ and D=3D=-3. We have $\kappa=3$ and $r_{1}=1$, $r_{2}=4$, $r_{3}=7$. In this case, we have
$G_{8,-3}^{(1)}$ $\;=\;\frac{E_{8}}{j}$
$G_{8,-3}^{(4)}$ $\;=\;-\frac{1}{13824}\cdot\frac{E_{8}}{j}+\frac{2}{9}\cdot\frac{E_{8}}{j^{2}}$
$G_{8,-3}^{(7)}$ $\;=\;\frac{227}{7739670528}\cdot\frac{E_{8}}{j}-\frac{295}{559872}\cdot\frac{E_{8}}{j^{2}}+\frac{80}{81}\cdot\frac{E_{8}}{j^{3}}.$
Here it is easy to check numerically that $G_{8,-3}^{(4)}$ is $3$-magnetic after clearing the denominators. As before, we choose $p=5$ and $p=7$, with $(\frac{-3}{5})=-1$ and $(\frac{-3}{7})=1$.
For $p=7$, we have
$\Omega_{7,-3}$ = $\;=\;1+5\cdot 7+5\cdot 7^{2}+4\cdot 7^{3}+7^{4}+6\cdot 7^{5}+4\cdot 7^{6}+O(7^{7})\;\equiv\;575492\pmod{7^{7}}$,
and one observes numerically that for all $n\in\mathbb{Z}^{+}$,
$a_{7n}(G_{8,-3}^{(1)})$  $\;\equiv\;575492^{3}\cdot a_{n}(G_{8,-3}^{(1)})$  $\;\equiv\;705608{\hskip 1.0pt}a_{n}(G_{8,-3}^{(1)})$ $\pmod{7^{7}}$
$a_{7n}(G_{8,-3}^{(4)})$  $\;\equiv\;7^{3}\cdot a_{n}(G_{8,-3}^{(4)})$ $\pmod{7^{8}}$
$a_{7n}(G_{8,-3}^{(7)})$  $\;\equiv\;7^{6}\cdot 575492^{-3}\cdot a_{n}(G_{8,-3}^{(7)})$  $\;\equiv\;7^{6}\cdot a_{n}(G_{8,-3}^{(7)})$ $\pmod{7^{7}}$.
Here 4.20 only predicts congruences modulo $7^{7}$, though we observe a stronger congruence for the middle one.
For $p=5$, we have
$\Omega_{5,-3}$ = $\;=\;4+5+3\cdot 5^{2}+2\cdot 5^{3}+5^{4}+3\cdot 5^{5}+5^{6}+O(5^{7})\;\equiv\;25959\pmod{5^{7}}$,
and one notes that $c_{-3}=-\frac{1}{576}$. Then, one observes numerically that for all $n\in\mathbb{Z}^{+}$,
$6!\cdot a_{5n}(G_{8,-3}^{(1)})$ $\;\equiv\;{\scriptstyle(-\frac{1}{576})^{-3}\cdot(-5)^{6}\cdot 25959^{3}}\cdot a_{n}(G_{8,-3}^{(7)})$  $\;\equiv\;5^{6}\cdot a_{n}(G_{8,-3}^{(7)})$ $\pmod{5^{7}}$
$a_{5n}(G_{8,-3}^{(4)})$ $\;\equiv\;(-5)^{3}\cdot a_{n}(G_{8,-3}^{(4)})$ $\pmod{5^{7}}$
$a_{5n}(G_{8,-3}^{(7)})$  $\;\equiv\;{\scriptstyle 6!\cdot(-\frac{1}{576})^{3}\cdot 25959^{-3}}\cdot a_{n}(G_{8,-3}^{(1)})$  $\;\equiv\;62445\cdot a_{n}(G_{8,-3}^{(1)})$ $\pmod{5^{7}}$.
In the first congruence, since $5$ divides 6!=(k2)!6!=(k-2)!, it is important that we copy the exact formula written in 4.21, i.e., put $6!$ in front of $G_{8,-3}^{(1)}$ rather than divide both sides by it as we always did previously. Otherwise, the congruence would only hold modulo $5^{6}$.
5. Generalization to other weights
We will now discuss the generalization of the whole setting to other weights. In general, the coefficients of meromorphic modular forms having precisely one pole at a non-cuspidal point are an interplay between cusp forms and the elliptic curve corresponding to the pole.
Let $g\in\mathbb{Z}[[q]]$ be a holomorphic modular form of weight $k$ and level $1$ and let $C/\mathbb{Q}$ be an elliptic curve (with $j(C)\notin\{0,1728\}$) such that $\frac{g}{j-j(C)}$ has a pole. For $r\in\mathbb{Z}^{+}$, write
$F_{g,C}^{(r)}\;:=\;\frac{g}{(j-j(C))^{r}}\;=\;\sum_{n=1}^{\infty}a_{n}(F_{g,C}^{(r)})q^{n}.$
In general, it is necessary to remove the influence from the cusp forms to obtain the same congruences observed for $F_{k,C}^{(r)}$ in Section 2 and Section 3. This is possible via the help of relations. An integer sequence $\lambda=(\lambda_{m})_{m\in\mathbb{Z}^{+}}$ is called a relation for $S_{k}=S_{k}(1)$ if
  1. (1)
    $\lambda_{m}=0$ for all but finitely many $m$;
  2. (2)
    $\sum_{m=1}^{\infty}\lambda_{m}a_{m}\;=\;0$ for any cusp form $\sum_{m=1}^{\infty}a_{m}q^{m}\in S_{k}$.
For a modular form $F$ of weight $k$ and a relation $\lambda$ for $S_{k}$, write
$F|\lambda\;:=\;\sum_{m=1}^{\infty}\lambda_{m}F|T_{m,k},$ (18)
where $T_{m,k}$ is the $m$-th Hecke operator of weight $k$. Note that $\sum_{m=1}^{\infty}\lambda_{m}(\sum_{m=1}^{\infty}a_{m}q^{m})|T_{m,k}=0$ for any cusp form $\sum_{m=1}^{\infty}a_{m}q^{m}\in S_{k}$, so replacing $F$ by $F|\lambda$ will “kill the cuspidal part” of FF. [^15] The consideration of relations first appeared in Gross–Zagier [GZ86, p. 316]. This idea was used in Löbrich–Schwagenscheidt [LS21, p. 2383] to construct examples of magnetic modular forms and in Brown–Fonseca [BF25, Remark 8.10] to study meromorphic modular forms.
Now, one simply fixes a relation $\lambda$ for $S_{k}$ and considers $F_{g,C}^{(r)}|\lambda$. Then, all the phenomena in Section 2 and Section 3 should still hold after replacing $F_{k,C}^{(r)}$ by $F_{g,C}^{(r)}|\lambda$. As an example, we have the following generalization of Theorem 1.7.
Theorem 5.1.
Let $g\in\mathbb{Z}[[q]]$ be a holomorphic modular form of weight $k$ and level $1$, $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$, and $\lambda$ be a relation for $S_{k}$. Let $p\geq 5$ be a good prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n\in\mathbb{Z}^{+}$,
$a_{np}(F_{g,C}^{(1)}|\lambda)\;\equiv\;a_{p}(C)^{k-2}a_{n}(F_{g,C}^{(1)}|\lambda)\pmod{p}$.
Remark 5.2.
The theorem is conjecturally true for $p=2,3$. Depending on $k\pmod{12}$, the theorem is also conjecturally true for primes $p$ with $v_{p}(j(C))\neq 0$ or $v_{p}(j(C)-1728)\neq 0$, which we will describe here. Let
$S_{0,C}\;:=\;\{p\mid v_{p}(j(C))>0\}\quad\text{ and }\quad S_{1728,C}\;:=\;\{p\mid v_{p}(j(C)-1728)>0\}.$
For $k\in 2\mathbb{Z}_{\geq 2}$, define
[…]
Then, the theorem should hold conjecturally for all good primes $p$ of $C$ with $p\notin S^{(k,C)}$. For example, when $k\equiv 14\pmod{12}$, the theorem should simply hold conjecturally for all good primes of $C$.
In the supersingular case, we also have the following generalization of 3.4.
Conjecture 5.3.
Let $g\in\mathbb{Z}[[q]]$ be a holomorphic modular form of weight $k$ and level $1$, $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$, and $\lambda$ be a relation for $S_{k}$. Let $p$ be a supersingular prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $1\leq r\leq k-1$ and for all $n,l\in\mathbb{Z}^{+}$,
$(k-1-r)!\cdot a_{np^{l}}(F_{g,C}^{(r)}|\lambda)\;\equiv\;(k-1-r)!\cdot p^{k-2}a_{np^{l-2}}(F_{g,C}^{(r)}|\lambda)\pmod{p^{(k-1)l-r}}$,
where $a_{m}(\;\cdot\;):=0$ if $m\notin\mathbb{Z}$.
Remark 5.4.
Here the extra factor (k1r)!(k-1-r)! should be related to the same extra factor that appears in 4.5, which is important for the congruence when $p\leq k-1-r$. The reason that this factor does not appear previously is because the weight is too small. If $k\leq 14$, then $p\leq k-2$ implies that $p\leq 11$. In this case, it is easy to check that the only supersingular $j$-invariants in $\overline{\mathbb{F}_{p}}$ are 0 or $1728$. Hence, if $p\leq 11$ is a supersingular prime of $C$, then we must have $v_{p}(j(C))\neq 0$ or $v_{p}(j(C)-1728)\neq 0$, which is then excluded by the assumption that $v_{p}(j(C))=0=v_{p}(j(C)-1728)$.
Example 5.1 ($k=12$).
Let $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$ and let $g$ be a holomorphic modular form of weight $12$ with integer coefficients such that
$F_{g,C}\;:=\;\frac{g}{j-j(C)}$
has a pole. Since
$\Delta\;=\;\sum_{m=1}^{\infty}\tau(m)q^{m}$
is the only cusp form of weight $12$, for every $m\geq 2$, there is an obvious choice of the relation $\lambda_{m}=(-\tau(m),0,\ldots,0,1,0,0,\ldots)$, where the $1$ is at the $m$-th entry, and one has that
$F_{g,C}|\lambda_{m}\;=\;F_{g,C}|_{12}(T_{m,12}-\tau(m)).$
Let $p\geq 5$ be a good prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, applying Theorem 5.1 with $n=1$ yields that
$a_{p}(F_{g,C}|\lambda_{m})\;\equiv\;a_{p}(C)^{10}{\hskip 1.0pt}a_{1}(F_{g,C}|\lambda_{m})\pmod{p}$,
or equivalently,
$a_{mp}(F_{g,C})-\tau(m)a_{p}(F_{g,C})\;\equiv\;a_{m}(F_{g,C})a_{p}(C)^{10}-\tau(m)a_{p}(C)^{10}\pmod{p}$.
Let $p$ be a supersingular prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, one observes numerically that
$a_{p}(F_{g,C}|\lambda_{m})\;\equiv\;0\pmod{p^{10}}$,
or equivalently,
$a_{mp}(F_{g,C})-\tau(m)a_{p}(F_{g,C})\;\equiv\;0\pmod{p^{10}}$.
This example essentially shows that the coefficients of $F_{g,C}$ are conjecturally “determined” (up to congruences) by the coefficients of $\Delta$ and the $a_{p}(C)$’s.
As in the cases when $k\in\{4,6,8,10,14\}$ where only $g=E_{k}$ is considered, for a general weight $k$, it is also sufficient to consider one particular holomorphic modular form. For any $k\in 2\mathbb{Z}$, define
$g_{k}\;:=\;E_{k^{\prime}}\Delta^{\delta},$
where $k^{\prime}\in\{0,4,6,8,10,14\}$ and $\delta\in\mathbb{Z}$ satisfying that $k=k^{\prime}+12\delta$. Indeed, $g_{k}=E_{k}$ when $k\in\{4,6,8,10,14\}$ and note also that $\delta=\dim S_{k}$ for $k\geq 4$. Before stating the next lemma, we recall that for a prime $p$, one defines
$\mathbb{Z}_{(p)}\;:=\;\{a\in\mathbb{Q}\mid v_{p}(a)\geq 0\},$
i.e., the ring of $p$-integral (rational) numbers.
Lemma 5.5.
Let $k\in 2\mathbb{Z}_{\geq 2}$. Let $p$ be a prime and let $\alpha\in\mathcal{H}$ satisfy that $j(\alpha)\in\mathbb{Z}_{(p)}$. Let $g\in\mathbb{Z}_{(p)}[[q]]$ be a holomorphic modular form of weight $k$ and level $1$. Then, for any $r\in\mathbb{Z}^{+}$, $\frac{g}{(j-j(\alpha))^{r}}$ is a $\mathbb{Z}_{(p)}$-linear combination of $\frac{g_{k}}{j-j(\alpha)},\frac{g_{k}}{(j-j(\alpha))^{2}},\ldots,\frac{g_{k}}{(j-j(\alpha))^{r}}$, and cusp forms of weight $k$ and level $1$.
Proof.
We will prove this by induction. By analyzing the location of the poles, it is easy to check that $g/g_{k}$ can be written as a polynomial in $j$, which has $p$-integral coefficients (see also [KZ98]). In particular, $A:=g(\alpha)/g_{k}(\alpha)\in\mathbb{Z}_{(p)}$. Write
$h\;=\;\frac{g-Ag_{k}}{j-j(\alpha)}\in\mathbb{Z}_{(p)}[[q]].$
Then, $h$ is holomorphic on $\mathcal{H}$ and vanishes at $\infty$, so $h\in S_{k}$. Now, for $r=1$,
$\frac{g}{j-j(\alpha)}\;=\;h+A\cdot\frac{g_{k}}{j-j(\alpha)},$
so it satisfies the requirement. For $r>1$,
$\frac{g}{(j-j(\alpha))^{r}}\;=\;\frac{h}{(j-j(\alpha))^{r-1}}+A\cdot\frac{g_{k}}{(j-j(\alpha))^{r}}.$
The result then follows by applying the induction hypothesis on $h$ and r1r-1. ∎
Remark 5.6.
It is easy to check that the same proof holds in the version of number fields, which we will need in Section 7.
By Lemma 5.5, it suffices to only investigate properties of $\frac{g_{k}}{(j-j(C))^{r}}$. This is especially useful in the CM case. In this case, one can similarly obtain suitable linear combinations of $\{\frac{g_{k}}{(j-j(C))^{r}}\}_{1\leq r\leq k-1}$ via differentiating an immediate generalization of $G_{k}(z,\tau)$ such that these linear combinations, after applying $\lambda$, would still satisfy the congruences in Section 4. We will illustrate this now, and we will adopt all the notations in Section 4 on negative discriminants.
For $k\in 2\mathbb{Z}_{\geq 2}$, write
$G_{k}(z,\tau)\;:=\;\frac{g_{k}(z)\cdot g_{2-k}(\tau)}{j(z)-j(\tau)}.$ (19)
For $1\leq r\leq k-1$, write
$G_{k,D}^{(r)}(z)\;:=\;\bigg(g_{2-k}(\tau)^{-1}\cdot\bigg(-\frac{E_{4}}{E_{6}}(\tau)\bigg)^{r-1}\cdot(\partial_{\tau}^{r-1}G_{k})(z,\tau)\bigg)\bigg|_{\tau=\alpha_{D}}.$ (20)
Now, one simply fixes a relation $\lambda$ for $S_{k}$ and considers $G_{k,D}^{(r)}|\lambda$. Then, all the phenomena in Section 4 should still hold after replacing $G_{k,D}^{(r)}$ with $G_{k,D}^{(r)}|\lambda$. As an example, we have the following conjectures.
Conjecture 5.7.
Let $k\in 2\mathbb{Z}_{\geq 2}$ and let $1\leq r\leq k-1$. Let D<4D<-4 be a discriminant with class number $1$, $C/\mathbb{Q}$ be an elliptic curve with $j(C)=j(\alpha_{D})$, and $\lambda$ be a relation for $S_{k}$. Let $p$ be a prime with $(\frac{D}{p})=1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(r)}|\lambda)\;\equiv\;p^{r-1}u_{p}(C)^{k-2r}a_{np^{l-1}}(G_{k,D}^{(r)}|\lambda)\pmod{p^{(k-1)l}}$,
where $u_{p}(C)$ is the root of $X^{2}-a_{p}(C)X+p$ which is a $p$-adic unit.
Conjecture 5.8.
Let $k\in 2\mathbb{Z}_{\geq 2}$, $r\in\mathbb{Z}$ with $1\leq r\leq k-1$, and $\lambda$ be a relation for $S_{k}$. Let $D<0$ be a discriminant with class number $1$ and write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}$ and some fundamental discriminant $D_{0}<0$. Suppose that $D\notin\{-3,-4,-12\}$. Let $p$ be a prime with $(\frac{D}{p})=-1$ and $v_{p}(j(\alpha_{D}))=0=v_{p}(j(\alpha_{D})-1728)$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$(r-1)!\cdot a_{np^{l}}(G_{k,D}^{(k-r)}|\lambda)\;\equiv\;(k-1-r)!{\hskip 1.0pt}c_{D}^{\frac{k}{2}-r}\cdot(-p)^{r-1}\Omega_{p,D_{0}}^{-(\frac{k}{2}-r)}a_{np^{l-1}}(G_{k,D}^{(r)}|\lambda)\pmod{p^{(k-1)l}}$,
where $c_{D}$ is the same as in 4.5 (see also Remark 4.6).
Conjecture 5.9.
Let $k\in 2\mathbb{Z}_{\geq 2}$, $r\in\mathbb{Z}$ with $1\leq r\leq k-1$, and $r^{\prime}=\min(r,k-r)$. Let D<4D<-4 be a discriminant with class number $1$ and $\lambda$ be a relation for $S_{k}$. Let $A=A_{k,D}^{(r)}$ be a nonzero integer such that $\widetilde{G}_{k,D}^{(r)}=AG_{k,D}^{(r)}$ has integer coefficients with no common divisors. Then, $\widetilde{G}_{k,D}^{(r)}|\lambda$ is $(r^{\prime}-1)$-magnetic, i.e., for all $n\in\mathbb{Z}^{+}$,
$n^{r^{\prime}-1}{\hskip 1.0pt}\big|{\hskip 1.0pt}a_{n}(\widetilde{G}_{k,D}^{(r)}|\lambda).$
For the middle element $G_{k,D}^{(k/2)}$ in the family, we also have a different and presumably better normalization, which yields a generalization of Theorem 1.10. For $k\in 2\mathbb{Z}_{\geq 2}$, let
$\mathcal{G}_{k,D}(z)\;:=\;\frac{2}{w_{D}}\cdot(\partial^{\frac{k-2}{2}}_{\tau}G_{k})(z,\alpha_{D}).$ (21)
Theorem 5.10.
Let $k\in 2\mathbb{Z}_{\geq 2}$ and $\lambda$ be a relation for $S_{k}$. Let $D<0$ be a discriminant of class number $1$ and write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}$ and some fundamental discriminant $D_{0}<0$. Let
[…]
Then, $\widetilde{\mathcal{G}}_{k,D}$ satisfies the following.
  1. (1)
    $\widetilde{\mathcal{G}}_{k,D}\in\mathbb{Z}[[q]]$, and $\widetilde{\mathcal{G}}_{k,D}|\lambda$ is $\frac{k-2}{2}$-magnetic, i.e., for all $n\in\mathbb{Z}^{+}$,
    $n^{\frac{k-2}{2}}{\hskip 1.0pt}\big|{\hskip 1.0pt}a_{n}(\widetilde{\mathcal{G}}_{k,D}|\lambda).$
  2. (2)
    Let $p$ be a prime with $p\nmid A$. Then, for all $n,l\in\mathbb{Z}^{+}$,
    $a_{np^{l}}(\widetilde{\mathcal{G}}_{k,D}|\lambda)\;\equiv\;\big(\big(\tfrac{D}{p}\big)p\big)^{\frac{k-2}{2}}a_{np^{l-1}}(\widetilde{\mathcal{G}}_{k,D}|\lambda)\pmod{p^{(k-1)l}}$.
We now end this section with some numerical examples illustrating some of the phenomena described above.
Example 5.2.
Let $k\in 2\mathbb{Z}_{\geq 2}$. Let D<4D<-4 be a discriminant of class number $1$ and write $j_{D}=j(\alpha_{D})$ and $\psi_{D}=\psi(\alpha_{D})$ as in Section 4.3. Now, write $g_{2-k}=E_{4}^{a}E_{6}^{b}\Delta^{\delta^{\prime}}$ for some $a,b,\delta^{\prime}\in\mathbb{Z}$. Then,
$G_{k,D}^{(2)}\;=\;\frac{g_{k}}{j-j_{D}}\cdot\left(\tfrac{b}{2}\cdot\tfrac{j_{D}}{j_{D}-1728}+\tfrac{a}{3}+\tfrac{k-2}{12}\cdot\psi_{D}\right)+\frac{g_{k}}{(j-j_{D})^{2}}\cdot j_{D}.$
Fix a relation $\lambda$ for $S_{k}$. Then, 5.9 predicts that a suitable multiple of $G_{k,D}^{(2)}|\lambda$ is $1$-magnetic.
For the supercongruence predicted by 5.7, we will write it in the form of Remark 4.2. Let $p$ be a prime with $(\frac{D}{p})=1$ and $v_{p}(j_{D})=0=v_{p}(j_{D}-1728)$. Write $p=\pi\overline{\pi}$ for some $\pi\in\mathcal{O}_{D}$. Then, 5.7 predicts that for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(G_{k,D}^{(2)}|\lambda)\;\equiv\;p{\hskip 1.0pt}\pi^{k-4}{\hskip 1.0pt}a_{np^{l-1}}(G_{k,D}^{(2)}|\lambda)\pmod{\overline{\pi}^{(k-1)l}}$.
Example 5.3.
Let $k=12$ so that $a=2$ and $b=1$ in Example 5.2, and let D=11D=-11 so that $j_{D}=-32^{3}$ and $\psi_{D}=32/77$. Then,
$G_{12,-11}^{(2)}\;=\;\frac{802}{539}\cdot\frac{\Delta}{j+32^{3}}-32^{3}\cdot\frac{\Delta}{(j+32^{3})^{2}}$
by Example 5.2. Note that primes that divide $D$, $j_{D}$, or $j_{D}-1728$ are exactly $2,7,11$.
Consider $p=5$ so that $(\frac{-11}{p})=1$ and fix a relation $\lambda$ for $S_{12}$. We will write the supercongruence explicitly using $\Omega_{p,D}$. In this case,
$\Omega_{5,-11}\;=\;4+4\cdot 5+5^{4}+2\cdot 5^{6}+5^{9}+4\cdot 5^{10}+O(5^{11})\;\equiv\;41047524\pmod{5^{11}}$.
Thus, for all $n\in\mathbb{Z}^{+}$, we have numerically that
$a_{5n}(G_{12,-11}^{(2)}|\lambda)\;\equiv\;5\cdot\Omega_{5,-11}^{4}\cdot a_{n}(G_{12,-11}^{(2)}|\lambda)\;\equiv\;46880755{\hskip 1.0pt}a_{n}(G_{12,-11}^{(2)}|\lambda)\pmod{5^{11}}$.
Now, consider explicitly $\lambda=\lambda_{5}=(-4830,0,0,0,1,0,\ldots)$ as defined in Example 5.1, which may also illustrate the independence between the relation $\lambda$ and the prime $p$. In this case,
$G_{12,-11}^{(2)}|\lambda_{5}\;=\;G_{12,-11}^{(2)}|(T_{5,12}-4830)\;\equiv\;G_{12,-11}^{(2)}|(U_{5}-4830)\pmod{5^{11}}$.
Unfolding the congruence above in this case, one obtains that
$a_{25n}(G_{12,-11}^{(2)})-46885585{\hskip 1.0pt}a_{5n}(G_{12,-11}^{(2)})+18031025{\hskip 1.0pt}a_{n}(G_{12,-11}^{(2)})\;\equiv\;0\pmod{5^{11}}$.
Example 5.4.
We give some examples of magnetic modular forms deduced from Theorem 5.10.
For $k=4$, one obtains $1$-magnetic modular forms that are integer multiples of examples in [PZ22, Table 1]. To give two new examples not documented in that table, the case when D=27D=-27 yields that
$64E_{4}\bigg(\frac{60083}{j+12288000}+\frac{64009\cdot(-12288000)}{(j+12288000)^{2}}\bigg)$
is $1$-magnetic, and the case when D=24D=-24 yields that
$48E_{4}\cdot\frac{3923j^{3}+1271674944j^{2}+33025786269696j-31712851087589376}{(j^{2}-4834944j+14670139392)^{2}}$
is $1$-magnetic (here the denominator is the square of the Hilbert class polynomial associated to 24-24). Indeed, numerical computations suggest that both examples are already $1$-magnetic without the pre-factors $64$ and $48$.
For higher weights, the expressions are already quite complicated. We will thus only provide the examples when D=3,4D=-3,-4 and $k=6,8,10,14$.
For $k=6$, we obtain that
$384\cdot\frac{E_{6}}{j}\hskip 14.22636pt\text{ and }\hskip 14.22636pt96E_{6}\bigg(\frac{13}{j-1728}+\frac{31104}{(j-1728)^{2}}\bigg)$
are $2$-magnetic.
For $k=8$, we obtain that
$8E_{8}\bigg(\frac{1}{j}-\frac{3072}{j^{2}}\bigg)\hskip 14.22636pt\text{ and }\hskip 14.22636pt12E_{8}\bigg(\frac{1}{j-1728}+\frac{5832}{(j-1728)^{2}}\bigg)$
are $3$-magnetic.
For $k=10$, we obtain that
$24E_{10}\bigg(\frac{13}{j}-\frac{110592}{j^{2}}\bigg)\hskip 14.22636pt\text{ and }\hskip 14.22636pt96E_{10}\bigg(\frac{17}{j-1728}+\frac{324864}{(j-1728)^{2}}+\frac{644972544}{(j-1728)^{3}}\bigg)$
are $4$-magnetic.
For $k=14$, we obtain that
$144E_{14}\bigg(\frac{1}{j}-\frac{89280}{j^{2}}+\frac{318504960}{j^{3}}\bigg)\qquad\text{ and }\qquad 96E_{14}\bigg(\frac{13}{j-1728}+\frac{2223744}{(j-1728)^{2}}+\frac{19707494400}{(j-1728)^{3}}+\frac{33435376680960}{(j-1728)^{4}}\bigg)$
are $6$-magnetic.
In all the cases, numerical computations suggest that all the modular forms are already magnetic without the corresponding pre-factors.
6. Notations and preliminaries for the proofs
In the last two sections of this paper, we will demonstrate the proofs of Theorem 5.1 and Theorem 5.10. For convenience and for the coherence of the proof, we define some notations and recall some preliminaries in this section.
6.1. Modular forms of integral weight
Fix $k\in\mathbb{Z}$. Let $M_{k}$ denote the space of holomorphic modular forms of weight $k$ and level $1$, $S_{k}$ denote the space of holomorphic cusp forms of weight $k$ and level $1$, and $M_{k}^{!}$ denote the space of weakly holomorphic modular forms of weight $k$ and level $1$.
Now, write $k=k^{\prime}+12\delta$ for some $k^{\prime}\in\{0,4,6,8,10,14\}$ and $\delta\in\mathbb{Z}$. For each $n\in\mathbb{Z}$ with $n\geq-\delta$, there exists a unique weakly holomorphic modular form $g_{k,n}$ of weight $k$ and level $1$ satisfying that
$g_{k,n}\;=\;q^{-n}+O(q^{\delta+1})\in\mathbb{Z}[[q,q^{-1}]].$
It is easy to check that $g_{k,-\delta}=g_{k}$ and that $\{g_{k,n}\}_{n\geq-\delta}$ forms a basis of $M_{k}^{!}$.
6.2. Modular forms of half-integral weight
Fix $s\in\mathbb{Z}$. Let $M_{s+1/2}^{!,+}$ denote the Kohnen plus space, i.e., the space of weakly holomorphic modular forms of weight $s+1/2$ and level $\Gamma_{0}(4)$ whose $n$-th Fourier coefficient is zero if $(-1)^{s}n\not\equiv 0,1\pmod{4}$. Let $S_{s+1/2}^{+}$ denote the Kohnen plus cuspidal subspace.
Now, write $2s=k^{\prime}+12\delta$ for some $k^{\prime}\in\{0,4,6,8,10,14\}$ and $\delta\in\mathbb{Z}$. Let
[…]
Then, for each $m\in\mathbb{Z}$ with $m\geq-A$ and $(-1)^{s-1}m\equiv 0,1\pmod{4}$, there exists a unique $f_{s+1/2,m}\in M_{s+1/2}^{!,+}$ such that
$f_{s+1/2,m}\;=\;q^{-m}+O(q^{A+1})\in\mathbb{Z}[[q,q^{-1}]].$
In particular, $\{f_{s+1/2,m}\}_{m}$ forms a basis of $M_{s+1/2}^{!,+}$ (see [DJ08b, Section 2 and Appendix] for a discussion of this fact).
6.3. Hecke operators (on $q$-series)
Let $f=\sum_{n\gg-\infty}a_{n}(f)q^{n}$ be a $q$-series. Let $k\in\mathbb{Z}$ and $p$ be a rational prime. Define the Hecke operators (on $q$-series) as
$f|U_{p}\;:=\;\sum_{n\gg-\infty}a_{np}(f)q^{n},\quad f|V_{p}\;:=\;\sum_{n\gg-\infty}a_{n}(f)q^{np},\quad f|T_{p,k}\;:=\;f|(U_{p}+p^{k-1}V_{p}).$
More generally, for $m\in\mathbb{Z}^{+}$, define
$f|T_{m,k}\;:=\;\sum_{n\gg-\infty}\bigg(\sum_{d>0,d|(m,n)}d^{k-1}a_{mn/d^{2}}(f)\bigg)q^{n}$
It is easy to check that $T_{m,k}$ and $T_{n,k}$ commute for any $m,n\in\mathbb{Z}^{+}$ and $k\in\mathbb{Z}$.
For a map $\chi:\mathbb{Z}\rightarrow\mathbb{C}$ (usually a Dirichlet character), define
$f|\chi\;:=\;\sum_{n\gg-\infty}\chi(n)a_{n}(f)q^{n}.$
Now, let $s\in\mathbb{Z}$ and let
$f\;=\;\sum_{(-1)^{s}n\equiv 0,1\;(\mathrm{mod}\;4)}a_{n}(f)q^{n}\in M_{s+1/2}^{+,!}.$
For a rational prime $p$, define
$f|T_{p,s+1/2}\;:=\;\sum_{(-1)^{s}n\equiv 0,1\;(\mathrm{mod}\;4)}\big(a_{np^{2}}(f)+p^{s-1}\big(\tfrac{(-1)^{s}n}{p}\big)a_{n}(f)+p^{2s-1}a_{n/p^{2}}(f)\big)q^{n}.$
Indeed, if $p$ is odd, then
$f|T_{p,s+1/2}\;=\;f|(U_{p}^{2}+p^{s-1}\chi_{p}^{\prime}+p^{2s-1}V_{p}^{2}),$
where $\chi_{p}^{\prime}(n)=\big(\frac{(-1)^{s}n}{p}\big)$.
6.4. Negative discriminants
Let $D<0$ be a discriminant, $\mathcal{O}_{D}$ be the imaginary quadratic order of discriminant $D$, and
$w_{D}\;:=\;|\mathcal{O}_{D}^{\times}|\;=\;\begin{cases}4&\text{ if $D=-4$,}\\ 6&\text{ if $D=-3$,}\\ 2&\text{ otherwise.}\end{cases}$
Let $\mathcal{Q}_{D}$ be a complete set of $\mathrm{SL}_{2}(\mathbb{Z})$-inequivalent positive definite quadratic forms $Q(X,Y)\in\mathbb{Z}[X,Y]$ of discriminant $D$ and let $\mathcal{Q}_{D}^{\mathrm{prim}}\subseteq\mathcal{Q}_{D}$ be the subset consisting of primitive elements, i.e., $aX^{2}+bXY+cY^{2}$ with $(a,b,c)=1$. For $Q(X,Y)=aX^{2}+bXY+cY^{2}\in\mathcal{Q}_{D}$, define
$\alpha_{Q}\;:=\;\frac{-b+\sqrt{D}}{2a}\in\mathcal{H}$
and
[…]
7. Proof of Theorem 1.6 and Theorem 1.7
We will now prove Theorem 5.1. As we will work with general number fields, we first introduce some notations.
Let $L$ be a number field and $\mathcal{O}_{L}$ be its ring of integers. For a prime $\mathfrak{p}$ of $L$, let $\mathbb{F}_{\mathfrak{p}}$ denote its residue field and let $\mathrm{N}(\mathfrak{p}):=|\mathbb{F}_{\mathfrak{p}}|$ denote its norm. For an elliptic curve $C/L$ and a good prime $\mathfrak{p}$ of $C$, define
$a_{\mathfrak{p}}\;:=\;\mathrm{N}(\mathfrak{p})+1-|C(\mathbb{F}_{\mathfrak{p}})|.$
Now, for a holomorphic modular form $g\in\mathcal{O}_{L}[[q]]$ of weight $k$ and level $1$ and an elliptic curve $C/L$, write
$F_{g,C}\;:=\;\frac{g}{j-j(C)}.$
We also adopt the notation $g_{k}$ in Section 5, and for $k\in 2\mathbb{Z}_{\geq 2}$, write
$F_{k,C}\;:=\;F_{g_{k},C}\;=\;\frac{g_{k}}{j-j(C)}.$
The main theorem that we will prove in this section is the following.
Theorem 7.1.
Let $k\in 2\mathbb{Z}_{\geq 2}$, $L$ be a number field, $g\in\mathcal{O}_{L}[[q]]$ be a holomorphic modular form of weight $k$ and level $1$, and $C/L$ be an elliptic curve. Let $\lambda$ be a relation for $S_{k}$ and let $\mathfrak{p}$ be a good prime of $C$ with $\mathfrak{p}\nmid 6$ and $v_{\mathfrak{p}}(j(C))=0=v_{\mathfrak{p}}(j(C)-1728)$. Then, for all $n\in\mathbb{Z}^{+}$,
$a_{n\cdot\mathrm{N}(\mathfrak{p})}(F_{g,C}|\lambda)\;\equiv\;a_{\mathfrak{p}}(C)^{k-2}a_{n}(F_{g,C}|\lambda)\pmod{\mathfrak{p}}$.
Remark 7.2.
By Lemma 5.5, it suffices to prove this for $g=g_{k}$, i.e., for $F_{k,C}$. Indeed, there exist some $\beta\in\mathcal{O}_{L}$ such that $F_{g,C}-\beta F_{k,C}\in S_{k}$. Hence, $F_{g,C}|\lambda=\beta\cdot F_{k,C}|\lambda$.
As mentioned in Remark 1.9, the proof depends on the hypergeometric interpretation of the coefficients in the case when $k=4$. In particular, we will show the following generalization of Theorem 1.6.
Theorem 7.3.
Let $p$ be a rational prime with $p\nmid 6$ and let $c\in\overline{\mathbb{Z}_{p}}$ with $v_{p}(c)=0$. Then, for all $l\in\mathbb{Z}^{+}$,
$a_{p^{l}}\bigg(\frac{E_{4}}{j-c}\bigg)$ $\;\equiv\;\big(c(c-1728)\big)^{\frac{p^{l}-1}{2}}\cdot\sum_{m=0}^{p^{l}-1}\frac{(6m)!}{(m!)^{3}(3m)!}{\hskip 1.0pt}c^{-m}\pmod{p}$.
7.1. The two variable function
From now on, we fix $k\in 2\mathbb{Z}_{\geq 2}$, and write $k=k^{\prime}+12\delta$ for some $k^{\prime}\in\{0,4,6,8,10,14\}$ and $\delta\in\mathbb{Z}$. Recall that
$G_{k}(z,\tau)\;=\;\frac{g_{k}(z)\cdot g_{2-k}(\tau)}{j(z)-j(\tau)}.$
Let $q_{z}=e^{2\pi iz}$ and $q_{\tau}=e^{2\pi i\tau}$. Then, we have the following result by Duke–Jenkins, from which they deduced Zagier’s duality.
Theorem 7.4 ([DJ08a, Theorem 2]).
For any even integer $k\in\mathbb{Z}^{+}$,
$\sum_{n=\delta+1}^{\infty}g_{2-k,n}(\tau)q_{z}^{n}\;=\;G_{k}(z,\tau)\;=\;-\sum_{n=-\delta}^{\infty}g_{k,n}(z)q_{\tau}^{n}.$ (22)
Fix a rational prime $p$ and let $l\in\mathbb{Z}^{+}$. It is easy to check from the definitions that
$T_{p^{l},k}\;=\;\sum_{a=0}^{l}p^{a(k-1)}U_{p}^{l-a}V_{p}^{a}.$
Now, fix $n\in\mathbb{Z}$ with $n\geq-\delta$ and consider $g_{k,n}(z)|T_{p^{l},k}$. If $n\geq 0$, then we have
$g_{k,n}(z)\big|T_{p^{l},k}\;=\;\sum_{a=0}^{l}p^{a(k-1)}q_{z}^{-np^{2a-l}}+O(q^{\delta+1}),$
and it follows that
$g_{k,n}(z)\big|T_{p^{l},k}-\sum_{a=0}^{l}p^{a(k-1)}g_{k,np^{2a-l}}(z)\in S_{k},$ (23)
where we abuse the notations by setting that $q_{z}^{\alpha}:=0$ if $\alpha\notin\mathbb{Z}$, and that $g_{k,\alpha}:=0$ if $\alpha\notin\mathbb{Z}$, for convenience. If $\delta\leq n<0$, then we have
$g_{k,n}(z)\in S_{k}\quad\text{ and }\quad g_{k,n}(z)\big|T_{p^{l},k}\in S_{k}.$ (24)
Now, let $\lambda$ be a relation for $S_{k}$. In the following computations, the relation $\lambda$ and all the Hecke operators are viewed as acting on the variable $z$. We have
[…]
$\;\;\equiv\;-\sum_{n=0}^{\infty}\big(g_{k,np^{-s}}(z)\big|\lambda\big)q_{\tau}^{n}$
[…]
as a double Fourier series in terms of $z$ and $\tau$. Also,
$G_{k}(z,p^{l}\tau)\;=\;\frac{g_{k}(z)\cdot g_{2-k}(p^{l}\tau)}{j(z)-j(p^{l}\tau)}\;\equiv\;\frac{g_{k}(z)\cdot g_{2-k}(\tau)^{p^{l}}}{j(z)-j(\tau)^{p^{l}}}\pmod{p}$
as a double Fourier series in terms of $z$ and $\tau$. We thus obtain the following.
Proposition 7.5.
For a rational prime $p$, some $l\in\mathbb{Z}^{+}$, and a relation $\lambda$ for $S_{k}$, we have
$\frac{g_{k}(z)}{j(z)-j(\tau)}\bigg|\lambda U_{p^{l}}\;\equiv\;g_{2-k}(\tau)^{p^{l}-1}\cdot\frac{g_{k}(z)}{j(z)-j(\tau)^{p^{l}}}\bigg|\lambda\pmod{p}$
as a double Fourier series in terms of $z$ and $\tau$.
7.2. Reducing to the case of weight 44
We will now prove that all the cases follow from the single case when $k=4$ and $\lambda=(1,0,0,\ldots)$.
Write
$\frac{g_{k}}{j-X}\bigg|\lambda\;=\;\sum_{n=1}^{\infty}P_{k,n,\lambda}(X)q^{n}$
and in particular,
$\frac{E_{4}}{j-X}\;=\;\sum_{n=1}^{\infty}P_{4,n}(X)q^{n},$
for some polynomials $P_{k,n,\lambda},P_{4,n}\in\mathbb{Z}[X]$.
Lemma 7.6.
For a rational prime $p$ and $l\in\mathbb{Z}^{+}$,
$P_{4,p^{l}}(j(\tau))\;\equiv\;(g_{-2}(\tau))^{p^{l}-1}\pmod{p}$
as $q_{\tau}$-series.
Proof.
This follows from specializing Proposition 7.5 at $k=4$ and $\lambda=(1,0,0,\ldots)$. ∎
Now, by analyzing the potential poles and zeros, it is easy to check that there always exist $a,b\in\mathbb{Z}_{\geq 0}$ such that
$\frac{g_{-2}(\tau)^{\frac{k-2}{2}}}{g_{2-k}(\tau)}\;=\;j(\tau)^{a}(j(\tau)-1728)^{b}.$
Then, Proposition 7.5 and Lemma 7.6 together imply that
$\bigg(\frac{g_{-2}(\tau)^{\frac{k-2}{2}}}{g_{2-k}(\tau)}\bigg)^{p^{l}-1}\cdot\frac{g_{k}(z)}{j(z)-j(\tau)}\bigg|\lambda U_{p^{l}}\;\equiv\;P_{4,p^{l}}(j(\tau))^{\frac{k-2}{2}}\cdot\frac{g_{k}(z)}{j(z)-j(\tau)^{p^{l}}}\bigg|\lambda\pmod{p}$
as a double Fourier series in terms of $z$ and $\tau$. In particular, taking the $n$-th coefficient of $q_{z}$ on both sides, one obtains that
$\big(j(\tau)^{a}(j(\tau)-1728)^{b}\big)^{p^{l}-1}P_{k,np^{l},\lambda}(j(\tau))\;\equiv\;P_{4,p^{l}}(j(\tau))^{\frac{k-2}{2}}P_{k,n,\lambda}(j(\tau)^{p^{l}})\pmod{p}$ (25)
as $q_{\tau}$-series, and hence also as polynomials in $j(\tau)$.
Now, let $L$ be a number field, $\mathfrak{p}$ be a prime of $L$, and $c\in L$ with $v_{\mathfrak{p}}(c)=0=v_{\mathfrak{p}}(c-1728)$. Let $p$ be a rational prime lying under $\mathfrak{p}$ and write $\mathrm{N}(\mathfrak{p})=p^{l}$. Then, plugging in $j(\tau)=c$ in (25) yields that
$P_{k,np^{l},\lambda}(c)\;\equiv\;P_{4,p^{l}}(c)^{\frac{k-2}{2}}P_{k,n,\lambda}(c)\pmod{\mathfrak{p}}$.
In particular, we obtain the following.
Lemma 7.7.
Let $k\in 2\mathbb{Z}_{\geq 2}$, $L$ be a number field, and $C/L$ be an elliptic curve. Let $\lambda$ be a relation for $S_{k}$ and let $\mathfrak{p}$ be a prime with $v_{\mathfrak{p}}(j(C))=0=v_{\mathfrak{p}}(j(C)-1728)$. Then, for all $n\in\mathbb{Z}^{+}$,
$a_{n\cdot\mathrm{N}(\mathfrak{p})}(F_{k,C}|\lambda)\;\equiv\;a_{\mathrm{N}(\mathfrak{p})}(F_{4,C})^{\frac{k-2}{2}}{\hskip 1.0pt}a_{n}(F_{k,C}|\lambda)\pmod{\mathfrak{p}}$.
From the lemma above, it suffices to show that
$a_{\mathrm{N}(\mathfrak{p})}(F_{4,C})\;\equiv\;a_{\mathfrak{p}}(C)^{2}\pmod{\mathfrak{p}}$
in order to show Theorem 5.1.
7.3. Hypergeometric reinterpretation of $a_{p^{l}}(F_{4,C})$
Now, we specialize to $a_{p^{l}}(F_{4,C})$. Specifically, we will prove Theorem 7.3. For this part, it is more convenient to forget the dependence on the elliptic curves and simply write
$F_{4,c}\;:=\;\frac{E_{4}}{j-c},$
for $c\in\overline{\mathbb{Q}}$ or $\overline{\mathbb{Z}_{p}}$.
Recall that $a_{p^{l}}(F_{4,c})=P_{4,p^{l}}(c)$ and
$P_{4,p^{l}}(j)\;\equiv\;g_{-2}^{p^{l}-1}\;\equiv\;(E_{10}/\Delta)^{p^{l}-1}\pmod{p}$
as $q$-series. In order to prove Theorem 7.3, we will use the hypergeometric intepretation of $E_{4}$.
We first recall the definition of hypergeometric functions. Given a hypergeometric datum $(\boldsymbol{\alpha},\boldsymbol{\beta})$ with
$\boldsymbol{\alpha}$ $\;=\;(\alpha_{1},\ldots,\alpha_{n-1},\alpha_{n})\in\mathbb{C}^{n}$
$\boldsymbol{\beta}$ $\;=\;(\beta_{1},\ldots,\beta_{n-1})\in\mathbb{C}^{n-1},$
the associated hypergeometric function is defined as
[…]
and for $r\in\mathbb{Z}_{\geq 0}$, the truncated hypergeometric sum is defined as
[…]
where $(a)_{m}:=\prod_{i=0}^{m-1}(a+i)$.
Recall the following identity by Fricke and Klein [Fri12, Chapter 5, Section 10]
[…]
and Clausen’s formula
[…]
It follows that
[…]
Now, let $p\nmid 6$ be a rational prime and let $l\in\mathbb{Z}^{+}$. Write
$Q_{n}(X)\;:=\;X^{n-1}P_{4,n}(X^{-1}).$
Then,
$Q_{p^{l}}(j^{-1})\;=\;j^{1-p^{l}}P_{4,p^{l}}(j)$ $\;\equiv\;j^{1-p^{l}}(E_{10}/\Delta)^{p^{l}-1}$
[…]
as $q$-series. Let $\alpha=j^{-1}=q+O(q^{2})$. Then,
[…]
$Q_{p^{l}}(\alpha)\;\equiv\;(1-1728\alpha)^{\frac{p^{l}-1}{2}}B(\alpha)^{1-p^{l}}\;\equiv\;(1-1728\alpha)^{\frac{p^{l}-1}{2}}B(\alpha)\cdot(B^{-1})(\alpha^{p^{l}})\pmod{p}$,
where $B^{-1}$ is the inverse of $B$ as a power series. Since $\deg Q_{p^{l}}=\deg P_{4,p^{l}}\leq p^{l}-1$ and $B(0)=1$, this implies that
[…]
Here for a power series $R(t)=\sum_{n=0}^{\infty}a_{n}t^{n}$, we write
$\mathrm{Trun}_{m}(R)(t)\;:=\;\sum_{n=0}^{m}a_{n}t^{n}\quad\text{ and }\quad\mathrm{Trun}_{m,t}\bigg(\sum_{n=0}^{\infty}a_{n}t^{n}\bigg)\;:=\;\sum_{n=0}^{m}a_{n}t^{n}.$
For $p\nmid 6$, it is easy to check that $\frac{(\frac{1}{2})_{r}(\frac{1}{6})_{r}(\frac{5}{6})_{r}}{(r!)^{3}}$ is divisible by $p$ for $\frac{p^{l}-1}{6}<r\leq p^{l}-1$. From this, one can easily deduce that
$\mathrm{Trun}_{p^{l}-1,\alpha}\big((1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot B(\alpha)\big)\;\equiv\;(1-1728\alpha)^{\frac{p^{l}-1}{2}}\cdot\mathrm{Trun}_{p^{l}-1}(B)(\alpha)\pmod{p}$.
To summarize, we obtain that
[…]
as polynomials in $\alpha$.
Now, for $c\in\overline{\mathbb{Z}_{p}}$ with $v_{p}(c)=0$,
$a_{p^{l}}(F_{4,c})\;=\;P_{4,p^{l}}(c)$ $\;=\;c^{p^{l}-1}Q_{p^{l}}(c^{-1})$
[…]
$\;\equiv\;\big(c(c-1728)\big)^{\frac{p^{l}-1}{2}}\cdot\sum_{n=0}^{p^{l}-1}\frac{(6m)!}{(m!)^{3}(3m)!}{\hskip 1.0pt}c^{-m}\pmod{p}$,
which proves Theorem 7.3.
7.4. Proof of Theorem 5.1
Let $L$ be a number field, $C/L$ be an elliptic curve, and $\mathfrak{p}$ be a good prime of $C$ with $v_{\mathfrak{p}}(j(C))=0=v_{\mathfrak{p}}(j(C)-1728)$. Let $\mathbb{F}_{\mathfrak{p}}$ denote the residue field of $\mathfrak{p}$ and let $p$ be the rational prime lying under $\mathfrak{p}$. For $a\in L$ with $v_{\mathfrak{p}}(a)\geq 0$, define a quadratic character
[…]
It is easy to check that if $\mathfrak{p}\nmid 2$, then
$\bigg(\frac{a}{\mathfrak{p}}\bigg)\;=\;a^{\frac{\mathrm{N}(\mathfrak{p})-1}{2}}\hskip 5.69046pt\text{ in }\mathbb{F}_{\mathfrak{p}}.$
By Theorem 7.3,
[…]
It thus suffices to show that
[…]
This congruence should follow from the hypergeometric machinery, though an explicit reference seems hard to find. Here we write a sketch of the proof and refer to the expository note [Zha25] for an explicit proof of this congruence.
Write $c=j(C)$ and consider the elliptic curve
$C_{1}:\;y^{2}+xy\;=\;x^{3}-\frac{t}{432}$
with $t=\frac{1-\sqrt{1-1728/c}}{2}$ (so that $j(C_{1})=c$). It is known that for a prime $\mathfrak{P}$ of $L(\sqrt{1-1728/c})$ that is a good prime of $C_{1}$,
[…]
Now, one can relate $a_{\mathfrak{p}}(C)$ and $a_{\mathfrak{P}}(C_{1})$ via the fact that $j(C)=j(C_{1})$ and use an appropriate version of the truncated Clausen’s formula to obtain the desired congruence.
8. Proof of Theorem 1.10
We will now prove Theorem 1.10, i.e. the supercongruences and the $\frac{k-2}{2}$-magnetic property of $G_{k,D}^{(k/2)}$, or rather its generalization Theorem 5.10. In fact, we will further generalize this to a result that holds for any discriminant $D<0$.
Recall that for $k\in 2\mathbb{Z}_{\geq 2}$,
$G_{k}(z,\tau)\;=\;\frac{g_{k}(z)\cdot g_{2-k}(\tau)}{j(z)-j(\tau)}.$
For a discriminant $D<0$, define
$\mathcal{G}_{k,D}(z)\;:=\;\frac{2}{w_{D}}\sum_{Q\in\mathcal{Q}_{D}^{\mathrm{prim}}}(\partial^{\frac{k-2}{2}}_{\tau}G_{k})(z,\alpha_{Q}).$ (27)
As before, since $(\partial^{\frac{k-2}{2}}_{\tau}G_{k})(z,\tau)$ is of weight 0 with respect to $\tau$, $\mathcal{G}_{k,D}(z)$ is independent of the choice of $\mathcal{Q}_{D}^{\mathrm{prim}}$.
Theorem 8.1.
Let $k\in 2\mathbb{Z}_{\geq 2}$ and $\lambda$ be a relation for $S_{k}$. Let $D<0$ be a discriminant and write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}$ and some fundamental discriminant $D_{0}<0$. Let
[…]
Then, $\widetilde{\mathcal{G}}_{k,D}$ satisfies the following.
  1. (1)
    $\widetilde{\mathcal{G}}_{k,D}\in\mathbb{Z}[[q]]$ and $\widetilde{\mathcal{G}}_{k,D}|\lambda$ is $\frac{k-2}{2}$-magnetic, i.e., for all $n\in\mathbb{Z}^{+}$,
    $n^{\frac{k-2}{2}}{\hskip 1.0pt}\big|{\hskip 1.0pt}a_{n}(\widetilde{\mathcal{G}}_{k,D}|\lambda).$
  2. (2)
    Let $p$ be a prime with $p\nmid A$. Then, for all $n,l\in\mathbb{Z}^{+}$,
    $a_{np^{l}}(\widetilde{\mathcal{G}}_{k,D}|\lambda)\;\equiv\;\big(\big(\tfrac{D}{p}\big)p\big)^{\frac{k-2}{2}}a_{np^{l-1}}(\widetilde{\mathcal{G}}_{k,D}|\lambda)\pmod{p^{(k-1)l}}$.
Remark 8.2.
This theorem should be viewed as a natural generalization of Li–Neururer [LN19, Theorem 1.5] and Paşol–Zudilin [PZ22, Theorems 1 and 2]. Indeed, the proof uses the same method: one considers the half-integral weight preimage of $\mathcal{G}_{k,D}$ under the Shimura lift, analyzes the $U_{p}$ action on the preimage, and transfers the properties of the preimage to the supercongruences and the magnetic property of $\mathcal{G}_{k,D}$.
8.1. The Shimura lift
Let $s\in\mathbb{Z}$, $f\in M_{s+1/2}^{!,+}$, and $d_{0}$ be a fundamental discriminant $d_{0}$ with $(-1)^{s}d_{0}>0$. Define the $d_{0}$-th Shimura lift of $f$ by
$\mathcal{S}_{d_{0}}(f)\;:=\;\frac{1}{2}\cdot L\big(1-s,(\tfrac{d_{0}}{\cdot})\big){\hskip 1.0pt}a_{0}(f)+\sum_{n=1}^{\infty}\bigg(\sum_{m{\hskip 1.0pt}|{\hskip 1.0pt}n}\big(\tfrac{d_{0}}{m}\big){\hskip 1.0pt}m^{s-1}{\hskip 1.0pt}a_{|d_{0}|n^{2}/m^{2}}(f)\bigg)q^{n}.$
Now, let $d$ be a discriminant and $d_{0}$ be a fundamental discriminant with $dd_{0}<0$. For an $\mathrm{SL}_{2}(\mathbb{Z})$-invariant function $f(\tau)$ on $\mathcal{H}$, recall the following definition from [DJ08b]
$\mathrm{Tr}_{d,d_{0}}(f)\;:=\;\sum_{Q\in\mathcal{Q}_{dd_{0}}}\frac{1}{w_{Q}}\cdot\chi(Q)f(\alpha_{Q}),$ (28)
where for each $Q(X,Y)=aX^{2}+bXY+cY^{2}\in\mathcal{Q}_{dd_{0}}$, we define
[…]
Now, let $s\in\mathbb{Z}_{\geq 2}$, $d$ be a discriminant, and $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$ and $dd_{0}<0$. Write $2s=k^{\prime}+12\delta$ for some $k^{\prime}\in\{0,4,6,8,10,14\}$ and $\delta\in\mathbb{Z}$. We define
$h_{s,d,d_{0}}\;:=\;\sum_{n=0}^{\delta}a_{n}\big(\mathcal{S}_{d_{0}}f_{s+1/2,|d|}\big)\cdot g_{2s,-n}.$
Lemma 8.3.
Let $s\in\mathbb{Z}_{\geq 2}$, $d$ be a discriminant, and $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$ and $dd_{0}<0$. Then, $h_{s,d,d_{0}}$ has integer coefficients and lies in $S_{2s}$.
Proof.
The integrality of $h_{s,d,d_{0}}$ simply follows from the integrality of $f_{s+1/2,|d|}$ and $g_{2s,-n}$. Now, since $s\in\mathbb{Z}_{\geq 2}$ and $|d|>0$, it is easy to check that $a_{0}(f_{s+1/2,|d|})=0$, so $a_{0}\big(\mathcal{S}_{d_{0}}f_{s+1/2,|d|}\big)=0$. Hence, $h_{s,d,d_{0}}\in S_{2s}$. ∎
Proposition 8.4.
Let $s\in\mathbb{Z}_{\geq 2}$, $d$ be a discriminant, and $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$ and $dd_{0}<0$. Then,
$\big(\mathcal{S}_{d_{0}}f_{s+1/2,|d|}\big)(z)-h_{s,d,d_{0}}(z)\;=\;-(-1)^{\lfloor\frac{s-1}{2}\rfloor}{\hskip 1.0pt}|d|^{-\frac{s}{2}}{\hskip 1.0pt}|d_{0}|^{\frac{s-1}{2}}\;\mathrm{Tr}_{d,d_{0}}\big((\partial^{s-1}_{\tau}G_{2s})(z,\tau)\big),$
where $\mathrm{Tr}_{d,d_{0}}$ is taken with respect to $\tau$.
Proof.
See [DJ08b, Proposition 10]. ∎
8.2. The Hecke actions
We will now analyze the actions of the Hecke operators on $f_{s+1/2,m}$ and transfer them to modular forms of integral weight.
Lemma 8.5.
Let $s\in\mathbb{Z}_{\geq 2}$, $m\in\mathbb{Z}^{+}$ with $(-1)^{s-1}m\equiv 0,1\pmod{4}$, and $p$ be a prime such that
  1. (1)
    $p^{2}\nmid m$, or
  2. (2)
    if $p=2$ and $4{\hskip 1.0pt}|{\hskip 1.0pt}m$, then $(-1)^{s-1}{\hskip 1.0pt}\frac{m}{4}\equiv 2,3\pmod{4}$.
Let
$g_{0}$ $\;:=\;f_{s+1/2,m}$
$g_{1}$ $\;:=\;g_{0}|T_{p,s+1/2}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot g_{0}$
$g_{i+1}$ $\;:=\;g_{i}|T_{p,s+1/2}-p^{2s-1}\cdot g_{i-1}$
Then, for all $i\in\mathbb{Z}^{+}$,
$g_{i}\;=\;p^{(2s-1)i}q^{-mp^{2i}}+O(q)\in\mathbb{Z}[[q,q^{-1}]].$
In particular, $g_{i}-p^{(2s-1)i}\cdot f_{s+1/2,mp^{2i}}$ has integer coefficients and lies in $S_{s+1/2}^{+}$.
Proof.
Write $f_{m}=f_{s+1/2,m}$. The case when $i=0$ follows from the definition of $f_{m}$. For $i=1$, it is easy to check that the assumptions on $p$ and $m$ imply that
$g_{1}$ $\;=\;g_{0}|T_{p,s+1/2}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot g_{0}$
$\;=\;\left(p^{2s-1}q^{-mp^{2}}+\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}q^{-m}+O(q)\right)-\left(\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}q^{-m}+O(q)\right)$
$\;=\;p^{2s-1}q^{-mp^{2}}+O(q),$
from which the result follows. Now, for $i\geq 1$, it is easy to check that by induction hypothesis,
$g_{i+1}\;=\;g_{i}|T_{p,s+1/2}-p^{2s-1}\cdot g_{i-1}\;=\;p^{(2s-1)(i+1)}q^{-mp^{2i+2}}+O(q).$
The result then follows. ∎
Now, recall that the Shimura lift is Hecke-equivariant. That is, for $s\in\mathbb{Z}$, $f\in M_{s+1/2}^{!,+}$, and a rational prime $p$
$\mathcal{S}_{d_{0}}(f|T_{p,s+1/2})\;=\;\mathcal{S}_{d_{0}}(f)|T_{p,2s},$
where $d_{0}$ is some fundamental discriminant with $(-1)^{s}d_{0}>0$.
Proposition 8.6.
Let $s\in\mathbb{Z}_{\geq 2}$ and let $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$. Let $m\in\mathbb{Z}^{+}$ with $(-1)^{s-1}m\equiv 0,1\pmod{4}$ and let $p$ be a prime such that
  1. (1)
    $p^{2}\nmid m$, or
  2. (2)
    if $p=2$ and $4{\hskip 1.0pt}|{\hskip 1.0pt}m$, then $(-1)^{s-1}{\hskip 1.0pt}\frac{m}{4}\equiv 2,3\pmod{4}$.
Let
$F_{m}\;:=\;\mathcal{S}_{d_{0}}f_{s+1/2,m}$
and let $\lambda$ be a relation for $S_{2s}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{m}|\lambda)\;\equiv\;\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot a_{np^{l-1}}(F_{m}|\lambda)\pmod{p^{(2s-1)l}}$.
In particular, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{m}|\lambda)\;\equiv\;0\pmod{p^{(s-1)l}}$.
Proof.
Let $g_{i}\in M_{s+1/2}^{!,+}$ be as in Lemma 8.5 and let $G_{i}=\mathcal{S}_{d_{0}}g_{i}$. In particular,
$G_{0}$ $\;=\;F_{m}$
$G_{1}$ $\;=\;G_{0}|T_{p,2s}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot G_{0}$
$G_{i+1}$ $\;=\;G_{i}|T_{p,2s}-p^{2s-1}\cdot G_{i-1}.$
Let $\widetilde{G}_{i}=G_{i}|\lambda$. Then, since $\lambda$ commutes with $T_{p,2s}$, we have
$\widetilde{G}_{i+1}$ $\;=\;\widetilde{G}_{i}|T_{p,2s}-p^{2s-1}\cdot\widetilde{G}_{i-1}.$
Write $H_{i}=\widetilde{G}_{i}|U_{p}-p^{2s-1}\widetilde{G}_{i-1}$. Then,
$H_{i+1}\;=\;\widetilde{G}_{i+1}|U_{p}-p^{2s-1}\widetilde{G}_{i}\;=\;\left(\widetilde{G}_{i}|(U_{p}+p^{2s-1}V_{p})-p^{2s-1}\widetilde{G}_{i-1}\right)\hskip-2.27626pt\big|U_{p}-p^{2s-1}\widetilde{G}_{i}\;=\;H_{i}|U_{p}.$
It follows that
$H_{1}|U_{p}^{l-1}\;=\;H_{l}.$
Now, since the Shimura lift preserves cusp forms and $\lambda$ annihilates cusp forms, it follows from Lemma 8.5 that
$\widetilde{G}_{i}\;=\;p^{(2s-1)i}(\mathcal{S}f_{s+1/2,mp^{2i}})|\lambda\;\equiv\;0\pmod{p^{(2s-1)i}}$.
Hence,
$H_{l}\;=\;\widetilde{G}_{l}|U_{p}-p^{2s-1}\widetilde{G}_{l-1}\;\equiv\;\widetilde{G}_{l}|T_{p,2s}-p^{2s-1}\widetilde{G}_{l-1}\;\equiv\;\widetilde{G}_{l+1}\;\equiv\;0\pmod{p^{(2s-1)(l+1)}}$. (29)
Now,
$F_{m}|\lambda U_{p}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot F_{m}|\lambda$ $\;=\;\widetilde{G}_{0}|U_{p}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}G_{0}$
$\;=\;\widetilde{G}_{1}-p^{2s-1}\widetilde{G}_{0}|V_{p}\;\equiv\;0\pmod{p^{2s-1}}$,
and for $l\geq 2$,
$F_{m}|\lambda U_{p}^{l}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot F_{m}|\lambda U_{p}^{l-1}$ $\;=\;\widetilde{G}_{1}|U_{p}^{l-1}-p^{2s-1}\widetilde{G}_{0}|U_{p}^{l-2}$
$\;=\;H_{1}|U_{p}^{l-2}$
$\;=\;H_{l-1}\;\equiv\;0\pmod{p^{(2s-1)(l+1)}}$.
The result then follows. ∎
Proposition 8.7.
Let $s\in\mathbb{Z}_{\geq 2}$ and let $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$. Let $m\in\mathbb{Z}^{+}$ with $(-1)^{s-1}m\equiv 0,1\pmod{4}$ and let $p$ be a prime such that
  1. (1)
    $p^{2}\nmid m$, or
  2. (2)
    if $p=2$ and $4{\hskip 1.0pt}|{\hskip 1.0pt}m$, then $(-1)^{s-1}{\hskip 1.0pt}\frac{m}{4}\equiv 2,3\pmod{4}$.
Let $t\in\mathbb{Z}$ with $t\geq 0$ and let
$F_{mp^{2t}}\;:=\;\mathcal{S}_{d_{0}}f_{s+1/2,mp^{2t}}.$
Let $\lambda$ be a relation for $S_{2s}$. Then, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{mp^{2t}}|\lambda)\;\equiv\;\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}\cdot a_{np^{l-1}}(F_{mp^{2t}}|\lambda)\pmod{p^{(2s-1)(l-t)}}$.
In particular, for all $n,l\in\mathbb{Z}^{+}$,
$a_{np^{l}}(F_{mp^{2t}}|\lambda)\;\equiv\;0\pmod{p^{(s-1)(l-t)}}$.
Proof.
We will prove this by induction on $t$. Note that the base case when $t=0$ follows from Proposition 8.6. $G_{i}$ and $\widetilde{G}_{i}=G_{i}|\lambda$ as in the proof of Proposition 8.6 with respect to $m$. Since the Shimura lift preserves cusp forms and $\lambda$ annihilates cusp forms, it follows from Lemma 8.5 that
$\widetilde{G}_{i}\;=\;p^{(2s-1)i}(\mathcal{S}_{d_{0}}f_{s+1/2,mp^{2i}})|\lambda\;=\;p^{(2s-1)i}F_{mp^{2i}}|\lambda.$
Let $H_{i}=\widetilde{G}_{i}|U_{p}-p^{2s-1}\widetilde{G}_{i-1}$ as before. Then,
$\widetilde{G}_{t}|U_{p}^{l}\;=\;(H_{t}+p^{2s-1}\widetilde{G}_{t-1})|U_{p}^{l-1}\;=\;H_{t+l-1}+p^{2s-1}\widetilde{G}_{t-1}|U_{p}^{l-1}.$
By (29),
$H_{t+l-1}\;\equiv\;0\pmod{p^{(2s-1)(t+l)}}$.
Hence,
$F_{mp^{2t}}|\lambda U_{p}^{l}\;=\;p^{-(2s-1)t}H_{t+l-1}+F_{mp^{2t-2}}|\lambda U_{p}^{l-1}\;\equiv\;F_{mp^{2t-2}}|\lambda U_{p}^{l-1}\pmod{p^{(2s-1)l}}$,
so
$F_{mp^{2t}}|\lambda U_{p}^{l}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}F_{mp^{2t}}|\lambda U_{p}^{l-1}\;\equiv\;$
$\hskip 56.9055ptF_{mp^{2t-2}}|\lambda U_{p}^{l-1}-\big(\tfrac{(-1)^{s-1}m}{p}\big)p^{s-1}F_{mp^{2t-2}}|\lambda U_{p}^{l-2}\pmod{p^{(2s-1)l}}$.
The result now follows from induction. ∎
Proposition 8.8.
Let $s\in\mathbb{Z}_{\geq 2}$ and let $d_{0}$ be a fundamental discriminant with $(-1)^{s}d_{0}>0$. Let $m\in\mathbb{Z}^{+}$ with $(-1)^{s-1}m\equiv 0,1\pmod{4}$ and let $A$ be the largest integer such that $A^{2}{\hskip 1.0pt}|{\hskip 1.0pt}m$ and $(-1)^{s-1}m/A^{2}\equiv 0,1\pmod{4}$. Let
$F\;:=\;A^{s-1}\mathcal{S}_{d_{0}}f_{s+1/2,m}$
and let $\lambda$ be a relation for $S_{2s}$. Then, $F|\lambda$ is (s1)(s-1)-magnetic, i.e., $n^{s-1}{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}(F|\lambda)$ for all $n\in\mathbb{Z}^{+}$.
Proof.
Fix a prime $p$. It suffices to prove that
$a_{np^{l}}(F|\lambda)\;\equiv\;0\pmod{p^{(s-1)l}}$
for all $n,l\in\mathbb{Z}^{+}$. Let $t=v_{p}(A)$ and let $m^{\prime}=m/p^{2t}$. Then, $m^{\prime}$ satisfies that $(-1)^{s-1}m^{\prime}\equiv 0,1\pmod{4}$. We will prove that the pair $(p,m^{\prime})$ satisfies the assumptions in Proposition 8.7.
Suppose that $p^{2}|m^{\prime}$. Then, $(Ap)^{2}|m$. If $p\geq 3$, then $(-1)^{s-1}m/(Ap)^{2}\equiv 0,1\pmod{4}$, contradicting the minimality of $A$. Hence, $p=2$. Write $A^{\prime}=A/p^{2t}$. By the maximality of $A$,
$(-1)^{s-1}\frac{m}{(2A)^{2}}\;=\;(-1)^{s-1}\frac{m^{\prime}}{4}\cdot\frac{1}{A^{\prime 2}}\;\equiv\;2,3\pmod{4}$.
Since $2\nmid A^{\prime}$, this implies that $(-1)^{s-1}m^{\prime}/4\equiv 2,3\pmod{4}$. Hence, the pair $(p,m^{\prime})$ satisfies the assumptions in Proposition 8.7.
Now, applying Proposition 8.7 with $m^{\prime}p^{2t}$ yields that
$a_{np^{l}}(F_{m^{\prime}p^{2t}}|\lambda)\;\equiv\;0\pmod{p^{(s-1)(l-t)}}$,
where $F_{m^{\prime}p^{2t}}=\mathcal{S}_{d_{0}}f_{s+1/2,m}$. The result then follows since $F=A^{s-1}F_{m^{\prime}p^{2t}}$ and $p^{t}|A$. ∎
8.3. Proof of Theorem 8.1 when $k/2$ is even
Adopt all the notations in Theorem 8.1. Assume that $s=k/2$ is even. As $s$ is even, one can apply Proposition 8.4 with $d_{0}=1$ and $d=D$ to obtain that
$\big(\mathcal{S}_{1}f_{s+1/2,-D}\big)(z)-h_{D}(z)\;=\;(-1)^{\frac{s}{2}}{\hskip 1.0pt}|D|^{-\frac{s}{2}}{\hskip 1.0pt}\mathrm{Tr}_{D,1}\big((\partial_{\tau}^{s-1}G_{2s})(z,\tau)\big),$
where $h_{D}:=h_{s,D,1}\in S_{2s}$ has integer coefficients by Lemma 8.3. Since $d_{0}=1$, we have that $\chi(Q)=1$ for all $Q\in\mathcal{Q}_{D}$. Hence, for an $\mathrm{SL}_{2}(\mathbb{Z})$-invariant function $f(\tau)$ on $\mathcal{H}$,
$\mathrm{Tr}_{D,1}(f)\;=\;\sum_{Q\in\mathcal{Q}_{D}}\frac{1}{w_{Q}}f(\alpha_{Q}).$ (30)
Write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}^{+}$ and some fundamental discriminant $D_{0}<0$. Then,
$\mathcal{Q}_{D}\;=\;\bigsqcup_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\frac{A}{A^{\prime}}\cdot\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}},$ (31)
where $\frac{A}{A^{\prime}}\cdot\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}=\{\frac{A}{A^{\prime}}\cdot Q\mid Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}\}$. One can thus further write (30) as
$\mathrm{Tr}_{D,1}(f)\;=\;\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\frac{2}{w_{A^{\prime 2}D_{0}}}f(\alpha_{Q}),$
where we use that $w_{A^{\prime 2}D_{0}}=2w_{Q}$ for all $Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}$.
To summarize, we obtain that
$\big(\mathcal{S}_{1}f_{s+1/2,-D}\big)(z)-h_{D}(z)$ $\;=\;(-1)^{\frac{s}{2}}{\hskip 1.0pt}|D|^{-\frac{s}{2}}{\hskip 1.0pt}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\frac{2}{w_{A^{\prime 2}D_{0}}}{\hskip 1.0pt}(\partial_{\tau}^{s-1}G_{2s})(z,\alpha_{Q})$
$\;=\;(-1)^{\frac{s}{2}}{\hskip 1.0pt}|D|^{-\frac{s}{2}}{\hskip 1.0pt}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\mathcal{G}_{2s,A^{\prime 2}D_{0}}(z).$
By the Möbius inversion formula,
$\mathcal{G}_{2s,A^{2}D_{0}}$ $\;=\;\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\mu\big(\tfrac{A}{A^{\prime}}\big)(-1)^{\frac{s}{2}}{\hskip 1.0pt}|A^{\prime 2}D_{0}|^{\frac{s}{2}}(\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}-h_{A^{\prime 2}D_{0}})$
$\;=\;(-1)^{\frac{s}{2}}|D_{0}|^{\frac{s}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\mu\big(\tfrac{A}{A^{\prime}}\big)A^{\prime s}(\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}-h_{A^{\prime 2}D_{0}}).$
Now, each $\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}$ and $h_{A^{\prime 2}D_{0}}$ have integer coefficients, so it follows that $\widetilde{\mathcal{G}}_{k,D}=|D_{0}|^{-\frac{k}{4}}\mathcal{G}_{k,D}=|D_{0}|^{-\frac{s}{2}}\mathcal{G}_{2s,A^{2}D_{0}}$ has integer coefficients. Since each $h_{A^{\prime 2}D_{0}}\in S_{2s}$ and $\lambda$ annihilates $S_{2s}$, we obtain that
$\widetilde{\mathcal{G}}_{2s,A^{2}D_{0}}|\lambda\;=\;(-1)^{\frac{s}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\mu\big(\tfrac{A}{A^{\prime}}\big)A^{\prime s}{\hskip 1.0pt}\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}|\lambda.$
As $s$ is even, each $A^{\prime s-1}{\hskip 1.0pt}\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}|\lambda$ is (s1)(s-1)-magnetic by Proposition 8.8, so it follows that $\widetilde{\mathcal{G}}_{k,D}|\lambda$ is $\frac{k-2}{2}$-magnetic. Part (2) of Theorem 8.1 now follows from Proposition 8.6.
8.4. Proof of Theorem 8.1 when $k/2$ is odd
Adopt all the notations in Theorem 8.1. Assume that $s=k/2$ is odd. Write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}^{+}$ and some fundamental discriminant $D_{0}<0$ as before. As $s$ is odd, one can apply Proposition 8.4 with $d_{0}=D_{0}$ and $d=A^{2}$ to obtain that
$\big(\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}\big)(z)-h_{D}(z)\;=\;(-1)^{\frac{s+1}{2}}A^{-s}{\hskip 1.0pt}|D_{0}|^{\frac{s-1}{2}}{\hskip 1.0pt}\mathrm{Tr}_{A^{2},D_{0}}\big((\partial_{\tau}^{s-1}G_{2s})(z,\tau)\big),$
where $h_{D}:=h_{s,A^{2},D_{0}}\in S_{2s}$ has integer coefficients by Lemma 8.3. In this case, by (31) and Lemma 8.9, we have that for an $\mathrm{SL}_{2}(\mathbb{Z})$-invariant function $f(\tau)$ on $\mathcal{H}$,
$\mathrm{Tr}_{A^{2},D_{0}}(f)\;=\;\sum_{Q\in\mathcal{Q}_{D}}\frac{1}{w_{Q}}\chi(Q)f(\alpha_{Q})\;=\;\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\frac{2}{w_{A^{\prime 2}D_{0}}}\big(\tfrac{D_{0}}{A/A^{\prime}}\big){\hskip 1.0pt}f(\alpha_{Q}).$
To summarize, we obtain that
$\big(\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}\big)(z)-h_{D}(z)$ $\;=\;(-1)^{\frac{s+1}{2}}A^{-s}{\hskip 1.0pt}|D_{0}|^{\frac{s-1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\hskip-5.69046pt\frac{2}{w_{A^{\prime 2}D_{0}}}\big(\tfrac{D_{0}}{A/A^{\prime}}\big){\hskip 1.0pt}(\partial_{\tau}^{s-1}G_{2s})(z,\alpha_{Q})$
$\;=\;(-1)^{\frac{s+1}{2}}A^{-s}{\hskip 1.0pt}|D_{0}|^{\frac{s-1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\big(\tfrac{D_{0}}{A/A^{\prime}}\big)\mathcal{G}_{2s,A^{\prime 2}D_{0}}(z).$
Hence,
$A^{s}(\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}-h_{D})\;=\;(-1)^{\frac{s+1}{2}}|D_{0}|^{\frac{s-1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\big(\tfrac{D_{0}}{A/A^{\prime}}\big)\mathcal{G}_{2s,A^{\prime 2}D_{0}}.$
Now, both $\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}$ and $h_{D}$ have integer coefficients, so it follows from induction (on $A$) that $\widetilde{\mathcal{G}}_{k,D}=|D_{0}|^{\frac{k-2}{4}}\mathcal{G}_{k,D}=|D_{0}|^{\frac{s-1}{2}}\mathcal{G}_{2s,A^{2}D_{0}}$ has integer coefficients. Similarly as before, we obtain that
$A^{s}{\hskip 1.0pt}\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}|\lambda\;=\;(-1)^{\frac{s+1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\big(\tfrac{D_{0}}{A/A^{\prime}}\big)\widetilde{\mathcal{G}}_{2s,A^{\prime 2}D_{0}}|\lambda.$
As $s$ is odd, $A^{s-1}\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}|\lambda$ is (s1)(s-1)-magnetic by Proposition 8.8, so it follows from induction that $\widetilde{\mathcal{G}}_{k,D}|\lambda$ is $\frac{k-2}{2}$-magnetic. Similarly, Part (2) of Theorem 8.1 follows from Proposition 8.6 and induction.
The following is a lemma which should follow from the theory of the genus characters of quadratic forms. We still include a proof for completeness.
Lemma 8.9.
Let $D=A^{2}D_{0}<0$ be a discriminant with $A\in\mathbb{Z}^{+}$ and $D_{0}<0$ being a fundamental discriminant. Let $Q\in\mathcal{Q}_{D}$ and suppose that $Q$ is $\mathrm{SL}_{2}(\mathbb{Z})$-equivalent to $A_{Q}\cdot Q^{\prime}$ for some $A_{Q}\in\mathbb{Z}^{+}$ with $A_{Q}{\hskip 1.0pt}|{\hskip 1.0pt}A$ and $Q^{\prime}\in\mathcal{Q}_{D/A_{Q}^{2}}^{\mathrm{prim}}$. Then, $\chi_{A^{2},D_{0}}(Q)=\big(\frac{D_{0}}{A_{Q}}\big)$.
Proof.
Since $Q^{\prime}$ is primitive, there exists a prime $p^{\prime}$ with $(p^{\prime},D)=1$ such that $p^{\prime}=Q^{\prime}(x^{\prime},y^{\prime})$ for some $x^{\prime},y^{\prime}\in\mathbb{Z}$. Then,
$\chi_{A^{2},D_{0}}(Q)\;=\;\big(\tfrac{D_{0}}{A_{Q}\cdot p^{\prime}}\big)\;=\;\big(\tfrac{D_{0}}{A_{Q}}\big)\big(\tfrac{D_{0}}{p^{\prime}}\big),$
so it suffices to prove that $\big(\frac{D_{0}}{p^{\prime}}\big)=1$. Since $\gcd(x^{\prime},y^{\prime})=1$ (as $p^{\prime}=Q^{\prime}(x^{\prime},y^{\prime})$ is a prime), there exist $u,v\in\mathbb{Z}$ such that $x^{\prime}u-y^{\prime}v=1$. Then,
$Q^{\prime}(x^{\prime}X+vY,y^{\prime}X+uY)\;=\;p^{\prime}X^{2}+BXY+CY^{2}$
for some $B,C\in\mathbb{Z}$. In particular,
$D_{0}\cdot(A/A_{Q})^{2}\;=\;D/A_{Q}^{2}\;=\;B^{2}-4p^{\prime}C,$
so $D_{0}$ is a square modulo $p^{\prime}$, i.e., $\big(\frac{D_{0}}{p^{\prime}}\big)=1$. ∎